### v2

# Everybody Prune Now: Structured Pruning of LLMs with only Forward Passes

Abstract
Given the generational gap in available hardware between lay practitioners and the most endowed institutions, LLMs are becoming increasingly inaccessible as they grow in size. Whilst many approaches have been proposed to compress LLMs to make their resource consumption manageable, these methods themselves tend to be resource intensive, putting them out of the reach of the very user groups they target. In this work, we explore the problem of structured pruning of LLMs using only forward passes. We seek to empower practitioners to prune models so large that their available hardware has just enough memory to run inference. We develop Bonsai, a gradient-free, perturbative pruning method capable of delivering small, fast, and accurate pruned models. We observe that Bonsai outputs pruned models that (i) outperform those generated by more expensive gradient-based structured pruning methods, and (ii) are twice as fast (with comparable accuracy) as those generated by semi-structured pruning methods requiring comparable resources as Bonsai. We also leverage Bonsai to produce a new sub-2B model using a single A6000 that yields state-of-the-art performance on 4/6 tasks on the Huggingface Open LLM leaderboard.
Code for Bonsai can be found here
Machine Learning, ICML
[…]
1 Introduction
As large language models (LLMs) (OpenAI et al., 2023; Touvron et al., 2023; Gemini Team et al., 2023) continue to grow in size, the gap between models that achieve state-of-the-art performance and those that every-day machine learning (ML) practitioners can feasibly run on their available hardware continues to widen (Bender et al., 2021; Samsi et al., 2023). With the goal of democratizing access to these powerful models, previous research has proposed approaches such as pruning (Xia et al., 2022; Sun et al., 2023), distillation (Hinton et al., 2015; Gu et al., 2023) and quantization (Xiao et al., 2023) to create smaller models from larger pre-trained language models.
Figure 1: Perplexity versus inference speed-up of pruned models for methods that use only forward passes on the parent model. At any given target perplexity, Bonsai produces the fastest model, resulting in improved latency and throughput. Post-pruning adaptation is only possible for models below a certain size, given the available hardware. Circle sizes are proportional to the model’s memory footprint.
Table 1: Landscape of resource consumption (memory and compute) of different model compression methods at training time and the inference time resource consumption of the models they deliver. ✗  means the method incurs a prohibitive cost to the lay practitioner whilst ✓  denotes that it is a viable option with respect to that resource.
Regime Resource Approaches
Quantization (Mixed Precision)
Distillation
Unstructured Pruning
Gradient-Based Structured Pruning
Bonsai (Ours)
Train Memory
Compute
Inference Memory
Compute
Unfortunately, so far, these methods have fallen short of the goal of truly democratizing access to LLMs. In the process of producing a smaller model out of an LLM, methods such as distillation and structured pruning are inaccessible to the everyday practitioner due to their prohibitive resource consumption at training time. Specifically, pure distillation-based techniques require running LLMs to generate large amounts of teacher data (Jiao et al., 2019; Hsieh et al., 2023) whilst existing structured pruning approaches like LLM-Pruner (Ma et al., 2023) and LoRAPrune (Zhang et al., 2023) require several times more memory than is needed to run inference on the model being pruned. Though unstructured pruning (Frantar & Alistarh, 2023; Sun et al., 2023) and quantization are less restrictive at training time, the models they produce are not faster except in the presence of specialized hardware for the former (Mishra et al., 2021), whilst the latter can actually slow down inference due to added overhead (Dettmers et al., 2022). This limits the usefulness of these options for practitioners who are concerned with latency-critical applications. Table 1 summarizes the landscape of existing methods and their limitations.
[…]
The practitioner only has enough memory on their hardware to run inference on the model to be pruned. [^1] Specifically we assume a forward pass with a batch size of at least 1. Given the hardware and open source LLM landscape as of January 2024, for an Nvidia A6000 (48GB) GPU, the largest (full-precision) model this would be applicable to is LLaMA-7B
The above setting is evergreen. As state-of-the-art models become more compute intensive over time, the generational gap in hardware available to the lay practitioner versus the most resource endowed institutions is expected to persist or possibly widen.
In light of the proposed setting, we present Bonsai, a forward pass-only structured pruning approach that is capable of delivering fast, compact, and accurate pruned models under the memory limitations that are typical of consumer hardware. To decide which modules (attention head, rows in feedforward projection, etc.) of the LLM to prune, Bonsai estimates module importances perturbatively by generating sub-models and evaluating their performance (running inference). We make this approach tractable by contributing multiple techniques.
First, we treat the problem of inferring each module’s importance from the performance of generated sub-models as an under-determined regression problem. This enables us to estimate the importance of a large number of modules by exploring a manageable number of random sub-models. This is unlike past perturbative approaches, which prohibitively require roughly as many sub-models as there are modules to select from (Ancona et al., 2020), making them intractable for LLMs. Next, instead of instantiating sub-models by dropping modules with equal likelihood (Kang et al., 2023), we use informative priors derived from work on unstructured pruning (Han et al., 2015; Sun et al., 2023). We thus obtain better estimates of module relevance with fewer evaluated sub-models. Finally, unlike past gradient-free approaches that greedily make pruning decisions layer-by-layer (Dekhovich et al., 2021; Nova et al., 2023; Sun et al., 2023), Bonsai takes a holistic view to preserve the accuracy of the pruned model: modules across layers are removed and evaluated together and relevance scores are computed globally to make pruning decisions. To the best of our knowledge, these ingredients taken together make Bonsai the first successful attempt at scaling gradient-free structured pruning to LLMs.
We conduct several experiments that demonstrate Bonsai’s efficacy. Bonsai, which uses only forward passes for pruning, achieves comparable performance to 2:4 semi-structured sparsity with Wanda (Sun et al., 2023) (0.75×0.75$\times$0.75 × speedup after fine-tuning with LoRA(Hu et al., 2021)) whilst also speeding up inference by up to 1.58×1.58$\times$1.58 × compared to the parent model. Even when compared to SoTA gradient-based structured pruning methods like LLM-Pruner and LoRAPrune, Bonsai outperforms these methods—on 4 out of 6 evaluation settings in our experiments. We also use Bonsai to prune the $\sim$3B Phi-2 (Li et al., 2023) model to a $\sim$1.8B model that outperforms the best sub-2B parameter model on the Huggingface Open LLM leaderboard on 4 out of 6 tasks as of the time of submission. Based on these strong results and its usability under real-world memory constraints, we view Bonsai as a significant contribution to unlocking the power of LLMs for a broader spectrum of practitioners facing diverse hardware constraints.
2 Background on Pruning, Problem Definition and Notation Setup
We are given an LLM, $\mathbf{M}_{\theta}$, parameterized by $\theta\in\mathbb{R}^{D}$. Also provided is $U$, a utility function that evaluates a model’s performance on a target task. We are interested in pruning $\mathbf{M}_{\theta}$ to produce a smaller and faster but still performant (with respect to $U$) model under the constraint that we only have enough memory to run inference on $\mathbf{M}_{\theta}$. Even though we assume we can run $\mathbf{M}_{\theta}$ on available hardware, pruning can be critical for achieving latency targets, reducing compute burden, or making the model small enough to adapt to new (out-of-domain) tasks by gradient-based fine-tuning.
Unstructured pruning approaches compress $\mathbf{M}_{\theta}$ by removing individual parameters $\theta_{j}$ from the model. This results in the updated model consisting of sparsified weight matrices with a smaller memory footprint. Unfortunately, the updated model does not enjoy inference speedups except when specialized hardware is available and thus poses a compute burden during inference. Whilst semi-structured variants – those that remove parameters in patterns like 2:4 or 4:8 (Mishra et al., 2021) – achieve some speedup, these are modest compared to those achieved with structured pruning.
Structured pruning takes a more modular view of the units to be removed from $\mathbf{M}_{\theta}$. Take that $\mathbf{M}_{\theta}$ is made up of modules $\mathbf{m}=\{m_{i}\}_{i\in[N]}$ each with corresponding parameter counts $\mathbf{s}=\{s_{i}\}_{i\in[N]}$ such that $\sum_{i}s_{i}=D$. For a transformer model, $\mathbf{m}$ could consist of attention heads, dimensions of fully connected layers, or even whole layers. For simplicity, we assume that $\mathbf{m}$ is made up of non-overlapping modules. Structured pruning compresses $\mathbf{M}_{\theta}$ by finding accurate sub-models defined by subsets of $\mathbf{m}$: provided with $\bar{\mathbf{m}}\subseteq\mathbf{m}$, we can construct an updated model $\mathbf{M}_{|\bar{\mathbf{m}}}$ that is produced by dropping the modules not in $\bar{\mathbf{m}}$ from $\mathbf{M}$. Thus, given a sparsity target $p$, structured pruning can be cast as the following combinatorial optimization problem:
$\begin{split}\mathbf{m}^{\ast}&=\mathrm{argmax}_{\bar{\mathbf{m}}\in\mathcal{F}_{p}}\quad U\left(\mathbf{M}_{|\bar{\mathbf{m}}}\right)\qquad\text{where}\\ \mathcal{F}_{p}&=\bigg{\{}\bar{\mathbf{m}}\subseteq\mathbf{m}~{}\bigg{|}~{}\bigg{(}\sum_{[j:m_{j}\in\bar{\mathbf{m}}]}s_{j}\bigg{)}\leq(1-p)D\bigg{\}}\end{split}$ (1)
$\mathcal{F}_{p}$ consists of all sub-models that meet the sparsity threshold. Note that, in general, not only does $\mathbf{M}_{|\mathbf{m}^{\ast}}$ have a smaller memory footprint than $\mathbf{M}$, it is also faster to run inference on it since it has fewer modules.
Many structured pruning methods attempt to solve Equation 1 by gradient-guided optimization (or search) over the space of sub-models. However, since we are interested in the memory-constrained setting where computing gradients is not feasible, these methods cannot be used. We will thus focus on developing a memory-friendly structured pruning technique, but we will compare to semi-structured pruning methods since they also have minimal memory overhead.
3 Methodology
In this section, we will develop a structured pruning algorithm that relies exclusively on forward passes through the parent model.
3.1 Estimating module relevance with only forward passes
We have motivated the setting of pruning a model that is so large (relative to the amount of memory available to the practitioner), such that we can only run forward passes through it. This means that we have to solve Equation 1 by relying on only evaluations of $U$, as opposed to gradient-based optimization with respect to $U$. A brute force approach would enumerate all members of $\mathcal{F}_{p}$, find their corresponding performances and pick the best one. Unfortunately, this would be computationally infeasible since $|\mathcal{F}_{p}|$ is combinatorial in the size of the model. [^2] The typical fully connected (FC) sub-layer in an LLM has size $>10^{4}$ units. Pruning just 1 of these FCs to 0.5 sparsity in this way requires evaluating over $\approx$ ($10^{4}\choose 10^{3}$ subsets.
We propose a computationally tractable approach where we first perform a small number, $n$, of evaluations, where $n\ll|\mathcal{F}_{p}|$, to gather data for estimating the relevance of each module in $\mathbf{M}$ with respect to the metric $U$. Upon carrying out this estimation, we can greedily choose the member of $\mathcal{F}_{p}$ that has the highest total module relevance. Specifically, let us assume that we have estimated $\mathbf{\beta}=\{\beta_{i}\}_{i\in[N]}$ to be the relevance of each of the $N$ modules. We can generate an approximate solution to Equation 1 as:
$\mathbf{m}^{\ast}\approx\mathbf{m}^{\mathrm{approx}}=\mathrm{argmax}_{\bar{\mathbf{m}}\in\mathcal{F}_{p}}\quad\sum_{j\in\bar{\mathbf{m}}}\beta_{j}$ (2)
We note that Equation 2 is straightforward to solve, as it simply requires sorting $\beta_{j}$s and greedily selecting the top modules until the parameter count constraint is met (this greedy approach may slightly overshoot the sparsity constraint but since $s_{i}\ll(1-p)D~{}\forall~{}i$ for our settings of interest, the difference is not significant).
Estimating $\mathbf{\beta}$:
To obtain estimates of the module relevance scores $\mathbf{\beta}=\{\beta_{i}\}_{i\in[N]}$, we propose to generate and evaluate $n\ll|\mathcal{F}_{p}|$ sub-models $\bar{\mathbf{m}}_{k}$, and construct a dataset of the sampled sub-models and their corresponding evaluated performances: $\mathbb{D}=\left\{\bar{\mathbf{m}}_{k},U_{k}\right\}_{k\in[n]}$ where $U_{k}=U(\mathbf{M}_{|\bar{\mathbf{m}}_{k}})$. Given $\mathbb{D}$, we treat the problem of estimating $\mathbf{\beta}$ as an under-specified regression problem :
$\begin{split}\hat{\mathbf{\beta}}&=\mathrm{argmin}_{\mathbf{\beta}\in\mathbb{R}^{N}}\quad\frac{1}{n}\sum_{\left(\bar{\mathbf{m}}_{k},U_{k}\right)\in\mathbb{D}}\left(U_{k}-\beta^{T}\alpha_{\bar{\mathbf{m}}_{k}}\right)^{2}+\gamma\|\mathbf{\beta}\|\\ \end{split}$ (3)
where $\alpha_{\bar{\mathbf{m}}_{k}}\in\mathbb{R}^{N}~{}\big{|}~{}\left(\alpha_{\bar{\mathbf{m}}_{k}}\right)_{i}=\mathbf{1}[i\in\bar{\mathbf{m}}_{k}]$, is the binary vector that has 0 at indices where modules have been dropped.
Implementing a sub-model $\bar{\mathbf{m}}_{k}$ as a binary mask $\alpha_{\mathbf{m}_{k}}$ is key to practically realizing our approach. We never actually fully instantiate sub-models as this would be prohibitively expensive. Instead, we create sub-models virtually by zeroing out the outputs of the components to be pruned so they have no effect on the final model output.
3.2 Selecting sub-models for evaluation
An as yet unexplored design choice is how to choose the $n$ candidate sub-models for evaluation. A naive approach would be to sample uniformly from the space of all feasible sub-models. However, we can show that this is sub-optimal. Let us assume $m_{i}$ is a module that is critical for good performance under evaluation with $U$. Since $n<N$, it means that some modules will never be ”turned on” in the list of $n$ chosen sub-models. If $m_{i}$ happens to be one of these masked-out modules, the resulting estimate of $\hat{\beta_{i}}=0$ would in turn result in $\mathbf{m}^{\mathrm{approx}}$ being a poor estimate for $\mathbf{m}^{\ast}$. We therefore need to make a more informed choice to ensure that our estimates of $\mathbf{\beta}$ are accurate and useful.
Given a module $m_{i}$, we can set the likelihood of it being present in any of the $n$ sampled sub-models to be proportional to some measure $\rho_{i}$ which captures our prior belief that $m_{i}$ is a useful module. To define $\rho_{i}$, we can turn to metrics from the unstructured pruning literature. Specifically, we set $\rho_{i}$ to be a module-level analogue of any of the parameter-level unstructured pruning metrics like Wanda (Sun et al., 2023) or output activation magnitude. As an illustrative example, consider that we have shallow network with 1 hidden layer (of dimension $d$) we wish to prune. If output activation magnitude is the metric for our prior, we would calculate the activation vector for the $d$ hidden dimensions – $\hat{\mathbf{a}}$ – averaged across multiple samples. The likelihood that the $i$th column of the matrix $W\in R^{d\times d}$ is left unpruned would be:
$\rho_{i}\propto\hat{\mathbf{a}}_{i}=\frac{1}{B}\sum_{b}\bigg{|}\sigma\bigg{(}\big{(}W^{T}[i,:]\big{)}x_{b}\bigg{)}\bigg{|}$
$\sigma\text{ is the nonlinearity}$. Sampling sub-models according to $\mathbf{\rho}$ means we are more likely to explore models that already have a good performance $U$. The priors $\rho_{i}$ can be computed via forward passes through the original unmodified model $\mathbf{M}_{\theta}$ which allows us to respect memory constraints.
To enhance the efficiency of our method, instead of considering all active modules for pruning, in each layer we consider the bottom $2p$ fraction of modules [^3] The choice of $2p$ is arbitrary. Practitioners can run hyper-parameter searches to find more optimal values for their settings. ranked according to our prior $\rho$. Thus, the top $1-2p$ fraction of modules remain fixed whilst the rest are perturbed and their relevance scores are computed and compared for pruning. This helps reduce the space of possible sub-models being considered for evaluation. For the bottom $2p$ fraction being compared, whenever we generate a mask $\alpha_{\mathbf{m}_{k}}$ with sparsity level $p$, we also generate its complement $\alpha^{c}_{\mathbf{m}_{k}}$, which is obtained by flipping the values in $\alpha_{\mathbf{m}_{k}}$ (except for the fixed $1-2p$ fraction of entries). Covert & Lee (2020) show that this technique can help reduce the variance of the estimator obtained from regression with binary inputs.
Algorithm 1 Bonsai
1:  Input: Model [$\mathbf{M}_{\theta}$], sub-models per iteration [$n_{\mathrm{iter}}$] Sparsity per iteration [$p_{\mathrm{iter}}$], Target sparsity [$p$] Module list [$\mathbf{m}$]
2:  for $l=1$ to $\lceil\frac{p}{p_{\mathrm{iter}}}\rceil$ do
3:     $\mathbb{\rho}^{l}\leftarrow$ Calculate unstructured pruning metric for all modules in $\mathbf{m}$
4:     $\bar{\mathbb{\rho}}^{l}\leftarrow$ Fix the top $(1-2p_{\mathrm{iter}})$ of $\mathbb{\rho}^{l}$ to $\infty$
5:     Sample $\{\mathbf{\bar{m}}_{i}\}_{[\frac{n_{\mathrm{iter}}}{2}]}$ sub-models according to $\bar{\mathbb{\rho}}^{l}$
6:     Generate $\{\mathbf{\bar{m}}^{c}_{i}\}_{[\frac{n_{\mathrm{iter}}}{2}]}$ complements of $\{\mathbf{\bar{m}}_{i}\}_{[\frac{n_{\mathrm{iter}}}{2}]}$
7:     Run forward pass on each sub-model and compute $U$. Construct $\mathbb{D}^{l}=\left\{\bar{\mathbf{m}}_{i},U_{i}\right\}_{[\frac{n_{\mathrm{iter}}}{2}]}\bigcup\left\{\bar{\mathbf{m}}^{c}_{k},U_{k}\right\}_{[\frac{n_{\mathrm{iter}}}{2}]}$
8:     $\beta^{\l}\leftarrow\mathrm{Regress}\left(\mathbb{D}^{l}\right)$
9:     $\{m_{\mathrm{pruned}}\}\leftarrow$ sort $\beta^{\l}$ and remove the bottom $k$ modules that correspond to $p_{\mathrm{iter}}$ fraction of the model.
10:     $\mathbf{m}\leftarrow$ update module list to exclude $\{m_{\mathrm{pruned}}\}$
11:  end for
12:  Output: Pruned model $\mathbf{M}_{|\mathbf{m}}$
3.3 Iterated Pruning
Previous work on gradient-based pruning (Anwar et al., 2017; Frankle & Carbin, 2018) have shown that taking an iterated approach to pruning yields improved results over pruning directly to the target sparsity $p$. Similarly, we define an updated pruning fraction $p_{\mathrm{iter}}<p$ and perform $\mathrm{iter}=\lceil\frac{p}{p_{\mathrm{iter}}}\rceil$ steps where we explore $n_{\mathrm{iter}}=\lceil\frac{n}{\mathrm{iter}}\rceil$ sub-models at a time. At the beginning of each iteration, we re-estimate the priors $\mathbf{\rho}$ for the unpruned modules and use them to select the $n_{\mathrm{iter}}$ sub-models to evaluate.
We combine the recipes developed in Sections 3.1, 3.2 and 3.3 together to produce Bonsai [^4] Structural pruning is a canonical way of giving a bonsai tree its shape hence the name., our gradient-free structural pruning algorithm. Algorithm 1 specifies Bonsai in detail.
4 Post-pruning adaptation
Depending on the sparsity level $p$, it is possible to obtain a pruned model on which it is feasible to run a parameter-efficient finetuning method like LoRA (Hu et al., 2021) with the available hardware memory. In this case, we can fine-tune the pruned model (result of Algorithm 1) on the downstream task in order to recover some of the parent model performance that was degraded by pruning.
Like many past works (Sanh et al., 2020; Xia et al., 2022), we combine pruning with distillation by incorporating a distillation loss in the training objective during fine-tuning of the pruned model. Let $\mathcal{L}_{\mathrm{task}}$ be the loss function over the task data and $\mathcal{L}_{\mathrm{distill}}$ be the distillation objective. We optimize the following post-pruning objective: $\mathcal{L}^{\mathrm{post-prune}}=\mathcal{L}_{\mathrm{task}}+\lambda\mathcal{L}_{\mathrm{distill}}$. Using $i$ to index the task data, we have:
$\mathcal{L}_{\mathrm{distill}}=\sum_{i}D_{\mathrm{KL}}\bigg{(}\mathrm{logits}^{i}\left(\mathbf{M}_{|\mathbf{m}^{\mathrm{approx}}}\right)~{}\|~{}\mathrm{logits}^{i}\left({\mathbf{M}}\right)\bigg{)}$, where $\lambda$ is a scalar weighting that can be optimized via cross-validation. Also, distillation can be performed without significant memory overhead by apriori caching the logits from the parent model $\mathbf{M}$ instead of hosting the model in memory during fine-tuning.
Table 2: Wikitext-2 perplexity for 50% sparsity of LLaMA-2 7B with end-to-end latency speedups (relative to LLaMA-2 7B). We use Phi-2 (Li et al., 2023) at the target size as a strong model for comparison. For semi-structured (Wanda 2:4) pruning, after fine-tuning, the learned LoRA weights cannot be merged with the primary model weights else the model reverts back to being dense; this leads to the reported slow downs. All methods use forward passes only on the parent model.
[…]
Phi-2 3B $8.69$ 1.24×1.24$\times$1.24 ×
LLaMA-2 7B Pruned
Wanda 2:4 3B $10.52$ 1.14×1.14$\times$1.14 ×
$8.34$ 0.75$\times$
Bonsai 3B $8.89$ $\boldsymbol{1.58}\times$
5 Experimental Details
All experiments are conducted on a single NVIDIA A6000 (48Gb) GPU. In all experiments with Bonsai, we prune (1) the heads in the self-attention layers (2) the dimensions of the fully connected layers.
Comparing Forward Pass Only Methods: We focus our first set of experiments on comparing methods that can be run without gradient-based optimization. We consider pruning the LLaMA-2 7B model (Touvron et al., 2023) to 50% sparsity. We evaluate on the Wikitext-2 (Merity et al., 2016) validation dataset and so our signal for pruning, $U$, is the language modelling performance on the training set. When measuring speedups, we consider end-to-end latency of running inference on model.sequence_length chunks of the Wikitext-2 validation set. See Appendix A.2 for details about the hyper-parameters used for all methods in this experiment and for specifics about fine-tuning.
Comparing Structured Pruning Approaches: We compare Bonsai to the following gradient-based structured pruning approaches: LLM-Pruner (Ma et al., 2023) and LoRA-Prune (Zhang et al., 2023). We prune the LLaMA-1 7B model (Touvron et al., 2023) to 50% sparsity since this was the model version available at time of release of the above methods. We compare these methods on Wikitext-2 and also on 6 tasks from the Eleuther LLM Evaluation Harness (Gao et al., 2023). Pruning signal for Wikitext-2 task is the same as the above experiment. For the Eleuther Harness tasks, we use language modelling performance on the C4 (Raffel et al., 2020) dataset as pruning signal. We also do parameter efficient finetuning on our pruned model with 30K 512-length sequences from this corpus. Specific hyper-parameters for Bonsai for this experiment can be found in Appendix A.3.
6 Main Results
6.1 Bonsai produces fast and performant models with only forward passes
In Table 2, we explore options that are available to practitioners when they can only run forward passes on the parent model. Here, LLaMA-2 7B is compressed to 50% sparsity.
We compare Bonsai to the semi-structured variant of Wanda (Sun et al., 2023). Before fine-tuning, the model produced by Wanda 2:4 achieves a speedup over the parent model (1.14$\times$) but is slower and less accurate than Bonsai’s pruned model (8.89ppl vs 10.52ppl). While the performance gap can be bridged by LoRA finetuning (10.52 $\rightarrow$ 8.34), the adapted semi-structured model experiences a drastic slowdown (0.75$\times$), since the learned low-rank matrices cannot be merged with the original sparsified ones without reverting back to dense computation. Thus, LoRA fine-tuned Wanda 2:4 is twice slower (0.48$\sim$ 0.48$\times$ 0.48 ×) than the model from Bonsai.
In this memory-constrained setting, practitioners could alternatively opt for a pre-existing model of the target size instead of pruning a larger model. We compare the Bonsai-pruned model to Phi-2 (Li et al., 2023), a strong representative pre-existing model of similar size. As can be seen in Table 2, Bonsai is able to generate a model that is as accurate (0.2 difference in ppl) yet significantly faster (1.58$\times$ vs. 1.24$\times$ speedup), thus making it a competitive option to consider even if a model already exists at the target size.
Table 3: LLaMA-1 (50% sparsity) after post-pruning adaptation with LoRA. ${}^{\dagger}$ indicate results as reported by Zhang et al. (2023).
Method Foward-only Wikitext-2 $\downarrow$ BoolQ HellaSwag WinoGrande ARC-e ARC-c Average $\uparrow$
LLaMA1-7B (Touvron et al., 2023) - 5.68 75.05 56.92 69.93 75.34 41.89 63.83
${}^{\dagger}$ (Ma et al., 2023) 16.41 60.28 47.06 53.43 45.96 29.18 47.18
${}^{\dagger}$ (Zhang et al., 2023) 11.60 61.88 47.86 55.01 45.13 31.62 48.30
Bonsai 10.92 67.22 43.09 61.64 54.92 26.28 50.63
6.2 Bonsai outperforms other structured pruning approaches
We next investigate the setting where the practitioner has enough memory to run gradient-based structured pruning methods. We compare Bonsai to recent SoTA methods: LLM-Pruner (Ma et al., 2023) and LoRA-Prune (Zhang et al., 2023). Since these approaches report their results for the LLaMA-1 only, we prune LLaMA-1 7B to 50% sparsity.
As evinced by Table 3, Bonsai outperforms existing structured pruning methods for LLMs even though it exclusively uses forward passes in the pruning stage. We attribute the superior performance of Bonsai to the fact that its pruning decisions are informed by directly exploring the space of sub-models whilst the other approaches resort on inaccurate proxies of module relevance in order to reduce the memory overhead of a fully gradient-based optimization approach. Specifically, LoRA-Prune’s criterion for deciding which modules to prune uses gradient signals not from the modules directly but from a surrogate low-rank matrix. On the other hand, LLM-Pruner attempts to model the impact of removing a module on the model loss via a Taylor expansion and substitutes the LLM’s intractable (memory-wise) Hessian term with the Fisher information matrix which we posit leads to sub-optimal pruning decisions.
6.3 Bonsai can produce compressed models with strong zero-shot abilities
Table 4: Phi-2 pruned to 35% sparsity compared to the best sub-2B parameter models on the Hugging Face OpenLLM Leaderboard. At the time of submission, $\dagger$ was the best sub-2B parameter model on the Leaderboard.
Generation Multiple Choice (MC)
Model Size GSM8k (5-shot) ARC-c (25-shot) Winogrande (5-shot) Hellaswag (10-shot) Truthful-QA (0-shot) MMLU (5-shot) MC Average $\uparrow$
Phi-2 (Li et al., 2023) 2.7B 54.81 61.09 74.35 75.11 44.47 58.11 62.63
StableLM-2-1_6$\dagger$ 1.6B $\boldsymbol{17.44}$ 43.34 64.56 $\boldsymbol{70.45}$ 36.78 38.95 50.82
Ours 1.8B 6.37 $\boldsymbol{47.44}$ $\boldsymbol{68.35}$ 65.09 $\boldsymbol{42.20}$ $\boldsymbol{40.53}$ $\boldsymbol{52.72}$
Considerable amounts of compute and data, beyond what is available to lay practitioners, are needed to train LLMs with strong zero-shot capabilities (OpenAI et al., 2023; Gemini Team et al., 2023). In this section, we demonstrate that Bonsai can empower everyday practitioners to produce strong and compact models with competitive zero-shot abilities by simply pruning bigger models on their available hardware.
We use Bonsai to prune a $\approx$3B Phi-2 model to $\approx$1.8B (35% sparsity). Values for the Bonsai hyper-parameters in this experiment are in Appendix A.4. Since its relatively small, the 1.8B pruned model can be fully fine-tuned on 1 A6000 GPU over 100k sequences of 2,048 tokens from the C4 dataset. As can be seen from Table 4, our pruned model achieves strong zero-shot performance compared to StableLM, the leading sub-2B parameter LLM on the Hugging Face OpenLLM leaderboard (Gao et al., 2023) — outperforming it on 4 out of 6 tasks as of January 2024. Interestingly, one exception to the general trend of Bonsai’s superior performance is the GSM-8K dataset, which is a mathematical reasoning dataset that requires generation of a long reasoning chains. We posit that this is because currently, Bonsai prunes with respect to language modeling likelihood, as opposed to reasoning accuracy. An interesting avenue for future work is to prune to improve maintain reasoning ability.
7 Analysis
In this section, we conduct various ablative experiments to understand how the methodological ingredients from Section 3 contribute to make Bonsai effective.
Do we need both perturbative and regressive components of Bonsai? Figure 2 shows that both components are key to obtaining a good pruned model. Removing the estimation of module importances via regression leads to a degradation in performance ($61.6$ ppl $\rightarrow$ $146.6$ ppl). More degradation ($146.6$ ppl $\rightarrow$ $405.7$ ppl) is encountered if we do not perform any perturbative evaluations on the parent model but simply just prune according to the prior $\rho$ as computed from the unperturbed parent model.
Figure 2: LLaMA-2 7B pruned to 50% sparsity. Perturbatively evaluating the impact of removing modules and then using regression to estimate module importances are both key in making Bonsai effective. Details of experiment configuration can be found in Appendix B. No post-pruning adaptation is performed.
Table 5: Varying the number of perturbative evaluations. Wikitext-2 perplexity of LLaMA-2 7B pruned to 50% sparsity. See Appendix G for details. No post-pruning adaptation is performed.
$\mathrm{ns}=50$ $\mathrm{ns}=200$ $\mathrm{ns}=1000$
PPL ($\downarrow$) NaN $61.63$ $22.09$
How many perturbative samples are sufficient? We investigate the number of perturbative samples required to obtain good estimates of module importances after performing regression as in Equation 3. Our results are shown in Table 5. As expected, performance improves as we increase the number of sub-models explored. We note that the number of samples we explore, $\mathrm{ns}$, is significantly less than the number of candidate modules at each iteration ($N\approx 70k$), nevertheless Bonsai is able to deliver a performant pruned model because of the recipes developed in Section 3.
How much performance is recovered by post-pruning adaptation? During iterative pruning, Bonsai damages the parent model by removing modules but does not perform intermittent retraining to recover lost performance since even intermediate models may be too large for fine-tuning. Even so, as Table 6 shows, the final model produced by Bonsai has reasonable performance without fine-tuning. We attribute this to the general robustness of LLMs, the redundancy of modules with respect to target end-tasks and Bonsai’s ability to identify good candidates for pruning. If the pruned model is small enough in size, we can perform either full fine-tuning or parameter-efficient fine-tunin to recover more performance, as can be seen from Table 6.
Table 6: Impact of post-pruning adaptation of LLaMA-2 7B pruned to 50% sparsity. See Appendix E for more details.
Method Wikitext-2 PPL
No Post-Pruning Adaptation $19.47$
Post-Pruning Finetuning $10.39$
   +  Distillation $8.89$
Figure 3: LLaMA-2 7B pruned to 50% sparsity. See Appendix F for details of experiment configuration and definitions of module-level analogues of Wanda and Activation Magnitude.
What is the impact of the choice of metric for the prior $\rho$? We investigate three different choices of metrics for defining $\rho$. Figure 3 shows that using the module-level analogue of Wanda (Sun et al., 2023) yields the best performance, both before and after post-pruning adaptation. This indicates that Wanda is a strong signal for efficiently estimating the importance of model units (whether at the parameter- or module-level).
Should Bonsai prune iteratively? Table 7 demonstrates the benefits of using Bonsai  in an iterative fashion. Pruning slowly ($p_{\mathrm{iter}}=0.05$) yields the best results, but this comes at the cost of increasing the total time to prune the model. The performance gap between values of $p_{\mathrm{iter}}$ persists even after post-pruning adaptation, indicating that slower pruning allows for more accurate estimates of module importance.
Table 7: Varying $p_{\mathrm{iter}}$. Wikitext-2 perplexity of LLaMA-2 7B pruned to 50% sparsity. See Appendix C for experiment details.
$p_{\mathrm{iter}}=0.05$ $p_{\mathrm{iter}}=0.1$ $p_{\mathrm{iter}}=0.2$
w/o Adapt $19.47$ $61.63$ $209.44$
w Adapt $8.89$ $9.15$ $9.57$
8 Related Work
8.1 Pruning
Unstructured pruning:
Whilst structured pruning approaches remove entire model components like layers (Xu et al., 2020; Xia et al., 2022), dimensions of linear layers (Wang et al., 2019) or attention heads (Michel et al., 2019; Held & Yang, 2022), unstructured pruning (Han et al., 2015; Frankle & Carbin, 2018; Benbaki et al., 2023; Sun et al., 2023) removes individual parameters of the model. These approaches achieve memory savings by inducing sparsity in the model weights, but they generally do not result in actual model speedups except when specialized hardware is available (Mishra et al., 2021). Proposed semi-structured sparsity methods (Mishra et al., 2021) such as 2:4 and 4:8 patterns do result in faster inference, but the speedup gains they achieve are far from the idealized 2×2$\times$2 ×. There are several gradient-free, unstructured pruning approaches. Since, as far as we know, there are no scalable gradient-free structured pruning alternatives outside of Bonsai, we compare it with Wanda (Sun et al., 2023), a SoTA unstructured/semi-structured pruning technique for LLMs.
Structured pruning without backward passes:
To the best of our knowledge, all existing structured pruning techniques for large (over 1B scale) language models, like LoRAPrune (Zhang et al., 2023) and LLM-Pruner (Ma et al., 2023), are gradient-based. Thus, under the memory setting we consider, these methods cannot be applied since their memory requirements well exceed the memory needed for inference on the model to be pruned. For smaller language models like BERT, Nova et al. (2023) recently proposed Kernelized Convex Masking (KCM) for gradient-free structured pruning. Unfortunately, to prune a fully connected layer with $K$ intermediate units, KCM requires instantiating a $K\times K$ coefficient matrix (for each layer). Whilst this is feasible for BERT models, a typical LLM (OpenAI et al., 2023; Touvron et al., 2023) has $K\sim 10^{4}-10^{5}$ which would make the size of each per-layer coefficient matrix comparable to the size of the LLM itself. In computer vision, there exist several perturbative gradient-free structured pruning techniques (Ancona et al., 2020; Dekhovich et al., 2021). However, these methods have been exclusively applied to small vision models (VGG; (Simonyan & Zisserman, 2014), WideResnet; (Zagoruyko & Komodakis, 2016)) and, taken as is, would not scale to LLMs. We introduce several techniques to make Bonsai scale to large language models. A key difference between Bonsai and these perturbative methods is that pruning decisions are made globally instead of on a local layer-by-layer basis (Ancona et al., 2020; Dekhovich et al., 2021). A failure mode of layer-by-layer pruning is that if a module, say $m_{i}$, which is in the top $p$ fraction of relevant modules when the whole model is considered but in the bottom $1-p$ fraction within its layer will not be included in the pruned model — a clearly sub-optimal choice. Bonsai enables us to obtain globally meaningful estimates of module relevance that are comparable across layers.
8.2 Model compression beyond pruning
Distillation:
Originally proposed by Hinton et al. (2015), Sanh et al. (2019) have shown that it is possible to reach similar performances on many downstream tasks using much smaller language models pre-trained with knowledge distillation. The resulting models are lighter and faster at inference time. Unfortunately, as noted by Xia et al. (2022), distilling a pre-trained model to a student model that has been initialized from scratch requires lots of compute and data to recover reasonable levels of teacher performance. To reduce this resource burden, Sanh et al. (2020); Lagunas et al. (2021); Xia et al. (2022) combine distillation with gradient-based structured pruning, so that the student model does not have to be initialized from scratch. Bonsai also leverages distillation, but only during the post-pruning adaptation phase. This maintains the constraint that the pruning process is gradient-free. And since the pruned model is smaller than the parent model, we can safely perform distillation and parameter efficient fine-tuning (Hu et al., 2021; He et al., 2021) within the bounds of available memory.
Quantization:
Li et al. (2020) introduced a training strategy that showed that heavily compressed, large models achieved higher accuracy than lightly compressed, small models, providing both efficiency and high-accuracy results on NLP tasks at once. Dettmers et al. (2022) introduced a new method called ‘LLM.int8()‘, which allows loading large models with 16 or 32-bit weights and using them immediately for inference without any performance degradation by combining vector-wise quantization with mixed-precision decomposition. Xiao et al. (2023) enabled 8-bit weight, 8-bit activation (W8A8) quantization for LLMs with >100B parameters, which smoothed activation outliers by migrating the quantization difficulty from activations to weights with a mathematically equivalent transformation. Quantization tends to be complementary to pruning, allowing independent exploration and subsequent combinations of both techniques. Some quantization approaches, specifically mixed precision ones like Dettmers et al. (2022) (Appendix D), can produce resulting models that are slower than the parent model, though it has also been shown that it can result in more accurate models than pruning (Kuzmin et al., 2023). This paper focuses on gradient-free structured pruning. We leave combining Bonsai with quantization for future work.
9 Conclusion, Limitations, and Future Work
In this work, we have presented Bonsai, the first tractable gradient-free method for structured pruning of LLMs. Bonsai allows practitioners to prune any LLM as long as they have enough memory to run inferences on the model. Through a battery of experiments, we have shown that the models produced by Bonsai are small, fast, and accurate.
A primary limitation of Bonsai is its runtime. As Section 7 shows, performance improves when we increase the number of sub-models explored, slow down pruning, and use more data samples; but this comes at the cost of increased runtime. Our result in Table 2 took $\sim 40$ hours to obtain, whilst Wanda-2:4 takes on the order of minutes. Since the model we produce is twice as fast, this trade-off is worthwhile when amortized over inference on the pruned models.
Bonsai presents several avenues for expansion by future work. First, though sub-models are sampled from an informative prior, $\rho$, the sampling process is not adaptive. Bonsai could further be strengthened by dynamically exploring the space of sub-models. Next, because of memory constraints, Bonsai does not fine-tune the model during iterative pruning to recover degraded performance. However, forward-pass-only fine-tuning approaches like MeZO (Malladi et al., 2023) exist that can be used to dynamically update the model during pruning.
10 Acknowlegements
This work was supported in part by the National Science Foundation grants IIS1705121, IIS1838017, IIS2046613, IIS2112471, and funding from Meta, Morgan Stanley, Amazon, and Google. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of any of these funding agencies. We are grateful for helpful feedback from Mingjie Sun, Victor Akinwande, Asher Trockman, Afshin Rostamizadeh and Daniel Glasner
11 Broader Impact Statement
Network pruning is an effective compression strategy to reduce the inference costs of large language models. However, existing techniques for structured pruning impose expensive memory constraints that may render them out of reach in practice. By reducing such constraints, work aims to democratize pruning to provide practitioners the ability to produce their own pruned LLMs for real-world applications. While we ultimately view this democratization as a benefit, we note that a necessary outcome is that it may also make it more feasible for malicious actors to readily use LLMs in practice. Finally, we note that our work has primarily focused on the axes of efficiency and utility (e.g., perplexity) in assessing performance; recent work has shown that compression may also have outsized effects on issues such as model fairness and robustness, which would be interesting additional aspects to consider in future study.
Appendix A Main Experiment Details
A.1 Hyper-parameters for all Bonsai regression during pruning
When using Bonsai, we estimate $\beta$ from $\mathbb{D}^{l}$ by performing linear regression via gradient descent with Adam (Kingma & Ba, 2014). We cross-validate over the following set of hyper-parameters. Note that doing this cross-validation takes much less time than the time needed to construct the dataset $\mathbb{D}^{l}$.
Table 8: Bonsai hyper-parameters for regression. This applies to all experiments unless otherwise specified
$\gamma$(Regression Weight) Learning rate Batch Size Epochs
{100, 0, 1e-4} {100, 10, 1, 0.1} {32, 64, 128} 50
During cross validation, we choose the model whose predictions have the best Kendall rank correlation co-efficient (Kendall, 1948) with the target. We do this because we do not care about matching $U_{k}$ exactly for each sub-model $k$; we rather care that our learned $\beta$ predicts the correct rankings amounts sub-models, which would denote that $\beta$ reasonably models relative module importances.
In general, we use $\ell_{1}$-norm regularization on $\beta$ for all experiments. For the Phi-2 experiment in Section 6.3, we find that $\ell_{2}$-norm works better.
A.2 Forward Pass Only Experiments
Table 10 show the Bonsai hyperparameters we used for the experiments in Section 6.1.
Table 9: Bonsai hyper-params for forward only experiments
Table 10: Bonsai  fine-tuning HP for pruned LLaMA family models
$p_{\mathrm{iter}}$ $\mathrm{ns}_{\mathrm{sub-models}}$ $\mathrm{ns}_{\mathrm{data}}$ Metric for $\rho$
0.05 200 32 (per-iter) Wanda
LR rank LoRA-$\alpha$ $\lambda$ (Distill Weight) LoRA Modules
1e-4 128 4$\times$×rank 0.01 All Modules
Table 10: Bonsai  fine-tuning HP for pruned LLaMA family models
For Wanda(Sun et al., 2023), we use the default hyper-parameters specified by the paper repo here for pruning. For fine-tuning, we use rank = 64. We apply LoRA to only the q_proj and v_proj matrices in each layer of the pruned LLaMA model – this is unlike with Bonsai where we fine-tune all modules. We cannot do same because since the Wanda model just produces sparse matrices, the matrices instantiated during the backward pass are the same sizes as the sparsified matrices and thus occupy more memory (compared to our approach that actually makes the matrices smaller in dimension instead of sparsifying). We are also unable to perform distillation on the Wanda models due to this reason. For fine-tuning the Phi-2 model on Wikitext-2, we use the same hyper-parameters as Bonsai in Table 10.
A.3 Experiments comparing to Gradient based structured pruning
We compare to LoRA-Prune and LLM-Pruner. We take their performance results directly from the LoRA-Prune paper. Whilst we use 1 A6000 GPU (48G) for all experiments, LoRA-Prune uses A100 GPU (80G) for pruning LLaMA-1 7B.
All Bonsai hyper-parameters are the same as Appendix A.2 except for $\mathrm{ns}_{\mathrm{sub-models}}$ which we set to $1000$.
A.4 Phi-2 pruning experiment details
For the experiment in Section 6.3, All Bonsai hyper-parameters are the same as Appendix A.2 except for the following changes:
  • $\mathrm{ns}_{\mathrm{sub-models}}=2000$
  • $p_{\mathrm{iter}}=0.35$. We thus perform 1-shot pruning directly to the target sparsity of 35%. We find that this seems to work best for the Phi-2 model. We posit that this might be because the Phi-2 models use LayerNorm(Ba et al., 2016) whilst the other models we explore, LLaMA and Mistral use RMSNorm.
  • Due to its relatively small size, the 1.8B pruned model can be fully fine-tuned on a single A6000 GPU over 100k sequences of length 2,048 tokens from the C4 dataset instead of using LoRA.
Appendix B Impact of regression and perturbation ablation details
For the experiment in Section 3, All Bonsai hyper-parameters are the same as Appendix A.2 except $p_{\mathrm{iter}}=0.1$ to speed up pruning.
A simple alternative to Bonsai is to leverage the prior $\rho$, computed from the unperturbed parent model, and make pruning decisions exclusively according to this. This is the No Perturbation + No Regression baseline in Figure 2. This approach has quite poor performance. We can further improve this baseline by adding back perturbative aspect where we prune the parent model according to $\rho^{\prime}$ which is computed by aggregating the $\rho$ metric computed over the perturbed models. Note that we use a Wanda based metric to define $\rho$ for this experiment. Module level analogues of the unstructured pruning metrics we explore are defined in Appendix F.
Table 11: Experiment on linear regression to estimate module importances. Wikitext-2 Perplexity. LLaMA-2 7B pruned to 50% sparsity
Linear Regression Relative Speepdup w/o Post-Pruning Adaptation w Post-Pruning Adaptation
No $2.06$ $146.57$ $9.68$
Yes $1.77$ $61.63$ $9.15$
Appendix C Varying the pruning fraction per-iteration
For the experiment in Section 3, All Bonsai hyper-parameters are the same as Appendix A.2 except we vary $p_{\mathrm{iter}}$.
Table 12: Varying the fraction pruned at a time. Wikitext-2 Perplexity. LLaMA-2 7B pruned to 50% sparsity
Prune Frac Relative Speepdup w/o Post-Pruning Adaptation w Post-Pruning Adaptation
0.05 $1.58$ $19.47$ $8.89$
0.1 $1.77$ $61.63$ $9.15$
0.20 $1.67$ $209.44$ $9.57$
Appendix D Varying the number of calibration data points for pruning
All Bonsai hyper-parameters are the same as Appendix A.2 except we vary $\mathrm{ns}_{\mathrm{data}}$ and $p_{\mathrm{iter}}=0.1$ to speed up pruning.
Table 13: How many data-points to consider during forward passes. Wikitext-2 Perplexity. Llama-2 7B pruned to 50% sparsity
$\mathrm{ns}_{\mathrm{data}}$ w/o Adapt w Adapt
8 $130.04$ $9.45$
32 $61.63$ $9.15$
Appendix E Post-pruning adaptation
For this experiment, All Bonsai hyper-parameters are the same as Appendix A.2.
Appendix F Impact of prior
For this experiment, All Bonsai hyper-parameters are the same as Appendix A.2 except we vary $\rho$
F.1 $\rho$ is Activation Magnitude
MLP / Fully Connected Module: Let $d$ be the intermediate dimension of the MLP to be pruned. Note that for all transformer models evaluate, the MLP components are 2 layer and thus have a single intermediate dimension. For any data-sample sequence $b$, we flatten model activation at this point $\mathbf{a}\in\mathbb{R}^{B\times S\times d}\rightarrow\mathbb{R}^{BS\times d}$ and then compute the following averaged activation magnitude :
$\left(\rho\in\mathbb{R}^{d}\right)\propto\hat{\mathbf{a}}=\frac{1}{B}\sum_{b}\mathrm{Mean}\bigg{(}\big{|}\mathbf{a}_{b}\big{|},\mathrm{axis=}0\bigg{)}$ (4)
Self-Attention Module: For any data-sample sequence $b$, the output of the self attention module before the final output projection is $\mathbf{a}\in\mathbb{R}^{B\times S\times d_{h}\times h}$ where $h$ is the number of attention heads and $d_{h}$ is the size of each head’s output. We can flatten $\mathbf{a}\in\mathbb{R}^{B\times S\times d_{h}\times h}\rightarrow\mathbb{R}^{BSd_{h}\times h}$ and then use the same formula as Equation 5 above to calculate $\rho$.
$\left(\rho\in\mathbb{R}^{h}\right)\propto\hat{\mathbf{a}}$ (5)
F.2 $\rho$ is Wanda (Sun et al., 2023)
MLP / Fully Connected Module: Let $d$ be the intermediate dimension of the MLP to be pruned. Let $W\in\mathbb{R}^{d\times o}$ be the output projection matrix for the MLP. For any data-sample sequence $b$, we flatten model activation before the final output, $\mathbf{a}\in\mathbb{R}^{B\times S\times d}\rightarrow\mathbb{R}^{BS\times d}$ and then compute the following metric which is a module level analogue of Wanda:
$\begin{split}\left(\rho\in\mathbb{R}^{d}\right)\propto\hat{\mathbf{a}}&=\frac{1}{o}\sum_{o}\mathbf{a}^{o}\\ \mathbf{a}^{o}&=\bigg{|}W[:,o]\bigg{|}\odot\mathrm{RootMeanSquare}\bigg{(}\mathbf{a},\mathrm{axis=}0\bigg{)}\end{split}$ (6)
Self-Attention Module: Let $W\in\mathbb{R}^{d\times o}$ be the output projection matrix for the self-attention module. For any data-sample sequence $b$, the output of the self attention module before the final output projection is $\mathbf{a}\in\mathbb{R}^{B\times S\times d_{h}\times h}$ where $h$ is the number of attention heads and $d_{h}$ is the size of each head’s output. We can flatten $\mathbf{a}\in\mathbb{R}^{B\times S\times d_{h}\times h}\rightarrow\mathbb{R}^{BSd_{h}\times h}$ and then use the same formula as Equation 5 above to calculate $\rho\in\mathbb{R}^{h}$.
Appendix G How many perturbative samples are reasonable?
For this experiment, All Bonsai hyper-parameters are the same as Appendix A.2 except $p_{\mathrm{iter}}=0.1$ to speed up pruning.
Table 14: Varying the number of sub-models generated. Wikitext-2 Perplexity. LLaMA-2 7B pruned to 50% sparsity
Num Samples w/o Post-Pruning Adaptation w Post-Pruning Adaptation
1000 $22.09$ $9.25$
200 $61.63$ $9.15$
50 NaN $9.24$
Using $\mathrm{ns}_{\mathrm{sub-models}}=50$ results in an model with NaN perplexity on the Wikitext validation set. We posit that this is because of the LLaMA models are half precision, and removing the wrong modules can result in activations going outside of the FP16 dynamic range for unique data-points. Note that we are able to recover good performance of the model after fine-tuning though (we do not observe NaNs with the Wikitext-2 training data). This indicates that Bonsai actually recovers good modules even using as few samples as 50 sub-models.
Appendix H Mistral-7B Experiment Details
In addition to the primary experiments on the LLaMA and Phi-2 models, supplementary experiments were performed on the Mistral-7B (Jiang et al., 2023) model in comparison with Wanda results on the stated model. We apply Bonsai with the same hardware and configuration settings as used for the LLaMA and Phi-2 experiments. We target different pruning fractions (0.05, 0.1, and 0.2) across different numbers of samples and masks per iteration to evaluate the method’s performance under varying sparsity conditions.
The Mistral-7B model architecture differs from the LLaMA architecture in its use of group query attention and sliding window attention in lieu of the standard self-attention used in most transformer-based models like LLaMA (Jiang et al., 2023). We factor these differences into consideration in the implementation of Bonsai for Mistral. For the experiments that produced the results below, all Bonsai hyper-parameters are the same as Appendix A.2.
Table 15 presents the test perplexity results for Mistral-7B under different pruning methods. Considering the fully-structured sparsity nature of Bonsai, it achieves a test perplexity of 47.5 without post-pruning adaptation, with 1.66$\times$ inference speedup. After performing post-pruning adaptation on our pruned Mistral-7B, perplexity dropped drastically to 10.08. Note that the reported results of Wanda-pruned Mistral-7B are not fine-tuned afterward; if they were, their results would be marginally better than Bonsai’s results. However, as shown in Table 2, latency speedup would have dropped rapidly, while Bonsai stays the same at 1.66$\times$.
Table 15: Test perplexity of Mistral-7B model on Wikitext-2 across fully-structured Bonsai and semi-structured Wanda methods.
Sparsity Level Method Test PPL
Original, unpruned Mistral-7B N/A N/A 5.245
Wanda semi-structured 2-4 magnitude 13.81
Wanda 12.38
SparseGPT 10.46
Bonsai (w/o Adaptation) structured 50% magnitude 67.48
Wanda 47.50
Bonsai (w/ Adaptation) structured 50% Wanda 10.08
We further investigate the pruning habits of Bonsai by examining the pruned layers of Mistral, as shown in Figure 4. We notice a recurring theme: when an attention layer is significantly altered, it leads to compensation in the next layers within the sequence. This adaptive behavior, termed the ”Hydra effect” by McGrath et al. (2023), implies that the layers within a language model interact in a way that changes in one layer prompt adjustments in another. McGrath et al. (2023) specifically mentioned that when one attention layer was removed from a language model, the model was still able to self-repair and produce similar outputs; but it did so by relying more heavily on other layers.
Figure 4: Mistral’s pruned attention layers. The heavily pruned layers are usually preceded by or sandwiched between lightly-pruned layers, exhibiting the self-repairing ”Hydra effect” (McGrath et al., 2023).

### v3

@@
 Abstract
-Given the generational gap in available hardware between lay practitioners and the most endowed institutions, LLMs are becoming increasingly inaccessible as they grow in size.
-Whilst many approaches have been proposed to compress LLMs to make their resource consumption manageable, these methods themselves tend to be resource intensive, putting them out of the reach of the very user groups they target.
-In this work, we explore the problem of structured pruning of LLMs using only forward passes.
-We seek to empower practitioners to prune models so large that their available hardware has just enough memory to run inference.
-We develop Bonsai, a gradient-free, perturbative pruning method capable of delivering small, fast, and accurate pruned models.
-We observe that Bonsai outputs pruned models that (i) outperform those generated by more expensive gradient-based structured pruning methods, and (ii) are twice as fast (with comparable accuracy) as those generated by semi-structured pruning methods requiring comparable resources as Bonsai.
-We also leverage Bonsai to produce a new sub-2B model using a single A6000 that yields state-of-the-art performance on 4/6 tasks on the Huggingface Open LLM leaderboard.
-Code for Bonsai can be found here
-Machine Learning, ICML
-[…]
-[…]
+Structured pruning is a promising approach to create smaller, faster LLMs.
+However, existing methods typically rely on backward passes, which can inflate memory requirements and compute costs.
+In this work we introduce Bonsai, a gradient-free structured pruning method that eliminates the need for backpropagation, significantly reducing memory requirements and compute costs while achieving state-of-the-art pruning performance.
+Bonsai uses forward-pass-only perturbative pruning to enable efficient compression of large models on a broader range of hardware configurations.
+Unlike existing structured pruning approaches, Bonsai not only achieves better compression with fewer resources, but also produces models that are twice as fast as those generated by semi-structured pruning.
+As a concrete demonstration, we use Bonsai to prune an 8B LLaMA-3 model to 50% sparsity on a single A6000 GPU–a task infeasible with backprop-based methods, which require $2$-3×3$\times$3 × memory.
+Our results show that removing backprop as a requirement not only enables pruning larger models on constrained hardware but can also lead to state-of-the-art efficiency and performance.
 1 Introduction
-As large language models (LLMs) (OpenAI et al., 2023; Touvron et al., 2023; Gemini Team et al., 2023) continue to grow in size, the gap between models that achieve state-of-the-art performance and those that every-day machine learning (ML) practitioners can feasibly run on their available hardware continues to widen (Bender et al., 2021; Samsi et al., 2023).
-With the goal of democratizing access to these powerful models, previous research has proposed approaches such as pruning (Xia et al., 2022; Sun et al., 2023), distillation (Hinton et al., 2015; Gu et al., 2023) and quantization (Xiao et al., 2023) to create smaller models from larger pre-trained language models.
 Figure 1: Perplexity versus inference speed-up of pruned models for methods that use only forward passes on the parent model.
 At any given target perplexity, Bonsai produces the fastest model, resulting in improved latency and throughput.
-Post-pruning adaptation is only possible for models below a certain size, given the available hardware.
-Circle sizes are proportional to the model’s memory footprint.
-Table 1: Landscape of resource consumption (memory and compute) of different model compression methods at training time and the inference time resource consumption of the models they deliver.
-✗  means the method incurs a prohibitive cost to the lay practitioner whilst ✓  denotes that it is a viable option with respect to that resource.
-Regime Resource Approaches
-Quantization (Mixed Precision)
-Distillation
-Unstructured Pruning
-Gradient-Based Structured Pruning
-Bonsai (Ours)
-Train Memory
-Compute
-Inference Memory
-Compute
-Unfortunately, so far, these methods have fallen short of the goal of truly democratizing access to LLMs.
-In the process of producing a smaller model out of an LLM, methods such as distillation and structured pruning are inaccessible to the everyday practitioner due to their prohibitive resource consumption at training time.
-Specifically, pure distillation-based techniques require running LLMs to generate large amounts of teacher data (Jiao et al., 2019; Hsieh et al., 2023) whilst existing structured pruning approaches like LLM-Pruner (Ma et al., 2023) and LoRAPrune (Zhang et al., 2023) require several times more memory than is needed to run inference on the model being pruned.
-Though unstructured pruning (Frantar & Alistarh, 2023; Sun et al., 2023) and quantization are less restrictive at training time, the models they produce are not faster except in the presence of specialized hardware for the former (Mishra et al., 2021), whilst the latter can actually slow down inference due to added overhead (Dettmers et al., 2022).
-This limits the usefulness of these options for practitioners who are concerned with latency-critical applications.
-Table 1 summarizes the landscape of existing methods and their limitations.
+Circle sizes $\propto$ model’s memory footprint.
 […]
-The practitioner only has enough memory on their hardware to run inference on the model to be pruned.
-[^1] Specifically we assume a forward pass with a batch size of at least 1.
-Given the hardware and open source LLM landscape as of January 2024, for an Nvidia A6000 (48GB) GPU, the largest (full-precision) model this would be applicable to is LLaMA-7B
-The above setting is evergreen.
-As state-of-the-art models become more compute intensive over time, the generational gap in hardware available to the lay practitioner versus the most resource endowed institutions is expected to persist or possibly widen.
-In light of the proposed setting, we present Bonsai, a forward pass-only structured pruning approach that is capable of delivering fast, compact, and accurate pruned models under the memory limitations that are typical of consumer hardware.
-To decide which modules (attention head, rows in feedforward projection, etc.) of the LLM to prune, Bonsai estimates module importances perturbatively by generating sub-models and evaluating their performance (running inference).
-We make this approach tractable by contributing multiple techniques.
-First, we treat the problem of inferring each module’s importance from the performance of generated sub-models as an under-determined regression problem.
-This enables us to estimate the importance of a large number of modules by exploring a manageable number of random sub-models.
-This is unlike past perturbative approaches, which prohibitively require roughly as many sub-models as there are modules to select from (Ancona et al., 2020), making them intractable for LLMs.
-Next, instead of instantiating sub-models by dropping modules with equal likelihood (Kang et al., 2023), we use informative priors derived from work on unstructured pruning (Han et al., 2015; Sun et al., 2023).
-We thus obtain better estimates of module relevance with fewer evaluated sub-models.
-Finally, unlike past gradient-free approaches that greedily make pruning decisions layer-by-layer (Dekhovich et al., 2021; Nova et al., 2023; Sun et al., 2023), Bonsai takes a holistic view to preserve the accuracy of the pruned model: modules across layers are removed and evaluated together and relevance scores are computed globally to make pruning decisions.
-To the best of our knowledge, these ingredients taken together make Bonsai the first successful attempt at scaling gradient-free structured pruning to LLMs.
-We conduct several experiments that demonstrate Bonsai’s efficacy.
-Bonsai, which uses only forward passes for pruning, achieves comparable performance to 2:4 semi-structured sparsity with Wanda (Sun et al., 2023) (0.75×0.75$\times$0.75 × speedup after fine-tuning with LoRA(Hu et al., 2021)) whilst also speeding up inference by up to 1.58×1.58$\times$1.58 × compared to the parent model.
-Even when compared to SoTA gradient-based structured pruning methods like LLM-Pruner and LoRAPrune, Bonsai outperforms these methods—on 4 out of 6 evaluation settings in our experiments.
-We also use Bonsai to prune the $\sim$3B Phi-2 (Li et al., 2023) model to a $\sim$1.8B model that outperforms the best sub-2B parameter model on the Huggingface Open LLM leaderboard on 4 out of 6 tasks as of the time of submission.
-Based on these strong results and its usability under real-world memory constraints, we view Bonsai as a significant contribution to unlocking the power of LLMs for a broader spectrum of practitioners facing diverse hardware constraints.
-2 Background on Pruning, Problem Definition and Notation Setup
+In light of this challenge, we present Bonsai, a structured pruning approach that operates exclusively with forward passes through the parent model.
+By eliminating the need for backward passes for pruning, Bonsai dramatically reduces memory requirements that enable pruning of models that would otherwise be infeasible on available hardware, and can do so while outperforming existing structured and unstructured pruning approaches (Fig 1).
+Rather than computing gradients to determine module importance, Bonsai employs an innovative perturbative approach, estimating module importances by evaluating the performance of sub-models using only inference.
+We make this perturbative approach tractable through several key innovations.
+First, we formulate module importance estimation as an underdetermined regression problem, allowing us to infer the importance of a large number of modules by exploring a manageable number of random sub-models.
+This fundamentally differs from previous pertubative approaches, which require evaluating approximately as many sub-models as there are modules (Ancona et al., 2020), making them intractable for LLMs.
+Second, we use informative priors derived from unstructured pruning (Han et al., 2015; Sun et al., 2023) to guide sub-model exploration, yielding better estimates of module relevance with fewer evaluations, instead of instantiating sub-models by dropping modules with equal likelihood (Kang et al., 2023).
+Finally, unlike prior gradient-free approaches that prune layer-by-layer (Dekhovich et al., 2021; Nova et al., 2023; Sun et al., 2023), Bonsai takes a holistic view across the entire model: modules across layers are removed and evaluated together, with relevance scores computed globally to make pruning decisions that better preserve accuracy.
+Through extensive experiments, we demonstrate Bonsai’s effectiveness.
+When compared to other structured pruning methods that also avoid backward passes, including a forward-only structured pruning version of Wanda (Sun et al., 2023), Bonsai consistently produces models with better perplexity at any given speedup target.
+Wen the resulting pruned model small enough to be fine-tuned on available hardware, Bonsai compared to semi-structured pruning methods like Wanda (Sun et al., 2023)—which is also fine-tuned—Bonsai not only achieves similar performance, but also delivers 2×2$\times$2 × faster inference (an overall inference speed-up of 1.58×1.58$\times$1.58 × compared to the parent model).
+Even when compared to the state-of-the-art gradient-based structured pruning methods like LLM-Pruner and LoRAPrune, Bonsai produces models that outperform these more memory-intensive approaches on 4 out of 6 evaluations, despite using significantly less memory during pruning.
+Our results lead us to explore Bonsai’s practical utility by pruning the  3B Phi-2 (Li et al., 2023) model to a  1.8B model that performs competitively against other sub-2B parameter models on the Huggingface Open LLM leaderboard.
+Overall, by reducing the memory and computational requirements of structured pruning, Bonsai enables more efficient adaptation of LLMs across a wider range of hardware configurations.
+This makes these LLMs more accessible while maintaining their performance advantages, representing an important step forward for practical LLM compression.
+2 Methodology
+We first cover background in LLM pruning, and then discuss our structured pruning method that exclusively performs inference on the parent model.
+2.1 Background on Pruning, Problem Definition and Notation Setup
 We are given an LLM, $\mathbf{M}_{\theta}$, parameterized by $\theta\in\mathbb{R}^{D}$.
@@
 Thus, given a sparsity target $p$, structured pruning can be cast as the following combinatorial optimization problem:
-$\begin{split}\mathbf{m}^{\ast}&=\mathrm{argmax}_{\bar{\mathbf{m}}\in\mathcal{F}_{p}}\quad U\left(\mathbf{M}_{|\bar{\mathbf{m}}}\right)\qquad\text{where}\\ \mathcal{F}_{p}&=\bigg{\{}\bar{\mathbf{m}}\subseteq\mathbf{m}~{}\bigg{|}~{}\bigg{(}\sum_{[j:m_{j}\in\bar{\mathbf{m}}]}s_{j}\bigg{)}\leq(1-p)D\bigg{\}}\end{split}$ (1)
+$\begin{split}\mathbf{m}^{\ast}&=\mathrm{argmax}_{\bar{\mathbf{m}}\in\mathcal{F}_{p}}\quad U\left(\mathbf{M}_{|\bar{\mathbf{m}}}\right)\quad\text{where}\quad\mathcal{F}_{p}=\big{\{}\bar{\mathbf{m}}\subseteq\mathbf{m}~{}\big{|}~{}\big{(}\sum_{[j:m_{j}\in\bar{\mathbf{m}}]}s_{j}\big{)}\leq(1-p)D\big{\}}\end{split}$ (1)
 $\mathcal{F}_{p}$ consists of all sub-models that meet the sparsity threshold.
@@
 However, since we are interested in the memory-constrained setting where computing gradients is not feasible, these methods cannot be used.
-We will thus focus on developing a memory-friendly structured pruning technique, but we will compare to semi-structured pruning methods since they also have minimal memory overhead.
-3 Methodology
-In this section, we will develop a structured pruning algorithm that relies exclusively on forward passes through the parent model.
+Figure 2: Overview of Bonsai.
+Perturbations to the original model are generated according to an intelligent prior.
+These perturbed models are evaluated on downstream task and the collected data is used to build a simple linear model of module importances.
+Bonsai iteratively drops modules of least importance until the target model size is reached.
 2.2 Estimating module relevance with only forward passes
-We have motivated the setting of pruning a model that is so large (relative to the amount of memory available to the practitioner), such that we can only run forward passes through it.
-This means that we have to solve Equation 1 by relying on only evaluations of $U$, as opposed to gradient-based optimization with respect to $U$.
-A brute force approach would enumerate all members of $\mathcal{F}_{p}$, find their corresponding performances and pick the best one.
-Unfortunately, this would be computationally infeasible since $|\mathcal{F}_{p}|$ is combinatorial in the size of the model.
-[^2] The typical fully connected (FC) sub-layer in an LLM has size $>10^{4}$ units.
-Pruning just 1 of these FCs to 0.5 sparsity in this way requires evaluating over $\approx$ ($10^{4}\choose 10^{3}$ subsets.
-We propose a computationally tractable approach where we first perform a small number, $n$, of evaluations, where $n\ll|\mathcal{F}_{p}|$, to gather data for estimating the relevance of each module in $\mathbf{M}$ with respect to the metric $U$.
+We consider pruning a model too large to fit in memory, allowing only forward passes.
+Thus, we must solve Eq 1 using evaluations of $U$ rather than gradient-based optimization.
+A brute-force search over $\mathcal{F}_{p}$ is infeasible, as its size grows combinatorially with the model.
+For instance, pruning a single FC sub-layer in an LLM to 50% sparsity requires evaluating $\approx$ ($10^{4}\choose 10^{3}$ subsets.
+We propose a computationally tractable approach where we first perform a small number, $n$, of evaluations, where $n\ll|\mathcal{F}_{p}|$, to gather data for estimating the relevance of each module in $\bm{M}$ with respect to the metric $U$.
 Upon carrying out this estimation, we can greedily choose the member of $\mathcal{F}_{p}$ that has the highest total module relevance.
-Specifically, let us assume that we have estimated $\mathbf{\beta}=\{\beta_{i}\}_{i\in[N]}$ to be the relevance of each of the $N$ modules.
+Specifically, let us assume that we have estimated $\bm{\beta}=\{\beta_{i}\}_{i\in[N]}$ to be the relevance of each of the $N$ modules.
 We can generate an approximate solution to Equation 1 as:
-$\mathbf{m}^{\ast}\approx\mathbf{m}^{\mathrm{approx}}=\mathrm{argmax}_{\bar{\mathbf{m}}\in\mathcal{F}_{p}}\quad\sum_{j\in\bar{\mathbf{m}}}\beta_{j}$ (2)
-We note that Equation 2 is straightforward to solve, as it simply requires sorting $\beta_{j}$s and greedily selecting the top modules until the parameter count constraint is met (this greedy approach may slightly overshoot the sparsity constraint but since $s_{i}\ll(1-p)D~{}\forall~{}i$ for our settings of interest, the difference is not significant).
-Estimating $\mathbf{\beta}$:
-To obtain estimates of the module relevance scores $\mathbf{\beta}=\{\beta_{i}\}_{i\in[N]}$, we propose to generate and evaluate $n\ll|\mathcal{F}_{p}|$ sub-models $\bar{\mathbf{m}}_{k}$, and construct a dataset of the sampled sub-models and their corresponding evaluated performances: $\mathbb{D}=\left\{\bar{\mathbf{m}}_{k},U_{k}\right\}_{k\in[n]}$ where $U_{k}=U(\mathbf{M}_{|\bar{\mathbf{m}}_{k}})$.
-Given $\mathbb{D}$, we treat the problem of estimating $\mathbf{\beta}$ as an under-specified regression problem :
-$\begin{split}\hat{\mathbf{\beta}}&=\mathrm{argmin}_{\mathbf{\beta}\in\mathbb{R}^{N}}\quad\frac{1}{n}\sum_{\left(\bar{\mathbf{m}}_{k},U_{k}\right)\in\mathbb{D}}\left(U_{k}-\beta^{T}\alpha_{\bar{\mathbf{m}}_{k}}\right)^{2}+\gamma\|\mathbf{\beta}\|\\ \end{split}$ (3)
-where $\alpha_{\bar{\mathbf{m}}_{k}}\in\mathbb{R}^{N}~{}\big{|}~{}\left(\alpha_{\bar{\mathbf{m}}_{k}}\right)_{i}=\mathbf{1}[i\in\bar{\mathbf{m}}_{k}]$, is the binary vector that has 0 at indices where modules have been dropped.
-Implementing a sub-model $\bar{\mathbf{m}}_{k}$ as a binary mask $\alpha_{\mathbf{m}_{k}}$ is key to practically realizing our approach.
-We never actually fully instantiate sub-models as this would be prohibitively expensive.
-Instead, we create sub-models virtually by zeroing out the outputs of the components to be pruned so they have no effect on the final model output.
+$\bm{m}^{\ast}\approx\bm{m}^{\mathrm{approx}}=\mathrm{argmax}_{\bar{\bm{m}}\in\mathcal{F}_{p}}\quad\sum_{j\in\bar{\bm{m}}}\beta_{j}$ (2)
+Eq.
+2 is easily solved by sorting $\beta_{j}$s and greedily selecting top modules until the constraint is met.
+This may slightly exceed the sparsity limit, but the overshoot is negligible since $s_{i}\ll(1-p)D~{}\forall~{}i$ in our settings.
+Estimating $\bm{\beta}$: To estimate of the module relevance scores $\bm{\beta}\in\mathbb{R}^{N}$, we generate and evaluate $n\ll|\mathcal{F}_{p}|$ sub-models, constructing a dataset $\mathbb{D}=\left\{\bar{\bm{m}}k,U{k}\right\}{k\in[n]}$ where $U_{k}=U(\bm{M}_{|\bar{\bm{m}}_{k}})$.
+We then frame the estimation of $\bm{\beta}$ as an under-specified regression problem:
+$\begin{split}\hat{\bm{\beta}}&=\mathrm{argmin}_{\bm{\beta}\in\mathbb{R}^{N}}\quad\frac{1}{n}\sum_{\left(\bar{\bm{m}}_{k},U_{k}\right)\in\mathbb{D}}\left(U_{k}-\beta^{T}\alpha_{\bar{\bm{m}}_{k}}\right)^{2}+\gamma\|\bm{\beta}\|\\ \end{split}$ (3)
+where $\left(\alpha_{\bar{\bm{m}}_{k}}\right)_{i}=\bm{1}[i\in\bar{\bm{m}}_{k}]$, is the binary vector that has 0 at indices with modules dropped.
+Implementing a sub-model $\bar{\bm{m}}_{k}$ as a binary mask $\alpha_{\bm{m}_{k}}$ is key to practically realizing our approach.
+We never actually instantiate sub-models as this would be prohibitively expensive.
+Instead, we create them virtually by zeroing out the outputs of the parts to be pruned so they have no effect on the model output.
 2.3 Selecting sub-models for evaluation
-An as yet unexplored design choice is how to choose the $n$ candidate sub-models for evaluation.
-A naive approach would be to sample uniformly from the space of all feasible sub-models.
-However, we can show that this is sub-optimal.
-Let us assume $m_{i}$ is a module that is critical for good performance under evaluation with $U$.
-Since $n<N$, it means that some modules will never be ”turned on” in the list of $n$ chosen sub-models.
-If $m_{i}$ happens to be one of these masked-out modules, the resulting estimate of $\hat{\beta_{i}}=0$ would in turn result in $\mathbf{m}^{\mathrm{approx}}$ being a poor estimate for $\mathbf{m}^{\ast}$.
-We therefore need to make a more informed choice to ensure that our estimates of $\mathbf{\beta}$ are accurate and useful.
-Given a module $m_{i}$, we can set the likelihood of it being present in any of the $n$ sampled sub-models to be proportional to some measure $\rho_{i}$ which captures our prior belief that $m_{i}$ is a useful module.
-To define $\rho_{i}$, we can turn to metrics from the unstructured pruning literature.
-Specifically, we set $\rho_{i}$ to be a module-level analogue of any of the parameter-level unstructured pruning metrics like Wanda (Sun et al., 2023) or output activation magnitude.
-As an illustrative example, consider that we have shallow network with 1 hidden layer (of dimension $d$) we wish to prune.
-If output activation magnitude is the metric for our prior, we would calculate the activation vector for the $d$ hidden dimensions – $\hat{\mathbf{a}}$ – averaged across multiple samples.
-The likelihood that the $i$th column of the matrix $W\in R^{d\times d}$ is left unpruned would be:
-$\rho_{i}\propto\hat{\mathbf{a}}_{i}=\frac{1}{B}\sum_{b}\bigg{|}\sigma\bigg{(}\big{(}W^{T}[i,:]\big{)}x_{b}\bigg{)}\bigg{|}$
-$\sigma\text{ is the nonlinearity}$.
-Sampling sub-models according to $\mathbf{\rho}$ means we are more likely to explore models that already have a good performance $U$.
-The priors $\rho_{i}$ can be computed via forward passes through the original unmodified model $\mathbf{M}_{\theta}$ which allows us to respect memory constraints.
-To enhance the efficiency of our method, instead of considering all active modules for pruning, in each layer we consider the bottom $2p$ fraction of modules [^3] The choice of $2p$ is arbitrary.
-Practitioners can run hyper-parameter searches to find more optimal values for their settings.
-ranked according to our prior $\rho$.
-Thus, the top $1-2p$ fraction of modules remain fixed whilst the rest are perturbed and their relevance scores are computed and compared for pruning.
-This helps reduce the space of possible sub-models being considered for evaluation.
-For the bottom $2p$ fraction being compared, whenever we generate a mask $\alpha_{\mathbf{m}_{k}}$ with sparsity level $p$, we also generate its complement $\alpha^{c}_{\mathbf{m}_{k}}$, which is obtained by flipping the values in $\alpha_{\mathbf{m}_{k}}$ (except for the fixed $1-2p$ fraction of entries).
-Covert & Lee (2020) show that this technique can help reduce the variance of the estimator obtained from regression with binary inputs.
+An as yet unexplored design choice is selecting the $n$ candidate sub-models for evaluation.
+A naive approach–uniform sampling–is shown here to be suboptimal.
+Take $m_{i}$ as a module that is critical for good performance under evaluation with $U$.
+Since $n<N$, it means that some modules may never be ”turned on” in the list of $n$ chosen sub-models.
+If $m_{i}$ happens to be one of these masked-out modules, the resulting estimate of $\hat{\beta_{i}}=0$ would in turn result in $\bm{m}^{\mathrm{approx}}$ being a poor estimate for $\bm{m}^{\ast}$.
+Thus, a more informed selection is necessary for accurate and useful $\bm{\beta}$ estimates.
+Given a module $m_{i}$, we set the likelihood of it being present in any of the $n$ sampled sub-models proportional to a prior $\rho_{i}$ which reflects its usefulness.
+To define $\rho_{i}$, we can turn to metrics from the pruning literature.
+Specifically, we set $\rho_{i}$ to be a module-level analogue of any of the pruning metrics like Wanda (Sun et al., 2023) or activation magnitude.
+For example, in one-hidden-layer network of dimension $d$, if activation magnitude is the prior, we compute the activation vector over multiple samples.
+The probability of retaining the $i$th column of $W\in R^{d\times d}$ is: $\rho_{i}\propto\hat{\bm{a}}_{i}=\frac{1}{B}\sum_{b}\big{|}\sigma\big{(}\big{(}W^{T}[i,:]\big{)}x_{b}\big{)}\big{|}$ where $\sigma\text{ is the nonlinearity}$.
+Sampling sub-models based on $\bm{\rho}$ biases selection toward high-performing models.
+Since $\rho_{i}$ can be computed via forward passes through the unmodified model $\bm{M}_{\theta}$, this approach remains memory-efficient.
+Appendix H details the priors we explored.
+[…]
+2.4 Iterated Pruning
+Previous work on gradient-based pruning (Anwar et al., 2017; Frankle & Carbin, 2018) have shown that taking an iterated approach to pruning yields improved results over pruning directly to the target sparsity $p$.
+Similarly, we define an updated pruning fraction $p_{\mathrm{iter}}<p$ and perform $\mathrm{iter}=\lceil\frac{p}{p_{\mathrm{iter}}}\rceil$ steps where we explore $n_{\mathrm{iter}}=\lceil\frac{n}{\mathrm{iter}}\rceil$ sub-models at a time.
+At the beginning of each iteration, we re-estimate the priors $\bm{\rho}$ for the unpruned modules and use them to select the $n_{\mathrm{iter}}$ sub-models to evaluate.
+We combine the methods from Sections 2.2, 2.3, and 2.4 to develop Bonsai, a gradient-free structural pruning algorithm (Figure 2). Algorithm 1 provides the full details.
+3 Experimental Details and Main Results
+Table 1: Reported memory consumption of different methods.
+The minimum amount of memory required to run a Llama-7B model at half precision (FP16) is 14Gb.
+Running a forward pass with a batch size of 1 using the default model sequence length of 4096 uses around 20Gb of memory.
+Forward Only Gradient-Based
+Llama-2-7B (Model Only) Forward with bsz=1 (Bonsai Min.) Bonsai (Faster) LoRA Prune (Zhang et al., 2023) Compresso (Guo et al., 2023) LLM-Pruner (Ma et al., 2023) ShearedLlaMA (Xia et al., 2023)
+14GB 20GB A6000 (48GB) 1 A100 (80GB) 4 V100 (128GB) 2 A100 (160GB) 8 A100 (640GB)
+In all Bonsai experiments, we prune (1) self-attention heads and (2) fully connected layer dimensions, focusing on LLMs around 7B parameters.
+Since our goal is to aid memory-constrained practitioners, Table 1 compares memory requirements for pruning Llama-2-7B across methods.
+As can be seen, for models of our size range of interest, we can run Bonsai on any device with $\approx 20$Gb of memory if we restrict our batch size to 1.
+To reduce variance in score estimates, we average over 32 data points.
+A practitioner with 20GB of memory would run 32 forward passes with batch size 1, but this would be too slow for our experiments.
+Instead, we use a 48GB A6000 GPU (although not required by Bonsai), enabling batch sizes of 4–8 for faster experimentation.
+3.1 Bonsai is competitive with other forward pass-only, structured pruning methods
+Figure 3: Perplexity at 50% sparsity of LLaMA-{1,2}-7B on Wikitext-2 and C4 datasets.
+Dataset Method Sparsity Llama-1/seqlen=1024 Llama-2/seqlen=1024
+Wikitext-2 Base Model 0% 5.68 5.11
+FLAP 50% 17.27 14.49
+Bonsai 50% 15.72 12.38
+C4 Base Model 0% 7.34 7.04
+FLAP 50% 24.98 24.0
+Bonsai 50% 22.31 20.7
+We focus our first set of experiments on comparing structured pruning methods that can be run without gradient-based optimization.
+We prune the LLaMA-2 7B model (Touvron et al., 2023) to 50% sparsity and evaluate on the Wikitext-2 (Merity et al., 2016) validation dataset.
+Our module importance signal for pruning, $U$, is the language modeling performance on the training set.
+When measuring speedups, we consider end-to-end latency of running inference on model.sequence_length chunks of the Wikitext-2 validation set.
+See Table 10 (Appendix C.3) for details about the hyper-parameters used for these experiments.
+Figure 1 shows the results of some of our experiments.
+Bonsai expectedly outperforms the structured variant of Wanda (Sun et al., 2023), producing models with much lower perplexity at all sparsity levels and at any fixed desired speedup over the parent model.
+The more competitive baseline is FLAP (An et al., 2024), which was published a little over a month before the first release of our work.
+Even here (as shown in Table 3),  Bonsai outperforms FLAP while maintaining comparable speedups.
+A key advantage of  Bonsai is its flexibility—it can incorporate various pruning methods as informative priors (magnitude, Wanda, or fluctuation-based metrics like those used in FLAP–see Appendix H for more details on how we estimate priors).
+Once these priors are established, Bonsai’s regression step refines the importance estimates, potentially yielding better pruning decisions.
+Bonsai offers practitioners flexible runtime-quality tradeoffs.
+While the reported results use a configuration that runs in $\leq~{}4hours$ (which is considerably more than FLAP’s $\approx 1$hr’s runtime), practitioners can configure  Bonsai  to produced structurally pruned models in as little as 15 minutes by reducing the number of perturbations, sample size, or pruning iterations—at at some cost to performance.
+Conversely, allocating more compute time generally yields better pruning outcomes.
+This flexibility allow practitioners to tailor Bonsai to their specific constraints, unlike methods with fixed computational profiles.
+3.2 Introducing Post-Pruning Adaptation (PPA)
+After pruning the parent model, depending on the sparsity level $p$ achieved, it is possible to obtain a pruned model on which we can run full fine-tuning or a parameter-efficient fine-tuning method like LoRA (Hu et al., 2021) with the available hardware memory.
+In this case, we can fine-tune the pruned model on the downstream task in order to recover more performance relative to that of the parent model.
+Like many past works (Sanh et al., 2020; Xia et al., 2022b), we can combine pruning with distillation by incorporating a distillation loss in the training objective during fine-tuning of the pruned model.
+Let $\mathcal{L}_{\mathrm{task}}$ be the loss function over the task data and $\mathcal{L}_{\mathrm{distill}}$ be the distillation objective.
+We optimize the following post-pruning objective: $\mathcal{L}^{\mathrm{post-prune}}=\mathcal{L}_{\mathrm{task}}+\lambda\mathcal{L}_{\mathrm{distill}}$.
+Using $i$ to index the task data, we have:
+$\mathcal{L}_{\mathrm{distill}}=\sum_{i}D_{\mathrm{KL}}\big{(}\mathrm{logits}^{i}\left(\mathbf{M}_{|\mathbf{m}^{\mathrm{approx}}}\right)~{}\|~{}\mathrm{logits}^{i}\left({\mathbf{M}}\right)\big{)}$, where $\lambda$ is a scalar weighting that can cross validated.
+Note, distillation can be performed without significant memory overhead by apriori caching the logits from the parent model $\mathbf{M}$ instead of hosting the model in memory during fine-tuning.
+In the sections ahead, we will apply PPA after pruning the parent model if the child model is small enough to allow for this.
+Table 2: Wikitext-2 perplexity at 50% sparsity of LLaMA-2 7B with end-to-end latency speedups.
+[…]
+LlaMA-2 7B $5.11$ 1×1$\times$1 ×
+Phi-2 3B $8.69$ 1.24×1.24$\times$1.24 ×
+Wanda 2:4 3B $10.52$ 1.14×1.14$\times$1.14 ×
+  + PPA $8.34$ 0.75$\times$
+Bonsai 3B $19.47$ $\boldsymbol{1.58}\times$
+  + PPA $8.89$ $\boldsymbol{1.58}\times$
+3.2.1 Bonsai is competitive with semi-structured pruning methods
+We compare Bonsai to the semi-structured variant of Wanda (Sun et al., 2023). In general, structured pruning methods under-perform semi-structured pruning methods, but compensate for this in speedup.
+Before fine-tuning, the Wanda 2:4 model is more accurate but slower (1.14$\times$ vs 1.58$\times$) than the model from Bonsai.
+Since the Bonsai child model is small enough, we can perform PPA on it, resulting in improved accuracy (8.89ppl) with unchanged speedup.
+Finetuning the semi-sparse Wanda 2:4 model is unfortunately less straightforward.
+It would require similar memory resources as needed to finetune the parent model222though the child tensors are sparse, the resulting gradients and cached intermediate tensors can be dense and have the same dimensions as those of the parent model (say $M\times M$).
+Since Bonsai does structured pruning, the actual tensor sizes are shrunk (say $N\times N~{}|~{}N<M$) which reduces memory during backward passes.
+but we are in the setting where we do not have enough memory for this.
+We therefore have to use a parameter efficient fine-tuning method like LoRA (Hu et al., 2021) instead.
+While the performance gap can be bridged by LoRA finetuning (10.52 $\rightarrow$ 8.34), the adapted semi-structured model experiences a drastic slowdown (0.75$\times$), since the learned low-rank matrices cannot be merged with the original sparsified ones without reverting back to dense computation.
+Thus, LoRA fine-tuned Wanda 2:4 is twice slower (0.48$\sim$ 0.48$\times$ 0.48 ×) than the model from Bonsai and similarly accurate.
+In a memory-constrained setting, practitioners could opt for a pre-existing model of the target size instead of pruning a larger model.
+We compare the Bonsai-pruned model to Phi-2 (Li et al., 2023), a strong representative pre-existing model of similar size.
+As can be seen in Table 2, Bonsai is able to generate a model that is as accurate (0.2 difference in ppl) yet significantly faster (1.58$\times$ vs.
+1.24$\times$ speedup), thus making it a competitive option to consider even if a model already exists at the target size.
+3.2.2 Bonsai is competitive with gradient based structured pruning
+Table 3: LLaMA-1 (50% sparsity) after post-pruning adaptation. are results as reported by (Zhang et al., 2023).
+Method Wikitext-2 $\downarrow$ BoolQ HellaSwag WinoGrande ARC-e ARC-c Average $\uparrow$
+LLaMA1-7B (Touvron et al., 2023) 5.68 75.05 56.92 69.93 75.34 41.89 63.83
+LLM-Pruner (Ma et al., 2023) 16.41 60.28 47.06 53.43 45.96 29.18 47.18
+LoRAPrune (Zhang et al., 2023) 11.60 61.88 47.86 55.01 45.13 31.62 48.30
+Bonsai + PPA 10.92 67.22 43.09 61.64 54.92 26.28 50.63
+We compare Bonsai to the following gradient-based structured pruning approaches: LLM-Pruner (Ma et al., 2023) and LoRA-Prune Zhang et al.
+(2023).
+We use the reported results from (Zhang et al., 2023) since none of these methods are actually runnable in our memory-constrained setting (Table 1).
+We choose to compare to these over Sheared LlaMA (Xia et al., 2023) since they have much lower memory requirements (Table 1).
+We prune the LLaMA-1 7B model (Touvron et al., 2023) to 50% sparsity since these approaches report their results for the LLaMA-1 model only.
+We compare these methods on Wikitext-2 and also on six tasks from the Eleuther LLM Evaluation Harness benchmark (Gao et al., 2023).
+The pruning signal used for the Wikitext-2 task is the same as the above experiments.
+For the Eleuther Harness tasks, we use language modeling performance on the C4 (Raffel et al., 2020) dataset as pruning signal.
+We also perform parameter-efficient finetuning on our pruned model using 30K 512-length sequences from this corpus.
+Bonsai and LoRAPrune use similar amounts of the C4 dataset for Table 3 (30K vs 20K samples, respectively) whilst LLM-Pruner is trained on instruction tuned data with nearly twice the amount of unique samples (50K).
+Find more details in Appendix C.4.
+As evinced by Table 3, Bonsai outperforms gradient-based methods even though it exclusively uses forward passes in the pruning stage.
+We attribute the superior performance of Bonsai to the fact that its pruning decisions are informed by directly exploring the space of sub-models whilst the other approaches resort to inaccurate proxies of module relevance in order to reduce the memory overhead of a fully gradient-based optimization approach (though not by enough to be runnable in our setting).
+3.2.3 Bonsai can produce compressed models with strong zero-shot abilities
+Table 4: Phi-2 pruned to 35% sparsity and fine-tune the pruned model on small amount of the C4.
+We achieve strong performance compared to Phi-1.5 (trained from scratch).
+Since Sheared LlaMA has values absent, its MC Average would be misleading and we refrain from adding it.
+Generation Multiple Choice (MC)
+Model Size GSM8k ARC-c Winogrande Hellaswag Truthful-QA MMLU MC Average $\uparrow$
+(5-shot) (25-shot) (5-shot) (10-shot) (0-shot) (5-shot)
+Phi-2 (Li et al., 2023) 2.7B 54.81 61.09 74.35 75.11 44.47 58.11 62.63
+Phi-1.5 (Li et al., 2023) 1.5B 12.43 52.9 72.22 63.79 40.89 43.89 54.74
+Sheared LlaMA (Xia et al., 2023) 1.3B Not Reported 33.5 57.9 60.7 Not Reported 25.7 *
+Bonsai (w PPA) 1.8B 6.37 47.44 68.35 65.09 42.20 40.53 52.72
+  + Reasoning Tuning 1.8B 27.67 45.56 68.82 64.51 42.58 40.97 52.49
+Considerable amounts of compute and data, beyond what is feasible for lay ML practitioners, are needed to train LLMs with strong zero-shot capabilities (AI, 2023; Gemini Team et al., 2023).
+In this section, we demonstrate that Bonsai can empower everyday practitioners to produce strong and compact models with competitive zero-shot abilities by simply pruning bigger models on their available hardware.
+We use Bonsai to prune a $\approx$3B Phi-2 model to $\approx$1.8B (35% sparsity).
+Bonsai hyper-parameters in this experiment are in Appendix C.5.
+Since it is small, the 1.8B pruned model can be fully fine-tuned on 1 A6000 GPU over 100k sequences of 2,048 tokens from the C4 dataset.
+As can be seen from Table 4, our pruned model achieves strong zero-shot performance on the Hugging Face OpenLLM leaderboard (Gao et al., 2023) compared to Phi-1.5, a smaller version in the Phi series of models that was trained from scratch.
+Figure 4: LLaMA-2 7B @ 50% sparsity (No PPA). Pertubation and regression components are both needed to make Bonsai effective. Experiment details in Appendix D.
+Interestingly, one exception to the general trend of Bonsai’s strong performance is the GSM-8K dataset, which is a mathematical reasoning dataset that requires generation of long reasoning chains.
+In our experiments, Bonsai prunes with respect to language modeling likelihood, as opposed to reasoning accuracy, and we posit that this mismatch may have contributed to our model’s underperformance on GSM8K.
+We attempt to remedy the drop in reasoning ability by including the GSM8K training data during post-pruning finetuning (resulting in a total of 108K fine-tuning samples).
+This boosts our model’s performance on the GSM8K with almost no degradation of performance on the other tasks.
+4 Analysis
+We conduct ablative experiments to understand the impact of the ingredients from Section 2.
+Do we need both perturbative and regressive components of Bonsai?
+Figure 4 shows that both components are key to obtaining a good pruned model.
+Removing the estimation of module importances via regression leads to a degradation in performance ($61.6$ ppl $\rightarrow$ $146.6$ ppl).
+Further degradation ($146.6$ ppl $\rightarrow$ $405.7$ ppl) is encountered if we do not perform perturbative evaluations on the parent model but simply prune according to the prior $\rho$ as computed from the unperturbed parent model.
+What is a reasonable number of perturbative samples?
+We investigate the number of perturbative samples required to obtain good regression estimates of module importance based on Equation 3.
+Our results are shown in Table 5.
+As expected, performance improves as we increase the number of sub-models explored.
+We note that the number of samples being explored, $\mathrm{ns}$, is significantly less than the number of candidate modules at each iteration ($N\approx 70k$).
+Nevertheless, Bonsai is able to deliver a performant pruned model because of the recipes developed in Section 2.
+$\mathrm{ns}=50$ $\mathrm{ns}=200$ $\mathrm{ns}=1000$
+PPL ($\downarrow$) NaN $61.63$ $22.09$
+Table 5: Wikitext-2 perplexity of LLaMA-2 7B @ 50% sparsity (No PPA). We vary the number of perturbative evaluations. Details in Appendix I.
+How much performance is recovered by post-pruning adaptation?
+During iterative pruning, Bonsai damages the parent model by removing modules but does not perform intermittent retraining to recover lost performance since even intermediate models may be too large for fine-tuning.
+Even so, as Table 6 shows, the final model produced by Bonsai has reasonable performance without fine-tuning.
+We attribute this to the redundancy of modules with respect to the target downstream tasks and Bonsai’s ability to identify good candidates for pruning.
+If the pruned model is small enough in size, we may perform post pruning adaptation to recover more performance, as can be seen from Table 6.
+Table 6: Impact of PPA on LLaMA-2 7B @ 50% sparsity. Details in Appendix G.
+Method Wikitext-2 PPL
+No Post-Pruning Adaptation $19.47$
+Post-Pruning Finetuning $10.39$
+   +  Distillation $8.89$
+Figure 5: LLaMA-2 7B pruned to 50% sparsity. See Appendix H for experiment details and definitions of Wanda and Activation Magnitude priors.
+What is the impact of the choice of metric for the prior $\rho$?
+We investigate three different choices of metrics for defining $\rho$.
+Figure 5 shows that using the module-level analogue of Wanda Sun et al.
+(2023) yields the best performance, both before and after post-pruning adaptation.
+This indicates that Wanda is a strong signal for efficiently estimating the importance of model units.
+5 Related Work
+Here we discuss relevant work in LLM pruning, and defer a longer discussion compression methods to Appendix B.
+Unstructured pruning:
+Whilst structured pruning approaches remove entire model components like layers (Xu et al., 2020; Xia et al., 2022b), dimensions of linear layers (Wang et al., 2019) or attention heads (Michel et al., 2019; Held & Yang, 2022), unstructured pruning (Han et al., 2015; Frankle & Carbin, 2018; Benbaki et al., 2023; Sun et al., 2023) removes individual parameters of the model.
+These approaches achieve memory savings by inducing sparsity in the model weights, but they generally do not result in actual model speedups except when specialized hardware is available (Mishra et al., 2021).
+Proposed semi-structured sparsity methods (Mishra et al., 2021) such as 2:4 and 4:8 patterns do result in faster inference, but the speedup gains they achieve are far from the idealized 2×2$\times$2 ×.
+Structured pruning without backward passes:
+To the best of our knowledge, all existing structured pruning techniques for large (over 1B scale) language models, like LoRAPrune (Zhang et al., 2023) and LLM-Pruner (Ma et al., 2023), are gradient-based.
+Thus, under the memory setting we consider, these methods cannot be applied since their memory requirements well exceed the memory needed for inference on the model to be pruned.
+For smaller language models like BERT, (Nova et al., 2023) recently proposed Kernelized Convex Masking (KCM) for gradient-free structured pruning.
+Unfortunately, to prune a fully connected layer with $K$ intermediate units, KCM requires instantiating a $K\times K$ coefficient matrix (for each layer).
+Whilst this is feasible for BERT models, a typical LLM (AI, 2023; Touvron et al., 2023) has $K\sim 10^{4}-10^{5}$ which would make the size of each per-layer coefficient matrix comparable to the size of the LLM itself.
+In computer vision, there exist several perturbative gradient-free structured pruning techniques (Ancona et al., 2020; Dekhovich et al., 2021).
+However, these methods have been exclusively applied to small vision models (VGG; Simonyan & Zisserman (2014), WideResnet; Zagoruyko & Komodakis (2016)) and, taken as is, would not scale to LLMs.
+We introduce several techniques to make Bonsai scale to large language models.
+A key difference between Bonsai and these perturbative methods is that pruning decisions are made globally instead of on a local layer-by-layer basis (Ancona et al., 2020; Dekhovich et al., 2021).
+A failure mode of layer-by-layer pruning is that if a module, say $m_{i}$, which is in the top $p$ fraction of relevant modules when the whole model is considered but in the bottom $1-p$ fraction within its layer will not be included in the pruned model — a clearly sub-optimal choice.
+Bonsai enables us to obtain globally meaningful estimates of module relevance that are comparable across layers.
+6 Conclusion, Limitations, and Future Work
+In this work, we have presented Bonsai, a gradient-free structured pruning method that enables efficient compression of LLMs using only forward passes.
+By eliminating the need for backward passes, Bonsai reduces memory requirements by over 2×2$\times$2 ×, allowing practitioners to prune and adapt models on hardware configurations where gradient-based approaches would be infeasible.
+Through extensive experiments, we have demonstrated that Bonsai produces models that are small, fast, and accurate—-outperforming even most of these gradient-based methods while using substantially fewer resources.
+A primary limitation of Bonsai is its computational runtime.
+As shown in Section 4, performance improves with more sub-model exploration, slower pruning rates, and larger data samples, but these improvements come at the cost of increased runtime.
+While competitors like FLAP complete pruning in $\approx 1$ hour, our configuration of Bonsai that achieves optimal results requires $\leq 4$ hours.
+However, Bonsai offers practitioners the flexibility to configure faster pruning schedules (as quick as 15 minutes) when speed is prioritized over achieving optimal compression quality.
+This runtime-quality tradeoff is a design feature that allows adaptation to various computational constraints.
+Additionally, since the models produced by Bonsai can guarantee optimal performance and speedups, this one-time investment becomes worthwhile when amortized over large-scale deployment.
+Bonsai presents several promising avenues for future research.
+First, while our current approach samples sub-models from an informative prior, the sampling process is not completely adaptive.
+Bonsai could be further enhanced by dynamically exploring the space of sub-models based on previous evaluations.
+Additionally, due to our focus on forward-pass-only operations, Bonsai does not fine-tune the model during iterative pruning.
+Integration with gradient-free approaches like MeZO (Malladi et al., 2023) could allow for continuous updates during the pruning process, potentially leading to better preservation of model performance.
+Appendix A Broader Impact Statement
+Network pruning is an effective compression strategy to reduce the inference costs of large language models.
+However, existing techniques for structured pruning impose expensive memory constraints that may render them out of reach in practice.
+By reducing such constraints, work aims to democratize pruning to provide practitioners the ability to produce their own pruned LLMs for real-world applications.
+While we ultimately view this democratization as a benefit, we note that a necessary outcome is that it may also make it more feasible for malicious actors to readily use LLMs in practice.
+Finally, we note that our work has primarily focused on the axes of efficiency and utility (e.g., perplexity) in assessing performance; recent work has shown that compression may also have outsized effects on issues such as model fairness and robustness, which would be interesting additional aspects to consider in future study.
+Appendix B Background: Additional Methods for LLM Compression
+Although we focus on pruning in this work, we note that prior research has explored various other compression schemes such as distillation (Hinton et al., 2015; Gu et al., 2023) and quantization (Xiao et al., 2023) to create smaller models from larger pre-trained ones.
+Similar to structured pruning, these compression methods themselves often impose significant computational burdens.
+For example, distillation-based techniques require using LLMs to generate large amounts of teacher data (Jiao et al., 2019; Hsieh et al., 2023).
+Although unstructured pruning (Frantar & Alistarh, 2023; Sun et al., 2023) and quantization have lower training-time resource demands, the models they produce either require specialized hardware to achieve speedups (Mishra et al., 2021) or may actually slow down inference due to additional computational overhead (Dettmers et al., 2022).
+Table 7 summarizes these existing limitations across different compression approaches.
+We also cover additional background in distillation and quantization below.
+Table 7: Landscape of resource consumption (memory and compute) of different model compression methods at training time and the inference time resource consumption of the models they deliver.
+✗  $\rightarrow$ a prohibitive cost to the lay practitioner whilst ✓  $\rightarrow$ a viable option with respect to that resource.
+Regime Resource Approaches
+Quantization (Mixed Precision) Distillation Unstructured Pruning Gradient-Based Structured Pruning Bonsai (Ours)
+Train Memory
+Compute
+Inference Memory
+Compute
+Distillation:
+Originally proposed by (Hinton et al., 2015; Sanh et al., 2019) have shown that it is possible to reach similar performances on many downstream tasks using much smaller language models pre-trained with knowledge distillation.
+The resulting models are lighter and faster at inference time.
+Unfortunately, as noted by (Xia et al., 2022b), distilling a pre-trained model to a student model that has been initialized from scratch requires lots of compute and data to recover reasonable levels of teacher performance.
+To reduce this resource burden, (Sanh et al., 2020; Lagunas et al., 2021; Xia et al., 2022b) combine distillation with gradient-based structured pruning, so that the student model does not have to be initialized from scratch.
+Bonsai also leverages distillation, but only during the post-pruning adaptation phase.
+This maintains the constraint that the pruning process is gradient-free.
+And since the pruned model is smaller than the parent model, we can safely perform distillation and parameter efficient fine-tuning (Hu et al., 2021; He et al., 2021) within the bounds of available memory.
+Quantization:
+(Li et al., 2020) introduced a training strategy that showed that heavily compressed, large models achieved higher accuracy than lightly compressed, small models, providing both efficiency and high-accuracy results on NLP tasks at once.
+(Dettmers et al., 2022) introduced a new method called ‘LLM.int8()‘, which allows loading large models with 16 or 32-bit weights and using them immediately for inference without any performance degradation by combining vector-wise quantization with mixed-precision decomposition.
+(Xiao et al., 2023) enabled 8-bit weight, 8-bit activation (W8A8) quantization for LLMs with >100B parameters, which smoothed activation outliers by migrating the quantization difficulty from activations to weights with a mathematically equivalent transformation.
+Quantization tends to be complementary to pruning, allowing independent exploration and subsequent combinations of both techniques.
+This paper focuses on gradient-free structured pruning.
+We leave combining Bonsai with quantization for future work.
+Appendix C Main Experiment Details
+C.1 Full Algorithm
 Algorithm 1 Bonsai
@@
 4:     $\bar{\mathbb{\rho}}^{l}\leftarrow$ Fix the top $(1-2p_{\mathrm{iter}})$ of $\mathbb{\rho}^{l}$ to $\infty$
-5:     Sample $\{\mathbf{\bar{m}}_{i}\}_{[\frac{n_{\mathrm{iter}}}{2}]}$ sub-models according to $\bar{\mathbb{\rho}}^{l}$
-6:     Generate $\{\mathbf{\bar{m}}^{c}_{i}\}_{[\frac{n_{\mathrm{iter}}}{2}]}$ complements of $\{\mathbf{\bar{m}}_{i}\}_{[\frac{n_{\mathrm{iter}}}{2}]}$
-7:     Run forward pass on each sub-model and compute $U$.
-Construct $\mathbb{D}^{l}=\left\{\bar{\mathbf{m}}_{i},U_{i}\right\}_{[\frac{n_{\mathrm{iter}}}{2}]}\bigcup\left\{\bar{\mathbf{m}}^{c}_{k},U_{k}\right\}_{[\frac{n_{\mathrm{iter}}}{2}]}$
-8:     $\beta^{\l}\leftarrow\mathrm{Regress}\left(\mathbb{D}^{l}\right)$
-9:     $\{m_{\mathrm{pruned}}\}\leftarrow$ sort $\beta^{\l}$ and remove the bottom $k$ modules that correspond to $p_{\mathrm{iter}}$ fraction of the model.
-10:     $\mathbf{m}\leftarrow$ update module list to exclude $\{m_{\mathrm{pruned}}\}$
-11:  end for
-12:  Output: Pruned model $\mathbf{M}_{|\mathbf{m}}$
-3.3 Iterated Pruning
-Previous work on gradient-based pruning (Anwar et al., 2017; Frankle & Carbin, 2018) have shown that taking an iterated approach to pruning yields improved results over pruning directly to the target sparsity $p$.
-Similarly, we define an updated pruning fraction $p_{\mathrm{iter}}<p$ and perform $\mathrm{iter}=\lceil\frac{p}{p_{\mathrm{iter}}}\rceil$ steps where we explore $n_{\mathrm{iter}}=\lceil\frac{n}{\mathrm{iter}}\rceil$ sub-models at a time.
-At the beginning of each iteration, we re-estimate the priors $\mathbf{\rho}$ for the unpruned modules and use them to select the $n_{\mathrm{iter}}$ sub-models to evaluate.
-We combine the recipes developed in Sections 3.1, 3.2 and 3.3 together to produce Bonsai [^4] Structural pruning is a canonical way of giving a bonsai tree its shape hence the name., our gradient-free structural pruning algorithm.
-Algorithm 1 specifies Bonsai in detail.
-4 Post-pruning adaptation
-Depending on the sparsity level $p$, it is possible to obtain a pruned model on which it is feasible to run a parameter-efficient finetuning method like LoRA (Hu et al., 2021) with the available hardware memory.
-In this case, we can fine-tune the pruned model (result of Algorithm 1) on the downstream task in order to recover some of the parent model performance that was degraded by pruning.
-Like many past works (Sanh et al., 2020; Xia et al., 2022), we combine pruning with distillation by incorporating a distillation loss in the training objective during fine-tuning of the pruned model.
-Let $\mathcal{L}_{\mathrm{task}}$ be the loss function over the task data and $\mathcal{L}_{\mathrm{distill}}$ be the distillation objective.
-We optimize the following post-pruning objective: $\mathcal{L}^{\mathrm{post-prune}}=\mathcal{L}_{\mathrm{task}}+\lambda\mathcal{L}_{\mathrm{distill}}$.
-Using $i$ to index the task data, we have:
-$\mathcal{L}_{\mathrm{distill}}=\sum_{i}D_{\mathrm{KL}}\bigg{(}\mathrm{logits}^{i}\left(\mathbf{M}_{|\mathbf{m}^{\mathrm{approx}}}\right)~{}\|~{}\mathrm{logits}^{i}\left({\mathbf{M}}\right)\bigg{)}$, where $\lambda$ is a scalar weighting that can be optimized via cross-validation.
-Also, distillation can be performed without significant memory overhead by apriori caching the logits from the parent model $\mathbf{M}$ instead of hosting the model in memory during fine-tuning.
-Table 2: Wikitext-2 perplexity for 50% sparsity of LLaMA-2 7B with end-to-end latency speedups (relative to LLaMA-2 7B).
-We use Phi-2 (Li et al., 2023) at the target size as a strong model for comparison.
-For semi-structured (Wanda 2:4) pruning, after fine-tuning, the learned LoRA weights cannot be merged with the primary model weights else the model reverts back to being dense; this leads to the reported slow downs.
-All methods use forward passes only on the parent model.
-[…]
-Phi-2 3B $8.69$ 1.24×1.24$\times$1.24 ×
-LLaMA-2 7B Pruned
-Wanda 2:4 3B $10.52$ 1.14×1.14$\times$1.14 ×
-$8.34$ 0.75$\times$
-Bonsai 3B $8.89$ $\boldsymbol{1.58}\times$
-5 Experimental Details
-All experiments are conducted on a single NVIDIA A6000 (48Gb) GPU. In all experiments with Bonsai, we prune (1) the heads in the self-attention layers (2) the dimensions of the fully connected layers.
-Comparing Forward Pass Only Methods: We focus our first set of experiments on comparing methods that can be run without gradient-based optimization.
-We consider pruning the LLaMA-2 7B model (Touvron et al., 2023) to 50% sparsity.
-We evaluate on the Wikitext-2 (Merity et al., 2016) validation dataset and so our signal for pruning, $U$, is the language modelling performance on the training set.
-When measuring speedups, we consider end-to-end latency of running inference on model.sequence_length chunks of the Wikitext-2 validation set.
-See Appendix A.2 for details about the hyper-parameters used for all methods in this experiment and for specifics about fine-tuning.
-Comparing Structured Pruning Approaches: We compare Bonsai to the following gradient-based structured pruning approaches: LLM-Pruner (Ma et al., 2023) and LoRA-Prune (Zhang et al., 2023).
-We prune the LLaMA-1 7B model (Touvron et al., 2023) to 50% sparsity since this was the model version available at time of release of the above methods.
-We compare these methods on Wikitext-2 and also on 6 tasks from the Eleuther LLM Evaluation Harness (Gao et al., 2023).
-Pruning signal for Wikitext-2 task is the same as the above experiment.
-For the Eleuther Harness tasks, we use language modelling performance on the C4 (Raffel et al., 2020) dataset as pruning signal.
-We also do parameter efficient finetuning on our pruned model with 30K 512-length sequences from this corpus.
-Specific hyper-parameters for Bonsai for this experiment can be found in Appendix A.3.
-6 Main Results
-6.1 Bonsai produces fast and performant models with only forward passes
-In Table 2, we explore options that are available to practitioners when they can only run forward passes on the parent model. Here, LLaMA-2 7B is compressed to 50% sparsity.
-We compare Bonsai to the semi-structured variant of Wanda (Sun et al., 2023).
-Before fine-tuning, the model produced by Wanda 2:4 achieves a speedup over the parent model (1.14$\times$) but is slower and less accurate than Bonsai’s pruned model (8.89ppl vs 10.52ppl).
-While the performance gap can be bridged by LoRA finetuning (10.52 $\rightarrow$ 8.34), the adapted semi-structured model experiences a drastic slowdown (0.75$\times$), since the learned low-rank matrices cannot be merged with the original sparsified ones without reverting back to dense computation.
-Thus, LoRA fine-tuned Wanda 2:4 is twice slower (0.48$\sim$ 0.48$\times$ 0.48 ×) than the model from Bonsai.
-In this memory-constrained setting, practitioners could alternatively opt for a pre-existing model of the target size instead of pruning a larger model.
-We compare the Bonsai-pruned model to Phi-2 (Li et al., 2023), a strong representative pre-existing model of similar size.
-As can be seen in Table 2, Bonsai is able to generate a model that is as accurate (0.2 difference in ppl) yet significantly faster (1.58$\times$ vs.
-1.24$\times$ speedup), thus making it a competitive option to consider even if a model already exists at the target size.
-Table 3: LLaMA-1 (50% sparsity) after post-pruning adaptation with LoRA. ${}^{\dagger}$ indicate results as reported by Zhang et al. (2023).
-Method Foward-only Wikitext-2 $\downarrow$ BoolQ HellaSwag WinoGrande ARC-e ARC-c Average $\uparrow$
-LLaMA1-7B (Touvron et al., 2023) - 5.68 75.05 56.92 69.93 75.34 41.89 63.83
-${}^{\dagger}$ (Ma et al., 2023) 16.41 60.28 47.06 53.43 45.96 29.18 47.18
-${}^{\dagger}$ (Zhang et al., 2023) 11.60 61.88 47.86 55.01 45.13 31.62 48.30
-Bonsai 10.92 67.22 43.09 61.64 54.92 26.28 50.63
-6.2 Bonsai outperforms other structured pruning approaches
-We next investigate the setting where the practitioner has enough memory to run gradient-based structured pruning methods.
-We compare Bonsai to recent SoTA methods: LLM-Pruner (Ma et al., 2023) and LoRA-Prune (Zhang et al., 2023).
-Since these approaches report their results for the LLaMA-1 only, we prune LLaMA-1 7B to 50% sparsity.
-As evinced by Table 3, Bonsai outperforms existing structured pruning methods for LLMs even though it exclusively uses forward passes in the pruning stage.
-We attribute the superior performance of Bonsai to the fact that its pruning decisions are informed by directly exploring the space of sub-models whilst the other approaches resort on inaccurate proxies of module relevance in order to reduce the memory overhead of a fully gradient-based optimization approach.
-Specifically, LoRA-Prune’s criterion for deciding which modules to prune uses gradient signals not from the modules directly but from a surrogate low-rank matrix.
-On the other hand, LLM-Pruner attempts to model the impact of removing a module on the model loss via a Taylor expansion and substitutes the LLM’s intractable (memory-wise) Hessian term with the Fisher information matrix which we posit leads to sub-optimal pruning decisions.
-6.3 Bonsai can produce compressed models with strong zero-shot abilities
-Table 4: Phi-2 pruned to 35% sparsity compared to the best sub-2B parameter models on the Hugging Face OpenLLM Leaderboard.
-At the time of submission, $\dagger$ was the best sub-2B parameter model on the Leaderboard.
-Generation Multiple Choice (MC)
-Model Size GSM8k (5-shot) ARC-c (25-shot) Winogrande (5-shot) Hellaswag (10-shot) Truthful-QA (0-shot) MMLU (5-shot) MC Average $\uparrow$
-Phi-2 (Li et al., 2023) 2.7B 54.81 61.09 74.35 75.11 44.47 58.11 62.63
-StableLM-2-1_6$\dagger$ 1.6B $\boldsymbol{17.44}$ 43.34 64.56 $\boldsymbol{70.45}$ 36.78 38.95 50.82
-Ours 1.8B 6.37 $\boldsymbol{47.44}$ $\boldsymbol{68.35}$ 65.09 $\boldsymbol{42.20}$ $\boldsymbol{40.53}$ $\boldsymbol{52.72}$
-Considerable amounts of compute and data, beyond what is available to lay practitioners, are needed to train LLMs with strong zero-shot capabilities (OpenAI et al., 2023; Gemini Team et al., 2023).
-In this section, we demonstrate that Bonsai can empower everyday practitioners to produce strong and compact models with competitive zero-shot abilities by simply pruning bigger models on their available hardware.
-We use Bonsai to prune a $\approx$3B Phi-2 model to $\approx$1.8B (35% sparsity).
-Values for the Bonsai hyper-parameters in this experiment are in Appendix A.4.
-Since its relatively small, the 1.8B pruned model can be fully fine-tuned on 1 A6000 GPU over 100k sequences of 2,048 tokens from the C4 dataset.
-As can be seen from Table 4, our pruned model achieves strong zero-shot performance compared to StableLM, the leading sub-2B parameter LLM on the Hugging Face OpenLLM leaderboard (Gao et al., 2023) — outperforming it on 4 out of 6 tasks as of January 2024.
-Interestingly, one exception to the general trend of Bonsai’s superior performance is the GSM-8K dataset, which is a mathematical reasoning dataset that requires generation of a long reasoning chains.
-We posit that this is because currently, Bonsai prunes with respect to language modeling likelihood, as opposed to reasoning accuracy.
-An interesting avenue for future work is to prune to improve maintain reasoning ability.
-7 Analysis
-In this section, we conduct various ablative experiments to understand how the methodological ingredients from Section 3 contribute to make Bonsai effective.
-Do we need both perturbative and regressive components of Bonsai?
-Figure 2 shows that both components are key to obtaining a good pruned model.
-Removing the estimation of module importances via regression leads to a degradation in performance ($61.6$ ppl $\rightarrow$ $146.6$ ppl).
-More degradation ($146.6$ ppl $\rightarrow$ $405.7$ ppl) is encountered if we do not perform any perturbative evaluations on the parent model but simply just prune according to the prior $\rho$ as computed from the unperturbed parent model.
-Figure 2: LLaMA-2 7B pruned to 50% sparsity.
-Perturbatively evaluating the impact of removing modules and then using regression to estimate module importances are both key in making Bonsai effective.
-Details of experiment configuration can be found in Appendix B.
-No post-pruning adaptation is performed.
-Table 5: Varying the number of perturbative evaluations. Wikitext-2 perplexity of LLaMA-2 7B pruned to 50% sparsity. See Appendix G for details. No post-pruning adaptation is performed.
-$\mathrm{ns}=50$ $\mathrm{ns}=200$ $\mathrm{ns}=1000$
-PPL ($\downarrow$) NaN $61.63$ $22.09$
-How many perturbative samples are sufficient?
-We investigate the number of perturbative samples required to obtain good estimates of module importances after performing regression as in Equation 3.
-Our results are shown in Table 5.
-As expected, performance improves as we increase the number of sub-models explored.
-We note that the number of samples we explore, $\mathrm{ns}$, is significantly less than the number of candidate modules at each iteration ($N\approx 70k$), nevertheless Bonsai is able to deliver a performant pruned model because of the recipes developed in Section 3.
-How much performance is recovered by post-pruning adaptation?
-During iterative pruning, Bonsai damages the parent model by removing modules but does not perform intermittent retraining to recover lost performance since even intermediate models may be too large for fine-tuning.
-Even so, as Table 6 shows, the final model produced by Bonsai has reasonable performance without fine-tuning.
-We attribute this to the general robustness of LLMs, the redundancy of modules with respect to target end-tasks and Bonsai’s ability to identify good candidates for pruning.
-If the pruned model is small enough in size, we can perform either full fine-tuning or parameter-efficient fine-tunin to recover more performance, as can be seen from Table 6.
-Table 6: Impact of post-pruning adaptation of LLaMA-2 7B pruned to 50% sparsity. See Appendix E for more details.
-Method Wikitext-2 PPL
-No Post-Pruning Adaptation $19.47$
-Post-Pruning Finetuning $10.39$
-   +  Distillation $8.89$
-Figure 3: LLaMA-2 7B pruned to 50% sparsity. See Appendix F for details of experiment configuration and definitions of module-level analogues of Wanda and Activation Magnitude.
-What is the impact of the choice of metric for the prior $\rho$?
-We investigate three different choices of metrics for defining $\rho$.
-Figure 3 shows that using the module-level analogue of Wanda (Sun et al., 2023) yields the best performance, both before and after post-pruning adaptation.
-This indicates that Wanda is a strong signal for efficiently estimating the importance of model units (whether at the parameter- or module-level).
-Should Bonsai prune iteratively?
-Table 7 demonstrates the benefits of using Bonsai  in an iterative fashion.
-Pruning slowly ($p_{\mathrm{iter}}=0.05$) yields the best results, but this comes at the cost of increasing the total time to prune the model.
-The performance gap between values of $p_{\mathrm{iter}}$ persists even after post-pruning adaptation, indicating that slower pruning allows for more accurate estimates of module importance.
-Table 7: Varying $p_{\mathrm{iter}}$. Wikitext-2 perplexity of LLaMA-2 7B pruned to 50% sparsity. See Appendix C for experiment details.
-$p_{\mathrm{iter}}=0.05$ $p_{\mathrm{iter}}=0.1$ $p_{\mathrm{iter}}=0.2$
-w/o Adapt $19.47$ $61.63$ $209.44$
-w Adapt $8.89$ $9.15$ $9.57$
-8 Related Work
-8.1 Pruning
-Unstructured pruning:
-Whilst structured pruning approaches remove entire model components like layers (Xu et al., 2020; Xia et al., 2022), dimensions of linear layers (Wang et al., 2019) or attention heads (Michel et al., 2019; Held & Yang, 2022), unstructured pruning (Han et al., 2015; Frankle & Carbin, 2018; Benbaki et al., 2023; Sun et al., 2023) removes individual parameters of the model.
-These approaches achieve memory savings by inducing sparsity in the model weights, but they generally do not result in actual model speedups except when specialized hardware is available (Mishra et al., 2021).
-Proposed semi-structured sparsity methods (Mishra et al., 2021) such as 2:4 and 4:8 patterns do result in faster inference, but the speedup gains they achieve are far from the idealized 2×2$\times$2 ×.
-There are several gradient-free, unstructured pruning approaches.
-Since, as far as we know, there are no scalable gradient-free structured pruning alternatives outside of Bonsai, we compare it with Wanda (Sun et al., 2023), a SoTA unstructured/semi-structured pruning technique for LLMs.
-Structured pruning without backward passes:
-To the best of our knowledge, all existing structured pruning techniques for large (over 1B scale) language models, like LoRAPrune (Zhang et al., 2023) and LLM-Pruner (Ma et al., 2023), are gradient-based.
-Thus, under the memory setting we consider, these methods cannot be applied since their memory requirements well exceed the memory needed for inference on the model to be pruned.
-For smaller language models like BERT, Nova et al.
-(2023) recently proposed Kernelized Convex Masking (KCM) for gradient-free structured pruning.
-Unfortunately, to prune a fully connected layer with $K$ intermediate units, KCM requires instantiating a $K\times K$ coefficient matrix (for each layer).
-Whilst this is feasible for BERT models, a typical LLM (OpenAI et al., 2023; Touvron et al., 2023) has $K\sim 10^{4}-10^{5}$ which would make the size of each per-layer coefficient matrix comparable to the size of the LLM itself.
-In computer vision, there exist several perturbative gradient-free structured pruning techniques (Ancona et al., 2020; Dekhovich et al., 2021).
-However, these methods have been exclusively applied to small vision models (VGG; (Simonyan & Zisserman, 2014), WideResnet; (Zagoruyko & Komodakis, 2016)) and, taken as is, would not scale to LLMs.
-We introduce several techniques to make Bonsai scale to large language models.
-A key difference between Bonsai and these perturbative methods is that pruning decisions are made globally instead of on a local layer-by-layer basis (Ancona et al., 2020; Dekhovich et al., 2021).
-A failure mode of layer-by-layer pruning is that if a module, say $m_{i}$, which is in the top $p$ fraction of relevant modules when the whole model is considered but in the bottom $1-p$ fraction within its layer will not be included in the pruned model — a clearly sub-optimal choice.
-Bonsai enables us to obtain globally meaningful estimates of module relevance that are comparable across layers.
-8.2 Model compression beyond pruning
-Distillation:
-Originally proposed by Hinton et al.
-(2015), Sanh et al.
-(2019) have shown that it is possible to reach similar performances on many downstream tasks using much smaller language models pre-trained with knowledge distillation.
-The resulting models are lighter and faster at inference time.
-Unfortunately, as noted by Xia et al.
-(2022), distilling a pre-trained model to a student model that has been initialized from scratch requires lots of compute and data to recover reasonable levels of teacher performance.
-To reduce this resource burden, Sanh et al.
-(2020); Lagunas et al.
-(2021); Xia et al.
-(2022) combine distillation with gradient-based structured pruning, so that the student model does not have to be initialized from scratch.
-Bonsai also leverages distillation, but only during the post-pruning adaptation phase.
-This maintains the constraint that the pruning process is gradient-free.
-And since the pruned model is smaller than the parent model, we can safely perform distillation and parameter efficient fine-tuning (Hu et al., 2021; He et al., 2021) within the bounds of available memory.
-Quantization:
-Li et al.
-(2020) introduced a training strategy that showed that heavily compressed, large models achieved higher accuracy than lightly compressed, small models, providing both efficiency and high-accuracy results on NLP tasks at once.
-Dettmers et al.
-(2022) introduced a new method called ‘LLM.int8()‘, which allows loading large models with 16 or 32-bit weights and using them immediately for inference without any performance degradation by combining vector-wise quantization with mixed-precision decomposition.
-Xiao et al.
-(2023) enabled 8-bit weight, 8-bit activation (W8A8) quantization for LLMs with >100B parameters, which smoothed activation outliers by migrating the quantization difficulty from activations to weights with a mathematically equivalent transformation.
-Quantization tends to be complementary to pruning, allowing independent exploration and subsequent combinations of both techniques.
-Some quantization approaches, specifically mixed precision ones like Dettmers et al.
-(2022) (Appendix D), can produce resulting models that are slower than the parent model, though it has also been shown that it can result in more accurate models than pruning (Kuzmin et al., 2023).
-This paper focuses on gradient-free structured pruning.
-We leave combining Bonsai with quantization for future work.
-9 Conclusion, Limitations, and Future Work
-In this work, we have presented Bonsai, the first tractable gradient-free method for structured pruning of LLMs.
-Bonsai allows practitioners to prune any LLM as long as they have enough memory to run inferences on the model.
-Through a battery of experiments, we have shown that the models produced by Bonsai are small, fast, and accurate.
-A primary limitation of Bonsai is its runtime.
-As Section 7 shows, performance improves when we increase the number of sub-models explored, slow down pruning, and use more data samples; but this comes at the cost of increased runtime.
-Our result in Table 2 took $\sim 40$ hours to obtain, whilst Wanda-2:4 takes on the order of minutes.
-Since the model we produce is twice as fast, this trade-off is worthwhile when amortized over inference on the pruned models.
-Bonsai presents several avenues for expansion by future work.
-First, though sub-models are sampled from an informative prior, $\rho$, the sampling process is not adaptive.
-Bonsai could further be strengthened by dynamically exploring the space of sub-models.
-Next, because of memory constraints, Bonsai does not fine-tune the model during iterative pruning to recover degraded performance.
-However, forward-pass-only fine-tuning approaches like MeZO (Malladi et al., 2023) exist that can be used to dynamically update the model during pruning.
-10 Acknowlegements
-This work was supported in part by the National Science Foundation grants IIS1705121, IIS1838017, IIS2046613, IIS2112471, and funding from Meta, Morgan Stanley, Amazon, and Google.
-Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of any of these funding agencies.
-We are grateful for helpful feedback from Mingjie Sun, Victor Akinwande, Asher Trockman, Afshin Rostamizadeh and Daniel Glasner
-11 Broader Impact Statement
-Network pruning is an effective compression strategy to reduce the inference costs of large language models.
-However, existing techniques for structured pruning impose expensive memory constraints that may render them out of reach in practice.
-By reducing such constraints, work aims to democratize pruning to provide practitioners the ability to produce their own pruned LLMs for real-world applications.
-While we ultimately view this democratization as a benefit, we note that a necessary outcome is that it may also make it more feasible for malicious actors to readily use LLMs in practice.
-Finally, we note that our work has primarily focused on the axes of efficiency and utility (e.g., perplexity) in assessing performance; recent work has shown that compression may also have outsized effects on issues such as model fairness and robustness, which would be interesting additional aspects to consider in future study.
-Appendix A Main Experiment Details
-A.1 Hyper-parameters for all Bonsai regression during pruning
+5:     Sample $\{\mathbf{\bar{m}}_{i}\}_{[n_{\mathrm{iter}}]}$ sub-models according to $\bar{\mathbb{\rho}}^{l}$
+6:     Run forward pass on each sub-model and compute $U$. Construct $\mathbb{D}^{l}=\left\{\bar{\mathbf{m}}_{i},U_{i}\right\}_{[n_{\mathrm{iter}}]}$
+7:     $\beta^{\l}\leftarrow\mathrm{Regress}\left(\mathbb{D}^{l}\right)$
+8:     $\{m_{\mathrm{pruned}}\}\leftarrow$ sort $\beta^{\l}$ and drop the bottom $k$ modules that make up $p_{\mathrm{iter}}$ fraction of the model.
+9:     $\mathbf{m}\leftarrow$ update module list to exclude $\{m_{\mathrm{pruned}}\}$
+10:  end for
+11:  Output: Pruned model $\mathbf{M}_{|\mathbf{m}}$
+A note about Line $5$ in our algorithm.
+Depending on the task, we find that sampling $\mathbf{\bar{m}}_{i}$ and its complement $\mathbf{\bar{m}}^{c}_{i}$ helps reduce the variance of our regression estimate and leads to better results.
+C.2 Hyper-parameters for all Bonsai regression during pruning
 When using Bonsai, we estimate $\beta$ from $\mathbb{D}^{l}$ by performing linear regression via gradient descent with Adam (Kingma & Ba, 2014).
@@
 We do this because we do not care about matching $U_{k}$ exactly for each sub-model $k$; we rather care that our learned $\beta$ predicts the correct rankings amounts sub-models, which would denote that $\beta$ reasonably models relative module importances.
-In general, we use $\ell_{1}$-norm regularization on $\beta$ for all experiments. For the Phi-2 experiment in Section 6.3, we find that $\ell_{2}$-norm works better.
-A.2 Forward Pass Only Experiments
-Table 10 show the Bonsai hyperparameters we used for the experiments in Section 6.1.
+In general, we use $\ell_{1}$-norm regularization on $\beta$ for all experiments. For the Phi-2 experiment in Section 3.2.3, we find that $\ell_{2}$-norm works better.
+C.3 Forward Pass Only / Semi-structured pruning Experiments
+Table 10 show the Bonsai hyperparameters we used for the experiments in Section 3.2.1.
 Table 9: Bonsai hyper-params for forward only experiments
-Table 10: Bonsai  fine-tuning HP for pruned LLaMA family models
 $p_{\mathrm{iter}}$ $\mathrm{ns}_{\mathrm{sub-models}}$ $\mathrm{ns}_{\mathrm{data}}$ Metric for $\rho$
 0.05 200 32 (per-iter) Wanda
+Table 10: Bonsai  fine-tuning HP for pruned LLaMA family models
 LR rank LoRA-$\alpha$ $\lambda$ (Distill Weight) LoRA Modules
 1e-4 128 4$\times$×rank 0.01 All Modules
-Table 10: Bonsai  fine-tuning HP for pruned LLaMA family models
 For Wanda(Sun et al., 2023), we use the default hyper-parameters specified by the paper repo here for pruning.
@@
 For fine-tuning the Phi-2 model on Wikitext-2, we use the same hyper-parameters as Bonsai in Table 10.
-A.3 Experiments comparing to Gradient based structured pruning
+C.4 Experiments comparing to Gradient based structured pruning
 We compare to LoRA-Prune and LLM-Pruner.
@@
 Whilst we use 1 A6000 GPU (48G) for all experiments, LoRA-Prune uses A100 GPU (80G) for pruning LLaMA-1 7B.
-All Bonsai hyper-parameters are the same as Appendix A.2 except for $\mathrm{ns}_{\mathrm{sub-models}}$ which we set to $1000$.
-A.4 Phi-2 pruning experiment details
-For the experiment in Section 6.3, All Bonsai hyper-parameters are the same as Appendix A.2 except for the following changes:
+All Bonsai hyper-parameters are the same as Appendix C.3 except for $\mathrm{ns}_{\mathrm{sub-models}}$ which we set to $1000$.
+C.5 Phi-2 pruning experiment details
+For the experiment in Section 3.2.3, All Bonsai hyper-parameters are the same as Appendix C.3 except for the following changes:
   • $\mathrm{ns}_{\mathrm{sub-models}}=2000$
@@
   • Due to its relatively small size, the 1.8B pruned model can be fully fine-tuned on a single A6000 GPU over 100k sequences of length 2,048 tokens from the C4 dataset instead of using LoRA.
-Appendix B Impact of regression and perturbation ablation details
-For the experiment in Section 3, All Bonsai hyper-parameters are the same as Appendix A.2 except $p_{\mathrm{iter}}=0.1$ to speed up pruning.
+Appendix D Impact of regression and perturbation ablation details
+For the experiment in Section 6, All Bonsai hyper-parameters are the same as Appendix C.3 except $p_{\mathrm{iter}}=0.1$ to speed up pruning.
 A simple alternative to Bonsai is to leverage the prior $\rho$, computed from the unperturbed parent model, and make pruning decisions exclusively according to this.
@@
 Note that we use a Wanda based metric to define $\rho$ for this experiment.
-Module level analogues of the unstructured pruning metrics we explore are defined in Appendix F.
+Module level analogues of the unstructured pruning metrics we explore are defined in Appendix H.
 Table 11: Experiment on linear regression to estimate module importances. Wikitext-2 Perplexity. LLaMA-2 7B pruned to 50% sparsity
@@
 Yes $1.77$ $61.63$ $9.15$
-Appendix C Varying the pruning fraction per-iteration
-For the experiment in Section 3, All Bonsai hyper-parameters are the same as Appendix A.2 except we vary $p_{\mathrm{iter}}$.
+Appendix E Varying the pruning fraction per-iteration
+For the experiment in Section 6, All Bonsai hyper-parameters are the same as Appendix C.3 except we vary $p_{\mathrm{iter}}$.
 Table 12: Varying the fraction pruned at a time. Wikitext-2 Perplexity. LLaMA-2 7B pruned to 50% sparsity
@@
 0.20 $1.67$ $209.44$ $9.57$
-Appendix D Varying the number of calibration data points for pruning
-All Bonsai hyper-parameters are the same as Appendix A.2 except we vary $\mathrm{ns}_{\mathrm{data}}$ and $p_{\mathrm{iter}}=0.1$ to speed up pruning.
+Appendix F Varying the number of calibration data points for pruning
+All Bonsai hyper-parameters are the same as Appendix C.3 except we vary $\mathrm{ns}_{\mathrm{data}}$ and $p_{\mathrm{iter}}=0.1$ to speed up pruning.
 Table 13: How many data-points to consider during forward passes. Wikitext-2 Perplexity. Llama-2 7B pruned to 50% sparsity
@@
 32 $61.63$ $9.15$
-Appendix E Post-pruning adaptation
-For this experiment, All Bonsai hyper-parameters are the same as Appendix A.2.
-Appendix F Impact of prior
-For this experiment, All Bonsai hyper-parameters are the same as Appendix A.2 except we vary $\rho$
-F.1 $\rho$ is Activation Magnitude
+Appendix G Post-pruning adaptation
+For this experiment, All Bonsai hyper-parameters are the same as Appendix C.3.
+Appendix H Impact of prior
+For this experiment, All Bonsai hyper-parameters are the same as Appendix C.3 except we vary $\rho$.
+H.1 $\rho$ is Activation Magnitude
 MLP / Fully Connected Module: Let $d$ be the intermediate dimension of the MLP to be pruned.
@@
 $\left(\rho\in\mathbb{R}^{h}\right)\propto\hat{\mathbf{a}}$ (5)
-F.2 $\rho$ is Wanda (Sun et al., 2023)
+H.2 $\rho$ is Wanda (Sun et al., 2023)
 MLP / Fully Connected Module: Let $d$ be the intermediate dimension of the MLP to be pruned.
@@
 We can flatten $\mathbf{a}\in\mathbb{R}^{B\times S\times d_{h}\times h}\rightarrow\mathbb{R}^{BSd_{h}\times h}$ and then use the same formula as Equation 5 above to calculate $\rho\in\mathbb{R}^{h}$.
-Appendix G How many perturbative samples are reasonable?
-For this experiment, All Bonsai hyper-parameters are the same as Appendix A.2 except $p_{\mathrm{iter}}=0.1$ to speed up pruning.
+H.3 $\rho$ is Fluctuation-based (FLAP) An et al. (2024)
+MLP / Fully Connected Module: Let $d$ be the intermediate dimension of the MLP to be pruned.
+Let $W\in\mathbb{R}^{d\times o}$ be the output projection matrix for the MLP.
+For any data-sample sequence $b$, we compute the sample variance of each input feature across batches and weight it with the squared norm of the corresponding column of the weight matrix:
+$\left(\rho\in\mathbb{R}^{d}\right)\propto S_{:,j}=\frac{1}{N-1}\sum_{n=1}^{N}(X_{n,j,:}-\bar{X}_{:,j,:})^{2}\cdot||W_{:,j}||_{2}^{2}$ (7)
+where $\bar{X}_{:,j,:}$ represents the average of the $j$-th channel for all samples, and $||W_{:,j}||_{2}^{2}$ denotes the squared norm of the $j$-th column of the weight matrix.
+Self-Attention Module: We compute the fluctuation at the head level, weighted by the corresponding weights in the output projection matrix:
+$\left(\rho\in\mathbb{R}^{h}\right)\propto S_{:,j}=\frac{1}{N-1}\sum_{n=1}^{N}(X_{n,j,:}-\bar{X}_{:,j,:})^{2}\cdot||W_{:,j}||_{2}^{2}$ (8)
+For our improved fluctuation metric (fluct.2.0), we track the mean and second moment separately over multiple batches, computing variance as:
+$\mathrm{Var}(X)=\mathbb{E}[X^{2}]-\mathbb{E}[X]^{2}$ (9)
+This approach provides greater numerical stability when accumulating statistics across multiple forward passes.
+Appendix I How many perturbative samples are reasonable?
+For this experiment, All Bonsai hyper-parameters are the same as Appendix C.3 except $p_{\mathrm{iter}}=0.1$ to speed up pruning.
 Table 14: Varying the number of sub-models generated. Wikitext-2 Perplexity. LLaMA-2 7B pruned to 50% sparsity
@@
 This indicates that Bonsai actually recovers good modules even using as few samples as 50 sub-models.
-Appendix H Mistral-7B Experiment Details
-In addition to the primary experiments on the LLaMA and Phi-2 models, supplementary experiments were performed on the Mistral-7B (Jiang et al., 2023) model in comparison with Wanda results on the stated model.
+Appendix J Mistral-7B Experiment Details
+In addition to the primary experiments on the LLaMA and Phi-2 models, supplementary experiments were performed on the Mistral-7B Jiang et al.
+(2023) model in comparison with Wanda results on the stated model.
 We apply Bonsai with the same hardware and configuration settings as used for the LLaMA and Phi-2 experiments.
 We target different pruning fractions (0.05, 0.1, and 0.2) across different numbers of samples and masks per iteration to evaluate the method’s performance under varying sparsity conditions.
-The Mistral-7B model architecture differs from the LLaMA architecture in its use of group query attention and sliding window attention in lieu of the standard self-attention used in most transformer-based models like LLaMA (Jiang et al., 2023).
+The Mistral-7B model architecture differs from the LLaMA architecture in its use of group query attention and sliding window attention in lieu of the standard self-attention used in most transformer-based models like LLaMA Jiang et al.
+(2023).
 We factor these differences into consideration in the implementation of Bonsai for Mistral.
-For the experiments that produced the results below, all Bonsai hyper-parameters are the same as Appendix A.2.
+For the experiments that produced the results below, all Bonsai hyper-parameters are the same as Appendix C.3.
 Table 15 presents the test perplexity results for Mistral-7B under different pruning methods.
@@
 We notice a recurring theme: when an attention layer is significantly altered, it leads to compensation in the next layers within the sequence.
-This adaptive behavior, termed the ”Hydra effect” by McGrath et al.
-(2023), implies that the layers within a language model interact in a way that changes in one layer prompt adjustments in another.
-McGrath et al.
-(2023) specifically mentioned that when one attention layer was removed from a language model, the model was still able to self-repair and produce similar outputs; but it did so by relying more heavily on other layers.
+This adaptive behavior, termed the ”Hydra effect” by (McGrath et al., 2023), implies that the layers within a language model interact in a way that changes in one layer prompt adjustments in another.
+(McGrath et al., 2023) specifically mentioned that when one attention layer was removed from a language model, the model was still able to self-repair and produce similar outputs; but it did so by relying more heavily on other layers.
 Figure 6: Mistral’s pruned attention layers.
-The heavily pruned layers are usually preceded by or sandwiched between lightly-pruned layers, exhibiting the self-repairing ”Hydra effect” (McGrath et al., 2023).
+The heavily pruned layers are usually preceded by or sandwiched between lightly-pruned layers, exhibiting the self-repairing ”Hydra effect” McGrath et al.
+(2023).
+Should Bonsai prune iteratively?
+Table 16 demonstrates the benefits of using Bonsai in an iterative fashion.
+Pruning slowly ($p_{\mathrm{iter}}=0.05$) yields the best results, but this comes at the cost of increasing the total time to prune the model.
+The performance gap between values of $p_{\mathrm{iter}}$ persists even after post-pruning adaptation, indicating that slower pruning allows for more accurate estimates of module importance.
+Table 16: Varying $p_{\mathrm{iter}}$. Wikitext-2 perplexity of LLaMA-2 7B pruned to 50% sparsity. See Appendix E for experiment details.
+$p_{\mathrm{iter}}=0.05$ $p_{\mathrm{iter}}=0.1$ $p_{\mathrm{iter}}=0.2$
+w/o Adapt $19.47$ $61.63$ $209.44$
+w Adapt $8.89$ $9.15$ $9.57$

### v4

@@
 Abstract
-Structured pruning is a promising approach to create smaller, faster LLMs.
-However, existing methods typically rely on backward passes, which can inflate memory requirements and compute costs.
+Structured pruning is a promising approach to create smaller, faster large language models.
+However, existing methods typically rely on computing the gradient via backward passes, which can inflate memory requirements and compute costs.
 In this work we introduce Bonsai, a gradient-free structured pruning method that eliminates the need for backpropagation, significantly reducing memory requirements and compute costs while achieving state-of-the-art pruning performance.
@@
 Unlike existing structured pruning approaches, Bonsai not only achieves better compression with fewer resources, but also produces models that are twice as fast as those generated by semi-structured pruning.
-As a concrete demonstration, we use Bonsai to prune an 8B LLaMA-3 model to 50% sparsity on a single A6000 GPU–a task infeasible with backprop-based methods, which require $2$-3×3$\times$3 × memory.
+As a concrete demonstration, we use Bonsai to prune 7B and 8B models to 50% sparsity on a single A6000 GPU—a task challenging for backprop-based methods in memory-constrained settings, as they require $2$-$3\times$ the memory.
 Our results show that removing backprop as a requirement not only enables pruning larger models on constrained hardware but can also lead to state-of-the-art efficiency and performance.
 1 Introduction
-Figure 1: Perplexity versus inference speed-up of pruned models for methods that use only forward passes on the parent model.
-At any given target perplexity, Bonsai produces the fastest model, resulting in improved latency and throughput.
-Circle sizes $\propto$ model’s memory footprint.
-[…]
-In light of this challenge, we present Bonsai, a structured pruning approach that operates exclusively with forward passes through the parent model.
-By eliminating the need for backward passes for pruning, Bonsai dramatically reduces memory requirements that enable pruning of models that would otherwise be infeasible on available hardware, and can do so while outperforming existing structured and unstructured pruning approaches (Fig 1).
+Figure 1: Memory requirements can make pruning methods inaccessible.
+Gradient-based structured pruning methods require substantially more memory relative to forward-pass-only approaches like Bonsai, which can operate in memory-constrained settings (e.g., $\leq$24GB) while achieving superior performance among accessible methods.
+Results for LLaMA-2 7B at 50% sparsity on Wikitext-2.
+The increasing scale of large language models (LLMs) has amplified the computational and memory requirements for their deployment and adaptation, presenting a significant barrier to widespread adoption.
+While structured pruning has emergeding methods require substantially more memory relative to as a popular approach to create smaller, faster models (Wang et al., 2019; Xia et al., 2022a), many existing methods rely on backpropagation, which significantly inflates memory and compute costs: backward passes consume $\gtrapprox 2\times$ (Bridger, 2023) the memory of a forward pass, with popular stateful optimizers like AdamW (Loshchilov and Hutter, 2017) requiring $\gtrapprox 3\times$ more memory.
+This often makes them impractical for a wide range of practitioners who operate under resource constraints, such as students, researchers, and small organizations with limited access to enterprise-grade, multi-GPU systems.
+In response to this challenge, we present Bonsai, a structured pruning approach that operates exclusively with forward passes through the parent model.
+By eliminating the need for backward passes for pruning, Bonsai dramatically reduces memory requirements to enable the compression of models on commodity hardware and in memory-constrained environments—potentially obviating the need for more expensive, enterprise-grade solutions.
+Notably, Bonsai maintains lower memory requirements while still outperforming existing structured and unstructured pruning approaches in terms of compression/accuracy trade-offs (Fig 1).
+The key challenge in structured pruning is determining which modules (e.g., attention heads, layer dimensions) are most important for model performance and can be safely removed.
 Rather than computing gradients to determine module importance, Bonsai employs an innovative perturbative approach, estimating module importances by evaluating the performance of sub-models using only inference.
@@
 First, we formulate module importance estimation as an underdetermined regression problem, allowing us to infer the importance of a large number of modules by exploring a manageable number of random sub-models.
-This fundamentally differs from previous pertubative approaches, which require evaluating approximately as many sub-models as there are modules (Ancona et al., 2020), making them intractable for LLMs.
-Second, we use informative priors derived from unstructured pruning (Han et al., 2015; Sun et al., 2023) to guide sub-model exploration, yielding better estimates of module relevance with fewer evaluations, instead of instantiating sub-models by dropping modules with equal likelihood (Kang et al., 2023).
-Finally, unlike prior gradient-free approaches that prune layer-by-layer (Dekhovich et al., 2021; Nova et al., 2023; Sun et al., 2023), Bonsai takes a holistic view across the entire model: modules across layers are removed and evaluated together, with relevance scores computed globally to make pruning decisions that better preserve accuracy.
-Through extensive experiments, we demonstrate Bonsai’s effectiveness.
-When compared to other structured pruning methods that also avoid backward passes, including a forward-only structured pruning version of Wanda (Sun et al., 2023), Bonsai consistently produces models with better perplexity at any given speedup target.
-Wen the resulting pruned model small enough to be fine-tuned on available hardware, Bonsai compared to semi-structured pruning methods like Wanda (Sun et al., 2023)—which is also fine-tuned—Bonsai not only achieves similar performance, but also delivers 2×2$\times$2 × faster inference (an overall inference speed-up of 1.58×1.58$\times$1.58 × compared to the parent model).
-Even when compared to the state-of-the-art gradient-based structured pruning methods like LLM-Pruner and LoRAPrune, Bonsai produces models that outperform these more memory-intensive approaches on 4 out of 6 evaluations, despite using significantly less memory during pruning.
+This fundamentally differs from previous perturbative approaches, which require evaluating approximately as many sub-models as there are modules (Ancona et al., 2020), making them intractable for LLMs.
+Second, we use informative priors derived from existing pruning approaches (Han et al., 2015; Sun et al., 2023; An et al., 2024) to guide sub-model exploration, yielding better estimates of module relevance with fewer evaluations, instead of instantiating sub-models by dropping modules with equal likelihood (Kang et al., 2023).
+Finally, unlike prior approaches that prune layer-by-layer (Dekhovich et al., 2021; Nova et al., 2023; Sun et al., 2023; Ling et al., 2024), Bonsai takes a holistic view across the entire model: modules across layers are removed and evaluated together, with relevance scores computed globally to make pruning decisions that better preserve accuracy than the sub-optimality of localized methods.
+Through extensive experimentation, we demonstrate that Bonsai’s approach is highly efficient and effective.
+When compared to other structured pruning methods that also avoid backward passes, including FLAP (An et al., 2024) and a forward-only structured pruning version of Wanda (Sun et al., 2023), Bonsai consistently produces models with better perplexity at any given speedup target.
+Critically, even when compared to the state-of-the-art gradient-based structured pruning methods like LLM-Pruner and LoRAPrune, Bonsai produces models that majorly outperform these more memory-intensive approaches, despite using significantly less memory during pruning.
 Our results lead us to explore Bonsai’s practical utility by pruning the 3B Phi-2 (Li et al., 2023) model to a 1.8B model that performs competitively against other sub-2B parameter models on the Huggingface Open LLM leaderboard.
-Overall, by reducing the memory and computational requirements of structured pruning, Bonsai enables more efficient adaptation of LLMs across a wider range of hardware configurations.
-This makes these LLMs more accessible while maintaining their performance advantages, representing an important step forward for practical LLM compression.
+A key advantage of Bonsai is that it not only reduces memory requirements during pruning but also unlocks the opportunity for efficient subsequent post-pruning adaptation.
+As the pruned model is small enough to fit on the same hardware that was used for inference, practitioners can then perform parameter-efficient fine-tuning to further recover performance.
+This capability addresses a fundamental misconception that gradient-based methods are always superior: in many cases, Bonsai is a prerequisite for reaching a stage where such fine-tuning is even possible, expanding access to the entire model adaptation pipeline and yielding superior end performance.
+Our approach thus provides a tangible solution for practitioners who do not have access to the high-cost resources typically associated with LLM development.
+2 Related Work
+We first discuss relevant work in LLM pruning and other complementary compression methods.
+2.1 Unstructured Pruning
+While structured pruning removes entire components like layers (Xu et al., 2020; Xia et al., 2022b), dimensions of linear layers (Wang et al., 2019) or attention heads (Michel et al., 2019; Held and Yang, 2022), unstructured pruning (Han et al., 2015; Frankle and Carbin, 2018; Benbaki et al., 2023; Sun et al., 2023) removes individual parameters of the model.
+These approaches achieve memory savings by inducing sparsity in the model weights, but they generally do not result in tangible model speedups except when specialized hardware is available (Mishra et al., 2021).
+Proposed semi-structured sparsity methods (Mishra et al., 2021) such as 2:4 and 4:8 patterns can provide faster inference, but the speedup gains they achieve are far from the idealized $2\times$.
+2.2 Gradient-Based Structured Pruning
+Most existing structured pruning techniques for large (over 1B scale) language models rely on gradient computation to estimate module importance.
+LLM-Pruner (Ma et al., 2023) uses first-order Taylor expansion to measure the importance of coupled structures in transformers, followed by LoRA fine-tuning for performance recovery.
+LoRAPrune (Zhang et al., 2024) reduces memory requirements compared to LLM-Pruner by using LoRA gradients rather than full model gradients for importance estimation, though it still requires gradient computation during the pruning process.
+Sheared LlaMA (Xia et al., 2024) takes a different approach by learning pruning masks during continued pre-training, integrating structured pruning with the training process.
+More recently, SparseLLM (Bai et al., 2024) decomposes the global pruning problem into manageable subproblems using auxiliary variables, though it maintains reliance on gradient information.
+Ultimately, a fundamental limitation of gradient-based approaches is their memory requirements.
+Computing gradients for large models during pruning significantly increases memory usage beyond what is needed for inference alone, making these methods challenging to apply on resource-constrained hardware.
+In this work we show that it is possible to remove these additional memory requirements while still achieving comparable or superior performance to gradient-based pruning methods.
+2.3 Memory-Efficient Structured Pruning Approaches
+The memory constraints of gradient-based methods have motivated research into gradient-free alternatives, though such work remains more limited for LLMs.
+SliceGPT (Ashkboos et al., 2024) utilizes PCA to remove up to 25% of weight matrices, although it selects embedding reduction directions based on variance, which may not correlate with actual model utility for downstream tasks.
+In the same vein, SlimGPT (Ling et al., 2024) extends the classical Optimal Brain Surgeon (OBS) framework (Hassibi and Stork, 1992) to LLMs to reduce memory costs; however, the work admits that its layer-wise nature can lead to ‘error accumulation’ and an inability to use global information when pruning; we find that this makes it substantially less performant than Bonsai’s global approach.
+For smaller language models like BERT, Nova et al.
+(2023) proposed Kernelized Convex Masking (KCM) for gradient-free structured pruning.
+Unfortunately, to prune a fully connected layer with $K$ intermediate units, KCM requires instantiating a $K\times K$ coefficient matrix for each layer.
+While this is feasible for BERT models, a typical LLM (AI, 2023; Touvron et al., 2023) has $K\sim 10^{4}-10^{5}$ which would make the size of each per-layer coefficient matrix comparable to the size of the LLM itself.
+In computer vision, there exist several perturbative gradient-free structured pruning techniques (Ancona et al., 2020; Dekhovich et al., 2021).
+However, these methods have been exclusively applied to small vision models (e.g., VGG (Simonyan and Zisserman, 2014) and WideResnet (Zagoruyko and Komodakis, 2016)) and would not scale to LLMs as-is.
+Most comparably, FLAP (An et al., 2024) introduces a fluctuation-based metric that operates without gradients using statistical fluctuations in activations to estimate module importance.
+We show in Section 4.4 and Appendix E.3 that Bonsai’s flexible framework outperforms FLAP and can in fact incorporate FLAP’s fluctuation-based importance metric as a special case, being a more general methodology.
+Finally, Probe Pruning (Le et al., 2025) introduces a different approach with “online, dynamic, structured pruning" that operates at inference time via model probing, and LoRAM (Zhang et al., 2025) trains LoRA adapters on a pruned model and uses recovered low-rank matrices with the original model for inference.
+Both of these methods tackle different problems to Bonsai—online pruning/fine-tuning rather than one-time permanent pruning—but they demonstrate the broader, shared trend toward memory-efficient model adaptation.
+2.4 Positioning of Bonsai
+Overall, the existing literature shows a clear gap: there is a need for methods that can operate within memory-constrained environments while achieving competitive structured pruning performance for LLMs.
+The highest-performing structured pruning approaches require computational resources that exceed what is available on single consumer GPUs.
+Our approach addresses this gap by formulating structured pruning as an underdetermined regression problem that can be solved using only forward passes through the model.
+Key distinctions of our approach include:
+  • Unlike methods requiring gradient computation or auxiliary mathematical operations like Hessian storage, Bonsai operates solely through model inference.
+  • Rather than making layer-wise pruning decisions, our method estimates module importance globally across the entire model through perturbative evaluation.
+  • Our approach requires only the memory needed for inference plus a small overhead for storing perturbation results, making it accessible on standard hardware.
+  • The regression-based formulation provides theoretical grounding for estimating module importance from limited samples, avoiding purely heuristic approaches.
+2.5 Additional Methods for LLM Compression
+Finally, although we focus on structured pruning in this work, we note that prior research has explored various other compression schemes such as distillation (Hinton et al., 2015; Gu et al., 2023) and quantization (Xiao et al., 2023) to create smaller models from larger pre-trained ones.
+Similar to structured pruning, these compression methods themselves often impose significant computational burdens.
+For example, distillation-based techniques require using LLMs to generate large amounts of teacher data (Jiao et al., 2019; Hsieh et al., 2023).
+Although unstructured pruning (Frantar and Alistarh, 2023; Sun et al., 2023) and quantization have lower training-time resource demands, the models they produce either require specialized hardware to achieve speedups (Mishra et al., 2021) or may actually slow down inference due to additional computational overhead (Dettmers et al., 2022).
+Table 1 summarizes these characteristics across different compression approaches.
+[…]
+Regime Resource Approaches
+Quantization (Mixed Precision) Distillation Unstructured Pruning Gradient-Based Structured Pruning Bonsai (Ours)
+Train Memory
+Compute
+Inference Memory
+Compute
+Overall, each approach has distinct hardware requirements and performance characteristics that make them suitable for different deployment scenarios.
+Our structured pruning approach could potentially be combined with these techniques; for instance, quantization could be applied post-pruning, and structured pruning could serve as initialization for distillation approaches.
+We leave such combinations for future work and focus specifically on the forward-pass-only structured pruning problem.
 3 Methodology
-We first cover background in LLM pruning, and then discuss our structured pruning method that exclusively performs inference on the parent model.
+We cover background in LLM pruning (§3.1), and then discuss Bonsai, our structured pruning method that exclusively performs inference on the parent model to prune.
+Figure 2 provides an overview of our approach as detailed in the following subsections: §3.2 describes our perturbative estimation; §3.3 explains our informative prior-based sampling strategy; and §3.4 discusses our iterative pruning procedure.
 3.1 Background on Pruning, Problem Definition and Notation Setup
-We are given an LLM, $\mathbf{M}_{\theta}$, parameterized by $\theta\in\mathbb{R}^{D}$.
-Also provided is $U$, a utility function that evaluates a model’s performance on a target task.
+We assume that we are given an LLM, $\mathbf{M}_{\theta}$, parameterized by $\theta\in\mathbb{R}^{D}$.
+Also provided is $U$, a utility function that evaluates the model’s performance on a target task.
+We instantiate $U$ as language modeling perplexity: Wikitext-2 training for Wikitext-2 experiments and C4 for zero-shot tasks.
 We are interested in pruning $\mathbf{M}_{\theta}$ to produce a smaller and faster but still performant (with respect to $U$) model under the constraint that we only have enough memory to run inference on $\mathbf{M}_{\theta}$.
-Even though we assume we can run $\mathbf{M}_{\theta}$ on available hardware, pruning can be critical for achieving latency targets, reducing compute burden, or making the model small enough to adapt to new (out-of-domain) tasks by gradient-based fine-tuning.
+Even though we assume we can run $\mathbf{M}_{\theta}$ on available hardware, pruning can be critical for achieving latency targets, reducing compute burden, or making the model small enough to adapt to new (out-of-domain) tasks by performing gradient-based fine-tuning.
 Unstructured pruning approaches compress $\mathbf{M}_{\theta}$ by removing individual parameters $\theta_{j}$ from the model.
 This results in the updated model consisting of sparsified weight matrices with a smaller memory footprint.
-Unfortunately, the updated model does not enjoy inference speedups except when specialized hardware is available and thus poses a compute burden during inference.
-Whilst semi-structured variants – those that remove parameters in patterns like 2:4 or 4:8 (Mishra et al., 2021) – achieve some speedup, these are modest compared to those achieved with structured pruning.
+Unfortunately, the updated model may not enjoy inference speedups except when specialized hardware is available and thus can pose a compute burden during inference Mishra et al.
+(2021).
+While semi-structured variants – those that remove parameters in patterns like 2:4 or 4:8 (Mishra et al., 2021) – achieve some speedup, these are modest compared to those achieved with structured pruning.
 Structured pruning takes a more modular view of the units to be removed from $\mathbf{M}_{\theta}$.
-Take that $\mathbf{M}_{\theta}$ is made up of modules $\mathbf{m}=\{m_{i}\}_{i\in[N]}$ each with corresponding parameter counts $\mathbf{s}=\{s_{i}\}_{i\in[N]}$ such that $\sum_{i}s_{i}=D$.
+Consider that $\mathbf{M}_{\theta}$ is made up of modules $\mathbf{m}=\{m_{i}\}_{i\in[N]}$ each with corresponding parameter counts $\mathbf{s}=\{s_{i}\}_{i\in[N]}$ such that $\sum_{i}s_{i}=D$.
 For a transformer model, $\mathbf{m}$ could consist of attention heads, dimensions of fully connected layers, or even whole layers.
@@
 Many structured pruning methods attempt to solve Equation 1 by gradient-guided optimization (or search) over the space of sub-models.
-However, since we are interested in the memory-constrained setting where computing gradients is not feasible, these methods cannot be used.
-Figure 2: Overview of Bonsai.
-Perturbations to the original model are generated according to an intelligent prior.
-These perturbed models are evaluated on downstream task and the collected data is used to build a simple linear model of module importances.
-Bonsai iteratively drops modules of least importance until the target model size is reached.
+In this work, we instead explore approaches suitable for memory-constrained settings, where computing gradients may be expensive or infeasible.
+Figure 2: Bonsai estimates module importance through regression on perturbative evaluations.
+Rather than requiring as many sub-models as there are modules (intractable for LLMs), we solve an underdetermined regression problem: sample $n\ll N$ sub-models according to informative priors, evaluate each via for forward pass, then regress to estimate global importance scores $\bm{\beta}$.
+Black squares indicate removed modules.
 3.2 Estimating module relevance with only forward passes
-We consider pruning a model too large to fit in memory, allowing only forward passes.
+We specifically aim to develop an approach for pruning that only relies on computing the forward pass over the model.
 Thus, we must solve Eq 1 using evaluations of $U$ rather than gradient-based optimization.
 A brute-force search over $\mathcal{F}_{p}$ is infeasible, as its size grows combinatorially with the model.
-For instance, pruning a single FC sub-layer in an LLM to 50% sparsity requires evaluating $\approx$ ($10^{4}\choose 10^{3}$ subsets.
-We propose a computationally tractable approach where we first perform a small number, $n$, of evaluations, where $n\ll|\mathcal{F}_{p}|$, to gather data for estimating the relevance of each module in $\bm{M}$ with respect to the metric $U$.
+For instance, pruning a single FC sub-layer in an LLM to 50% sparsity requires evaluating $\approx$ $10^{4}\choose 10^{3}$ subsets.
+We instead propose a computationally tractable approach where we first perform a small number, $n$, of evaluations, where $n\ll|\mathcal{F}_{p}|$, to gather data for estimating the relevance of each module in $\bm{M}$ with respect to the metric $U$.
 Upon carrying out this estimation, we can greedily choose the member of $\mathcal{F}_{p}$ that has the highest total module relevance.
@@
 We then frame the estimation of $\bm{\beta}$ as an under-specified regression problem:
-$\begin{split}\hat{\bm{\beta}}&=\mathrm{argmin}_{\bm{\beta}\in\mathbb{R}^{N}}\quad\frac{1}{n}\sum_{\left(\bar{\bm{m}}_{k},U_{k}\right)\in\mathbb{D}}\left(U_{k}-\beta^{T}\alpha_{\bar{\bm{m}}_{k}}\right)^{2}+\gamma\|\bm{\beta}\|\\ \end{split}$ (3)
+[…]
 where $\left(\alpha_{\bar{\bm{m}}_{k}}\right)_{i}=\bm{1}[i\in\bar{\bm{m}}_{k}]$, is the binary vector that has 0 at indices with modules dropped.
@@
 Instead, we create them virtually by zeroing out the outputs of the parts to be pruned so they have no effect on the model output.
+Algorithm 1 Bonsai Pruning Method
+1:Input:
+2:Model $\mathbf{M}_{\theta}$], sub-models per iteration $n_{\mathrm{iter}}$]
+3:Sparsity per iteration $p_{\mathrm{iter}}$], Target sparsity [$p$]
+4:Module list $\mathbf{m}$]
+5:
+6:for $l=1$ to $\lceil\frac{p}{p_{\mathrm{iter}}}\rceil$ do
+7:  $\mathbb{\rho}^{l}\leftarrow$ Calculate unstructured pruning metric for all modules in $\mathbf{m}$
+8:  $\bar{\mathbb{\rho}}^{l}\leftarrow$ Fix the top $(1-2p_{\mathrm{iter}})$ of $\mathbb{\rho}^{l}$ to $\infty$
+9:  Sample $\{\mathbf{\bar{m}}_{i}\}_{[n_{\mathrm{iter}}]}$ sub-models according to $\bar{\mathbb{\rho}}^{l}$
+10:  Run forward pass on each sub-model and compute $U$. Construct $\mathbb{D}^{l}=\left\{\bar{\mathbf{m}}_{i},U_{i}\right\}_{[n_{\mathrm{iter}}]}$
+11:  $\beta^{\l }\leftarrow\mathrm{Regress}\left(\mathbb{D}^{l}\right)$
+12:  $\{m_{\mathrm{pruned}}\}\leftarrow$ sort $\beta^{\l }$ and drop the bottom $k$ modules that make up $p_{\mathrm{iter}}$ fraction of the model.
+13:  $\mathbf{m}\leftarrow$ update module list to exclude $\{m_{\mathrm{pruned}}\}$
+14:end for
+15:Output: Pruned model $\mathbf{M}_{|\mathbf{m}}$
 3.3 Selecting sub-models for evaluation
-An as yet unexplored design choice is selecting the $n$ candidate sub-models for evaluation.
-A naive approach–uniform sampling–is shown here to be suboptimal.
+An important design choice is selecting the $n$ candidate sub-models for evaluation.
+A naive approach such as uniform sampling would be suboptimal.
 Take $m_{i}$ as a module that is critical for good performance under evaluation with $U$.
-Since $n<N$, it means that some modules may never be ”turned on” in the list of $n$ chosen sub-models.
+Since $n<N$, it means that some modules may never be "turned on" in the list of $n$ chosen sub-models.
 If $m_{i}$ happens to be one of these masked-out modules, the resulting estimate of $\hat{\beta_{i}}=0$ would in turn result in $\bm{m}^{\mathrm{approx}}$ being a poor estimate for $\bm{m}^{\ast}$.
 Thus, a more informed selection is necessary for accurate and useful $\bm{\beta}$ estimates.
-Given a module $m_{i}$, we set the likelihood of it being present in any of the $n$ sampled sub-models proportional to a prior $\rho_{i}$ which reflects its usefulness.
+Given a module $m_{i}$, we instead set the likelihood of it being present in any of the $n$ sampled sub-models proportional to a prior $\rho_{i}$ which reflects its usefulness.
 To define $\rho_{i}$, we can turn to metrics from the pruning literature.
@@
 Since $\rho_{i}$ can be computed via forward passes through the unmodified model $\bm{M}_{\theta}$, this approach remains memory-efficient.
-Appendix H details the priors we explored.
-[…]
+Appendix E details the priors we explored.
+To enhance efficiency, we prune only the bottom $2p$ fraction of modules per layer, ranked by prior $\rho$, while keeping the top 12p1-2p fraction fixed.1112p2p is arbitrary; practitioners can tune it for optimal performance.
+This reduces the search space for sub-model evaluation.
+For the bottom $2p$ fraction, whenever we generate a mask $\alpha_{\bm{m}_{k}}$ with sparsity $p$, we also generate its complement $\alpha^{c}_{\bm{m}_{k}}$, obtained by flipping the values in $\alpha_{\bm{m}_{k}}$ (excluding the fixed 12p1-2p fraction).
+Covert and Lee (2020) show that this technique helps lower variance in regression with binary inputs.
 3.4 Iterated Pruning
-Previous work on gradient-based pruning (Anwar et al., 2017; Frankle & Carbin, 2018) have shown that taking an iterated approach to pruning yields improved results over pruning directly to the target sparsity $p$.
+Previous works on gradient-based pruning (Anwar et al., 2017; Frankle and Carbin, 2018) have shown that taking an iterated approach to pruning yields improved results over pruning directly to the target sparsity $p$.
 Similarly, we define an updated pruning fraction $p_{\mathrm{iter}}<p$ and perform $\mathrm{iter}=\lceil\frac{p}{p_{\mathrm{iter}}}\rceil$ steps where we explore $n_{\mathrm{iter}}=\lceil\frac{n}{\mathrm{iter}}\rceil$ sub-models at a time.
 At the beginning of each iteration, we re-estimate the priors $\bm{\rho}$ for the unpruned modules and use them to select the $n_{\mathrm{iter}}$ sub-models to evaluate.
-We combine the methods from Sections 2.2, 2.3, and 2.4 to develop Bonsai, a gradient-free structural pruning algorithm (Figure 2). Algorithm 1 provides the full details.
+We combine the methods from Sections 3.2, 3.3, and 3.4 to develop Bonsai, a gradient-free structural pruning algorithm (Figure 2).
+Algorithm 1 provides the full details.
+A note about Line $5$ in our algorithm: depending on the task, we find that sampling $\mathbf{\bar{m}}_{i}$ and its complement $\mathbf{\bar{m}}^{c}_{i}$ helps reduce the variance of our regression estimate and leads to better results.
 4 Experimental Details and Main Results
@@
 The minimum amount of memory required to run a LlaMA-7B model at half precision (FP16) is 14GB.
-Running a forward pass with a batch size of 1 using the default model sequence length of 4096 uses around 20Gb of memory.
+Running a forward pass with batch size of 1 using the default model sequence length of 4096 uses around 20GB of memory.
 Forward Only Gradient-Based
-Llama-2-7B (Model Only) Forward with bsz=1 (Bonsai Min.) Bonsai (Faster) LoRA Prune (Zhang et al., 2023) Compresso (Guo et al., 2023) LLM-Pruner (Ma et al., 2023) ShearedLlaMA (Xia et al., 2023)
-14GB 20GB A6000 (48GB) 1 A100 (80GB) 4 V100 (128GB) 2 A100 (160GB) 8 A100 (640GB)
+Base Model Bonsai (Min.) Bonsai (Faster) LoRA Prune Compresso LLM-Pruner Sheared LlaMA
+LlaMA-2-7B Forward + Bsz=1 (Zhang et al., 2024) (Guo et al., 2023) (Ma et al., 2023) (Xia et al., 2024)
+14GB 20GB 48GB 80GB 128GB 160GB 640GB
+FP16 only A6000 1×A6000 1×A100 4×V100 2×A100 8×A100
+In this section, we empirically evaluate Bonsai across multiple structured pruning settings.
+We compare Bonsai + PPA against both semi-structured pruning (Section 4.1) and gradient-based pruning (Section 4.2).
+Next, we demonstrate that Bonsai can yield compact models with strong zero-shot abilities (Section 4.3).
+Finally, we compare Bonsai to other forward-pass-only pruning methods (Section 4.4) to establish competitiveness under memory constraints.
 In all Bonsai experiments, we prune (1) self-attention heads and (2) fully connected layer dimensions, focusing on LLMs around 7B parameters.
@@
 To reduce variance in score estimates, we average over 32 data points.
-A practitioner with 20GB of memory would run 32 forward passes with batch size 1, but this would be too slow for our experiments.
+A practitioner with 20GB of memory would run 32 forward passes with batch size 1, but this extends experiment runtime.
 Instead, we use a 48GB A6000 GPU (although not required by Bonsai), enabling batch sizes of 4–8 for faster experimentation.
-3.1 Bonsai is competitive with other forward pass-only, structured pruning methods
-Figure 3: Perplexity at 50% sparsity of LLaMA-{1,2}-7B on Wikitext-2 and C4 datasets.
-Dataset Method Sparsity Llama-1/seqlen=1024 Llama-2/seqlen=1024
-Wikitext-2 Base Model 0% 5.68 5.11
-FLAP 50% 17.27 14.49
-Bonsai 50% 15.72 12.38
-C4 Base Model 0% 7.34 7.04
-FLAP 50% 24.98 24.0
-Bonsai 50% 22.31 20.7
-We focus our first set of experiments on comparing structured pruning methods that can be run without gradient-based optimization.
-We prune the LLaMA-2 7B model (Touvron et al., 2023) to 50% sparsity and evaluate on the Wikitext-2 (Merity et al., 2016) validation dataset.
-Our module importance signal for pruning, $U$, is the language modeling performance on the training set.
-When measuring speedups, we consider end-to-end latency of running inference on model.sequence_length chunks of the Wikitext-2 validation set.
-See Table 10 (Appendix C.3) for details about the hyper-parameters used for these experiments.
-Figure 1 shows the results of some of our experiments.
-Bonsai expectedly outperforms the structured variant of Wanda (Sun et al., 2023), producing models with much lower perplexity at all sparsity levels and at any fixed desired speedup over the parent model.
-The more competitive baseline is FLAP (An et al., 2024), which was published a little over a month before the first release of our work.
-Even here (as shown in Table 3),  Bonsai outperforms FLAP while maintaining comparable speedups.
-A key advantage of  Bonsai is its flexibility—it can incorporate various pruning methods as informative priors (magnitude, Wanda, or fluctuation-based metrics like those used in FLAP–see Appendix H for more details on how we estimate priors).
-Once these priors are established, Bonsai’s regression step refines the importance estimates, potentially yielding better pruning decisions.
-Bonsai offers practitioners flexible runtime-quality tradeoffs.
-While the reported results use a configuration that runs in $\leq~{}4hours$ (which is considerably more than FLAP’s $\approx 1$hr’s runtime), practitioners can configure  Bonsai  to produced structurally pruned models in as little as 15 minutes by reducing the number of perturbations, sample size, or pruning iterations—at at some cost to performance.
-Conversely, allocating more compute time generally yields better pruning outcomes.
-This flexibility allow practitioners to tailor Bonsai to their specific constraints, unlike methods with fixed computational profiles.
-3.2 Introducing Post-Pruning Adaptation (PPA)
-After pruning the parent model, depending on the sparsity level $p$ achieved, it is possible to obtain a pruned model on which we can run full fine-tuning or a parameter-efficient fine-tuning method like LoRA (Hu et al., 2021) with the available hardware memory.
-In this case, we can fine-tune the pruned model on the downstream task in order to recover more performance relative to that of the parent model.
+Introducing Post-Pruning Adaptation (PPA).
+A central strength of Bonsai is that, by reducing models to a size that fits on commodity inference hardware depending on the sparsity level $p$ achieved, it unlocks the ability to further fine-tune (full fine-tuning or a parameter-efficient fine-tuning method like LoRA (Hu et al., 2021)) the pruned model on the same hardware.
+PPA is not required for Bonsai to be effective; but it shows that Bonsai does more than prune, since it expands access to to the full model adaptation pipeline in settings where gradient-based pruning is infeasible from the start.
 Like many past works (Sanh et al., 2020; Xia et al., 2022b), we can combine pruning with distillation by incorporating a distillation loss in the training objective during fine-tuning of the pruned model.
@@
 Using $i$ to index the task data, we have:
-$\mathcal{L}_{\mathrm{distill}}=\sum_{i}D_{\mathrm{KL}}\big{(}\mathrm{logits}^{i}\left(\mathbf{M}_{|\mathbf{m}^{\mathrm{approx}}}\right)~{}\|~{}\mathrm{logits}^{i}\left({\mathbf{M}}\right)\big{)}$, where $\lambda$ is a scalar weighting that can cross validated.
+$\mathcal{L}_{\mathrm{distill}}=\sum_{i}D_{\mathrm{KL}}\big(\mathrm{logits}^{i}\left(\mathbf{M}_{|\mathbf{m}^{\mathrm{approx}}}\right)~\|~\mathrm{logits}^{i}\left({\mathbf{M}}\right)\big)$, where $\lambda$ is a scalar weighting that can be cross-validated.
 Note, distillation can be performed without significant memory overhead by a priori caching the logits from the parent model $\mathbf{M}$ instead of hosting the model in memory during fine-tuning.
-In the sections ahead, we will apply PPA after pruning the parent model if the child model is small enough to allow for this.
-Table 2: Wikitext-2 perplexity at 50% sparsity of LLaMA-2 7B with end-to-end latency speedups.
+In the subsections below, we will apply PPA after pruning the parent model if the child model is small enough to allow for this.
+Table 3: Pruning can match highly optimized small, pretrained models while delivering better speedups.
+Bonsai-pruned LLaMA-2 achieves accuracy comparable to Phi-2 (a carefully engineered 3B model) with superior inference efficiency.
+Semi-structured pruning loses speedup benefits due to incompatible sparsity patterns.
 […]
-LlaMA-2 7B $5.11$ 1×1$\times$1 ×
-Phi-2 3B $8.69$ 1.24×1.24$\times$1.24 ×
-Wanda 2:4 3B $10.52$ 1.14×1.14$\times$1.14 ×
-  + PPA $8.34$ 0.75$\times$
-Bonsai 3B $19.47$ $\boldsymbol{1.58}\times$
-  + PPA $8.89$ $\boldsymbol{1.58}\times$
-3.2.1 Bonsai is competitive with semi-structured pruning methods
-We compare Bonsai to the semi-structured variant of Wanda (Sun et al., 2023). In general, structured pruning methods under-perform semi-structured pruning methods, but compensate for this in speedup.
-Before fine-tuning, the Wanda 2:4 model is more accurate but slower (1.14$\times$ vs 1.58$\times$) than the model from Bonsai.
-Since the Bonsai child model is small enough, we can perform PPA on it, resulting in improved accuracy (8.89ppl) with unchanged speedup.
-Finetuning the semi-sparse Wanda 2:4 model is unfortunately less straightforward.
-It would require similar memory resources as needed to finetune the parent model222though the child tensors are sparse, the resulting gradients and cached intermediate tensors can be dense and have the same dimensions as those of the parent model (say $M\times M$).
-Since Bonsai does structured pruning, the actual tensor sizes are shrunk (say $N\times N~{}|~{}N<M$) which reduces memory during backward passes.
-but we are in the setting where we do not have enough memory for this.
-We therefore have to use a parameter efficient fine-tuning method like LoRA (Hu et al., 2021) instead.
-While the performance gap can be bridged by LoRA finetuning (10.52 $\rightarrow$ 8.34), the adapted semi-structured model experiences a drastic slowdown (0.75$\times$), since the learned low-rank matrices cannot be merged with the original sparsified ones without reverting back to dense computation.
-Thus, LoRA fine-tuned Wanda 2:4 is twice slower (0.48$\sim$ 0.48$\times$ 0.48 ×) than the model from Bonsai and similarly accurate.
-In a memory-constrained setting, practitioners could opt for a pre-existing model of the target size instead of pruning a larger model.
-We compare the Bonsai-pruned model to Phi-2 (Li et al., 2023), a strong representative pre-existing model of similar size.
-As can be seen in Table 2, Bonsai is able to generate a model that is as accurate (0.2 difference in ppl) yet significantly faster (1.58$\times$ vs.
-1.24$\times$ speedup), thus making it a competitive option to consider even if a model already exists at the target size.
-3.2.2 Bonsai is competitive with gradient based structured pruning
-Table 3: LLaMA-1 (50% sparsity) after post-pruning adaptation. are results as reported by (Zhang et al., 2023).
+LlaMA-2 7B $5.11$ $1\times$
+Phi-2 3B $8.69$ $1.24\times$
+Wanda 2:4 3B $10.52$ $1.14\times$
+[…]
+Bonsai 3B $19.47$ $\bm{1.58}\times$
+  + PPA $8.89$ $\bm{1.58}\times$
+4.1 Bonsai is competitive with semi-structured pruning methods
+We compare Bonsai to the semi-structured variant of Wanda (Sun et al., 2023). In general, structured pruning under-perform semi-structured pruning, but compensate for this in speedup.
+[…]
+[…]
+[…]
+4.2 Bonsai is competitive with gradient-based structured pruning
+[…]
 Method Wikitext-2 $\downarrow$ BoolQ HellaSwag WinoGrande ARC-e ARC-c Average $\uparrow$
@@
 LLM-Pruner (Ma et al., 2023) 16.41 60.28 47.06 53.43 45.96 29.18 47.18
-LoRAPrune (Zhang et al., 2023) 11.60 61.88 47.86 55.01 45.13 31.62 48.30
+LoRAPrune (Zhang et al., 2024) 11.60 61.88 47.86 55.01 45.13 31.62 48.30
 Bonsai + PPA 10.92 67.22 43.09 61.64 54.92 26.28 50.63
-We compare Bonsai to the following gradient-based structured pruning approaches: LLM-Pruner (Ma et al., 2023) and LoRA-Prune Zhang et al.
+Next we compare Bonsai to the following gradient-based structured pruning approaches: LLM-Pruner (Ma et al., 2023) and LoRA-Prune Zhang et al.
 (2024).
-We use the reported results from (Zhang et al., 2023) since none of these methods are actually runnable in our memory-constrained setting (Table 1).
-We choose to compare to these over Sheared LlaMA (Xia et al., 2023) since they have much lower memory requirements (Table 1).
+We use the reported results from Zhang et al.
+(2024) since none of these methods are runnable in our memory-constrained setting (Table 2).
+We choose to compare to these over Sheared LlaMA (Xia et al., 2024) since they have much lower memory requirements (Table 2).
 We prune the LlaMA-1 7B model (Touvron et al., 2023) to 50% sparsity since these approaches report their results for the LlaMA-1 model only.
@@
 For the Eleuther Harness tasks, we use language modeling performance on the C4 (Raffel et al., 2020) dataset as pruning signal.
-We also perform parameter-efficient finetuning on our pruned model using 30K 512-length sequences from this corpus.
+We also perform parameter-efficient fine-tuning on our pruned model using 30K 512-length sequences from this corpus.
 Bonsai and LoRAPrune use similar amounts of the C4 dataset for Table 4 (30K vs 20K samples, respectively) whilst LLM-Pruner is trained on instruction tuned data with nearly twice the amount of unique samples (50K).
-Find more details in Appendix C.4.
-As evinced by Table 3, Bonsai outperforms gradient-based methods even though it exclusively uses forward passes in the pruning stage.
+Find more details in Appendix A.3.
+As seen in Table 4, Bonsai outperforms gradient-based methods even though it exclusively uses forward passes in the pruning stage.
 We attribute the superior performance of Bonsai to the fact that its pruning decisions are informed by directly exploring the space of sub-models whilst the other approaches resort to inaccurate proxies of module relevance in order to reduce the memory overhead of a fully gradient-based optimization approach (though not by enough to be runnable in our setting).
-3.2.3 Bonsai can produce compressed models with strong zero-shot abilities
+4.3 Bonsai can produce compressed models with strong zero-shot abilities
 Table 5: Phi-2 pruned to 35% sparsity and fine-tune the pruned model on small amount of the C4.
@@
 Phi-1.5 (Li et al., 2023) 1.5B 12.43 52.9 72.22 63.79 40.89 43.89 54.74
-Sheared LlaMA (Xia et al., 2023) 1.3B Not Reported 33.5 57.9 60.7 Not Reported 25.7 *
+Sheared LlaMA (Xia et al., 2024) 1.3B Not Reported 33.5 57.9 60.7 Not Reported 25.7 *
 Bonsai (w PPA) 1.8B 6.37 47.44 68.35 65.09 42.20 40.53 52.72
   + Reasoning Tuning 1.8B 27.67 45.56 68.82 64.51 42.58 40.97 52.49
-Considerable amounts of compute and data, beyond what is feasible for lay ML practitioners, are needed to train LLMs with strong zero-shot capabilities (AI, 2023; Gemini Team et al., 2023).
+Considerable amounts of compute and data, beyond what is feasible for many practitioners, are needed to train LLMs with strong zero-shot capabilities (AI, 2023; Gemini Team et al., 2023).
 In this section, we demonstrate that Bonsai can empower everyday practitioners to produce strong and compact models with competitive zero-shot abilities by simply pruning bigger models on their available hardware.
 We use Bonsai to prune a $\approx$3B Phi-2 model to $\approx$1.8B (35% sparsity).
-Bonsai hyper-parameters in this experiment are in Appendix C.5.
+Bonsai hyper-parameters in this experiment are in Appendix A.4.
 Since it is small, the 1.8B pruned model can be fully fine-tuned on 1 A6000 GPU over 100k sequences of 2,048 tokens from the C4 dataset.
 As can be seen from Table 5, our pruned model achieves strong zero-shot performance on the Hugging Face OpenLLM leaderboard (Gao et al., 2023) compared to Phi-1.5, a smaller version in the Phi series of models that was trained from scratch.
-Figure 4: LLaMA-2 7B @ 50% sparsity (No PPA). Pertubation and regression components are both needed to make Bonsai effective. Experiment details in Appendix D.
-Interestingly, one exception to the general trend of Bonsai’s strong performance is the GSM-8K dataset, which is a mathematical reasoning dataset that requires generation of long reasoning chains.
-In our experiments, Bonsai prunes with respect to language modeling likelihood, as opposed to reasoning accuracy, and we posit that this mismatch may have contributed to our model’s underperformance on GSM8K.
-We attempt to remedy the drop in reasoning ability by including the GSM8K training data during post-pruning finetuning (resulting in a total of 108K fine-tuning samples).
+Interestingly, one exception to the general trend of Bonsai’s strong performance is GSM8K, a mathematical reasoning dataset that requires generation of long reasoning chains.
+In our experiments, GSM8K performance is initially lower, consistent with task-agnostic pruning behavior in prior work (Xia et al., 2024; Reda et al., 2025); pruning based on language modeling can deprioritize modules needed for specialized reasoning.
+We attempt to remedy the drop in reasoning ability by adding 8K GSM8K samples during PPA (resulting in a total of 108K fine-tuning samples).
 This boosts our model’s performance on the GSM8K with almost no degradation of performance on the other tasks.
+4.4 Bonsai is competitive with other forward pass-only, structured pruning methods
+Table 6: Perplexity at 50% sparsity of LlaMA-{1,2}-7B on Wikitext-2 and C4.
+Dataset Method Sparsity LlaMA-1/seqlen=1024 LlaMA-2/seqlen=1024
+Wikitext-2 Base Model 0% 5.68 5.11
+FLAP 50% 17.27 14.49
+Bonsai 50% 15.72 12.38
+C4 Base Model 0% 7.34 7.04
+FLAP 50% 24.98 24.0
+Bonsai 50% 22.31 20.7
+We focus our last set of experiments on comparing structured pruning methods that can be run without gradient-based optimization.
+We prune the LlaMA-2 7B model (Touvron et al., 2023) to 50% sparsity and evaluate on the Wikitext-2 (Merity et al., 2016) validation dataset.
+Our module importance signal for pruning, $U$, is the language modeling performance on the training set.
+When measuring speedups, we consider end-to-end latency of running inference on model.sequence_length chunks of the Wikitext-2 validation set.
+See Table 11 (Appendix A.2) for details about the hyper-parameters used for these experiments.
+Figure 1 shows that at 50% sparsity, Bonsai achieves the lowest perplexity among methods accessible in memory-constrained settings ($\leq 48$GB).
+Bonsai expectedly outperforms the structured variant of Wanda (Sun et al., 2023), producing models with much lower perplexity at any fixed desired speedup over the parent model.
+A more competitive baseline is FLAP (An et al., 2024), which can be seen as a special case of our more general framework; here (as shown in Table 6),  Bonsai outperforms FLAP while maintaining comparable speedups.
+A key advantage of  Bonsai is its flexibility—it can incorporate various pruning methods as informative priors.
+In practice, we explored three classes of priors: activation magnitude (measuring the average size of activations per module), Wanda (a weight-activation product metric shown to be effective in unstructured pruing), and fluctuation-based metrics (as in FLAP, which track statistical variation across activations.
+These priors provide a a quick, forward-pass-only signal of which modules are likely to be important, and Bonsai’s regression step then refines them into more accurate importance estimates.
+See Appendix E for full details on how we estimate priors.
+An important outcome of this is that Bonsai offers practitioners flexible runtime-quality tradeoffs.
+While the reported results use a configuration that runs in $\leq~4hours$ (which is considerably more than FLAP’s $\approx 1$ hour runtime), practitioners can configure  Bonsai  to produce structurally pruned models in as little as 15 minutes by reducing the number of perturbations, sample size, or pruning iterations, potentially at some cost to performance.
+Conversely, allocating more compute time generally yields better pruning outcomes.
+This flexibility allows practitioners to tailor Bonsai to their specific constraints, unlike methods with fixed computational profiles.
 5 Analysis
-We conduct ablative experiments to understand the impact of the ingredients from Section 2.
+Here we conduct ablative experiments to understand the impact of the ingredients from Section 3.
 Do we need both perturbative and regressive components of Bonsai?
+Figure 3: LLaMA-2 7B @ 50% sparsity (No PPA). Perturbation and regression components are both needed to make Bonsai effective. Experiment details in Appendix B.
 Figure 3 shows that both components are key to obtaining a good pruned model.
@@
 PPL ($\downarrow$) NaN $61.63$ $22.09$
-Table 5: Wikitext-2 perplexity of LLaMA-2 7B @ 50% sparsity (No PPA). We vary the number of perturbative evaluations. Details in Appendix I.
+Table 7: Wikitext-2 perplexity of LLaMA-2 7B @ 50% sparsity (No PPA). We vary the number of perturbative evaluations. Details in Appendix F.
 How much performance is recovered by post-pruning adaptation?
@@
 If the pruned model is small enough in size, we may perform post pruning adaptation to recover more performance, as can be seen from Table 8.
-Table 6: Impact of PPA on LLaMA-2 7B @ 50% sparsity. Details in Appendix G.
+Table 8: Impact of PPA on LLaMA-2 7B @ 50% sparsity. Details in Appendix D.
 Method Wikitext-2 PPL
@@
    +  Distillation $8.89$
-Figure 5: LLaMA-2 7B pruned to 50% sparsity. See Appendix H for experiment details and definitions of Wanda and Activation Magnitude priors.
+Figure 4: LLaMA-2 7B pruned to 50% sparsity. See Appendix E for experiment details and definitions of Wanda and Activation Magnitude priors.
 What is the impact of the choice of metric for the prior $\rho$?
@@
 This indicates that Wanda is a strong signal for efficiently estimating the importance of model units.
-5 Related Work
-Here we discuss relevant work in LLM pruning, and defer a longer discussion compression methods to Appendix B.
-Unstructured pruning:
-Whilst structured pruning approaches remove entire model components like layers (Xu et al., 2020; Xia et al., 2022b), dimensions of linear layers (Wang et al., 2019) or attention heads (Michel et al., 2019; Held & Yang, 2022), unstructured pruning (Han et al., 2015; Frankle & Carbin, 2018; Benbaki et al., 2023; Sun et al., 2023) removes individual parameters of the model.
-These approaches achieve memory savings by inducing sparsity in the model weights, but they generally do not result in actual model speedups except when specialized hardware is available (Mishra et al., 2021).
-Proposed semi-structured sparsity methods (Mishra et al., 2021) such as 2:4 and 4:8 patterns do result in faster inference, but the speedup gains they achieve are far from the idealized 2×2$\times$2 ×.
-Structured pruning without backward passes:
-To the best of our knowledge, all existing structured pruning techniques for large (over 1B scale) language models, like LoRAPrune (Zhang et al., 2023) and LLM-Pruner (Ma et al., 2023), are gradient-based.
-Thus, under the memory setting we consider, these methods cannot be applied since their memory requirements well exceed the memory needed for inference on the model to be pruned.
-For smaller language models like BERT, (Nova et al., 2023) recently proposed Kernelized Convex Masking (KCM) for gradient-free structured pruning.
-Unfortunately, to prune a fully connected layer with $K$ intermediate units, KCM requires instantiating a $K\times K$ coefficient matrix (for each layer).
-Whilst this is feasible for BERT models, a typical LLM (AI, 2023; Touvron et al., 2023) has $K\sim 10^{4}-10^{5}$ which would make the size of each per-layer coefficient matrix comparable to the size of the LLM itself.
-In computer vision, there exist several perturbative gradient-free structured pruning techniques (Ancona et al., 2020; Dekhovich et al., 2021).
-However, these methods have been exclusively applied to small vision models (VGG; Simonyan & Zisserman (2014), WideResnet; Zagoruyko & Komodakis (2016)) and, taken as is, would not scale to LLMs.
-We introduce several techniques to make Bonsai scale to large language models.
-A key difference between Bonsai and these perturbative methods is that pruning decisions are made globally instead of on a local layer-by-layer basis (Ancona et al., 2020; Dekhovich et al., 2021).
-A failure mode of layer-by-layer pruning is that if a module, say $m_{i}$, which is in the top $p$ fraction of relevant modules when the whole model is considered but in the bottom $1-p$ fraction within its layer will not be included in the pruned model — a clearly sub-optimal choice.
-Bonsai enables us to obtain globally meaningful estimates of module relevance that are comparable across layers.
+Memory and Runtime Considerations.
+Bonsai requires $\approx 20$GB memory versus 80-160GB for gradient-based methods (Table 2).
+Runtime is $\approx 4$ hours for optimal quality or $\approx 15$ minutes with reduced perturbations.
+Pre-adaptation results appear in Table 6 and 8; note that gradient-based baselines in Table 4 also use adaptations (Section 4.2).
 6 Conclusion, Limitations, and Future Work
 In this work, we have presented Bonsai, a gradient-free structured pruning method that enables efficient compression of LLMs using only forward passes.
-By eliminating the need for backward passes, Bonsai reduces memory requirements by over 2×2$\times$2 ×, allowing practitioners to prune and adapt models on hardware configurations where gradient-based approaches would be infeasible.
-Through extensive experiments, we have demonstrated that Bonsai produces models that are small, fast, and accurate—-outperforming even most of these gradient-based methods while using substantially fewer resources.
+By eliminating the need for backward passes, Bonsai reduces memory requirements by over $2\times$, allowing practitioners to prune and adapt models on hardware configurations where gradient-based approaches would be infeasible.
+Through extensive experiments, we have demonstrated that Bonsai produces models that are small, fast, and accurate—outperforming even most of these gradient-based methods while using substantially fewer resources.
 A primary limitation of Bonsai is its computational runtime.
 As shown in Section 5, performance improves with more sub-model exploration, slower pruning rates, and larger data samples, but these improvements come at the cost of increased runtime.
-While competitors like FLAP complete pruning in $\approx 1$ hour, our configuration of Bonsai that achieves optimal results requires $\leq 4$ hours.
-However, Bonsai offers practitioners the flexibility to configure faster pruning schedules (as quick as 15 minutes) when speed is prioritized over achieving optimal compression quality.
+While competitors like FLAP complete pruning in $\approx 1$ hour, Bonsai’s optimal configuration requires $\leq 4$ hours.
+However, Bonsai offers practitioners the flexibility to configure faster pruning schedules ($\approx$15 minutes) when speed is prioritized over achieving optimal compression quality.
 This runtime-quality tradeoff is a design feature that allows adaptation to various computational constraints.
@@
 Additionally, due to our focus on forward-pass-only operations, Bonsai does not fine-tune the model during iterative pruning.
-Integration with gradient-free approaches like MeZO (Malladi et al., 2023) could allow for continuous updates during the pruning process, potentially leading to better preservation of model performance.
-Appendix A Broader Impact Statement
+Integration with gradient-free approaches like MeZO (Malladi et al., 2023) could allow for continuous updates during the pruning process, preserving model performance.
+MoE models also provide an interesting direction: while they enable sparse activation, all experts must remain in memory.
+Bonsai could prune individual experts to reduce memory footprint while preserving conditional computation benefits.
+Overall, by reducing the memory and computational requirements of structured pruning, Bonsai contributes to the growing trend of practical, efficiency-focused AI research, aligning with the community-wide push to make powerful models more accessible.
+The ability that Bonsai enables to prune and adapt large models on a wider range of hardware configurations fosters innovation beyond resource-rich institutions to represent a crucial move toward a more inclusive, sustainable future for the field.
+7 Acknowlegements
+This work was supported in part by the National Science Foundation grants IIS1705121, IIS1838017, IIS2046613, IIS2112471, and funding from Meta, Morgan Stanley, Amazon, and Google.
+Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of any of these funding agencies.
+We are grateful for helpful feedback from Mingjie Sun, Victor Akinwande, Asher Trockman, Afshin Rostamizadeh and Daniel Glasner.
+8 Broader Impact Statement
 Network pruning is an effective compression strategy to reduce the inference costs of large language models.
@@
 Finally, we note that our work has primarily focused on the axes of efficiency and utility (e.g., perplexity) in assessing performance; recent work has shown that compression may also have outsized effects on issues such as model fairness and robustness, which would be interesting additional aspects to consider in future study.
-Appendix B Background: Additional Methods for LLM Compression
-Although we focus on pruning in this work, we note that prior research has explored various other compression schemes such as distillation (Hinton et al., 2015; Gu et al., 2023) and quantization (Xiao et al., 2023) to create smaller models from larger pre-trained ones.
-Similar to structured pruning, these compression methods themselves often impose significant computational burdens.
-For example, distillation-based techniques require using LLMs to generate large amounts of teacher data (Jiao et al., 2019; Hsieh et al., 2023).
-Although unstructured pruning (Frantar & Alistarh, 2023; Sun et al., 2023) and quantization have lower training-time resource demands, the models they produce either require specialized hardware to achieve speedups (Mishra et al., 2021) or may actually slow down inference due to additional computational overhead (Dettmers et al., 2022).
-Table 7 summarizes these existing limitations across different compression approaches.
-We also cover additional background in distillation and quantization below.
-Table 7: Landscape of resource consumption (memory and compute) of different model compression methods at training time and the inference time resource consumption of the models they deliver.
-✗  $\rightarrow$ a prohibitive cost to the lay practitioner whilst ✓  $\rightarrow$ a viable option with respect to that resource.
-Regime Resource Approaches
-Quantization (Mixed Precision) Distillation Unstructured Pruning Gradient-Based Structured Pruning Bonsai (Ours)
-Train Memory
-Compute
-Inference Memory
-Compute
-Distillation:
-Originally proposed by (Hinton et al., 2015; Sanh et al., 2019) have shown that it is possible to reach similar performances on many downstream tasks using much smaller language models pre-trained with knowledge distillation.
-The resulting models are lighter and faster at inference time.
-Unfortunately, as noted by (Xia et al., 2022b), distilling a pre-trained model to a student model that has been initialized from scratch requires lots of compute and data to recover reasonable levels of teacher performance.
-To reduce this resource burden, (Sanh et al., 2020; Lagunas et al., 2021; Xia et al., 2022b) combine distillation with gradient-based structured pruning, so that the student model does not have to be initialized from scratch.
-Bonsai also leverages distillation, but only during the post-pruning adaptation phase.
-This maintains the constraint that the pruning process is gradient-free.
-And since the pruned model is smaller than the parent model, we can safely perform distillation and parameter efficient fine-tuning (Hu et al., 2021; He et al., 2021) within the bounds of available memory.
-Quantization:
-(Li et al., 2020) introduced a training strategy that showed that heavily compressed, large models achieved higher accuracy than lightly compressed, small models, providing both efficiency and high-accuracy results on NLP tasks at once.
-(Dettmers et al., 2022) introduced a new method called ‘LLM.int8()‘, which allows loading large models with 16 or 32-bit weights and using them immediately for inference without any performance degradation by combining vector-wise quantization with mixed-precision decomposition.
-(Xiao et al., 2023) enabled 8-bit weight, 8-bit activation (W8A8) quantization for LLMs with >100B parameters, which smoothed activation outliers by migrating the quantization difficulty from activations to weights with a mathematically equivalent transformation.
-Quantization tends to be complementary to pruning, allowing independent exploration and subsequent combinations of both techniques.
-This paper focuses on gradient-free structured pruning.
-We leave combining Bonsai with quantization for future work.
-Appendix C Main Experiment Details
-C.1 Full Algorithm
-Algorithm 1 Bonsai
-1:  Input: Model [$\mathbf{M}_{\theta}$], sub-models per iteration [$n_{\mathrm{iter}}$] Sparsity per iteration [$p_{\mathrm{iter}}$], Target sparsity [$p$] Module list [$\mathbf{m}$]
-2:  for $l=1$ to $\lceil\frac{p}{p_{\mathrm{iter}}}\rceil$ do
-3:     $\mathbb{\rho}^{l}\leftarrow$ Calculate unstructured pruning metric for all modules in $\mathbf{m}$
-4:     $\bar{\mathbb{\rho}}^{l}\leftarrow$ Fix the top $(1-2p_{\mathrm{iter}})$ of $\mathbb{\rho}^{l}$ to $\infty$
-5:     Sample $\{\mathbf{\bar{m}}_{i}\}_{[n_{\mathrm{iter}}]}$ sub-models according to $\bar{\mathbb{\rho}}^{l}$
-6:     Run forward pass on each sub-model and compute $U$. Construct $\mathbb{D}^{l}=\left\{\bar{\mathbf{m}}_{i},U_{i}\right\}_{[n_{\mathrm{iter}}]}$
-7:     $\beta^{\l}\leftarrow\mathrm{Regress}\left(\mathbb{D}^{l}\right)$
-8:     $\{m_{\mathrm{pruned}}\}\leftarrow$ sort $\beta^{\l}$ and drop the bottom $k$ modules that make up $p_{\mathrm{iter}}$ fraction of the model.
-9:     $\mathbf{m}\leftarrow$ update module list to exclude $\{m_{\mathrm{pruned}}\}$
-10:  end for
-11:  Output: Pruned model $\mathbf{M}_{|\mathbf{m}}$
-A note about Line $5$ in our algorithm.
-Depending on the task, we find that sampling $\mathbf{\bar{m}}_{i}$ and its complement $\mathbf{\bar{m}}^{c}_{i}$ helps reduce the variance of our regression estimate and leads to better results.
-C.2 Hyper-parameters for all Bonsai regression during pruning
-When using Bonsai, we estimate $\beta$ from $\mathbb{D}^{l}$ by performing linear regression via gradient descent with Adam (Kingma & Ba, 2014).
+Appendix A Main Experiment Details
+A.1 Hyper-parameters for all Bonsai regression during pruning
+When using Bonsai, we estimate $\beta$ from $\mathbb{D}^{l}$ by performing linear regression via gradient descent with Adam (Kingma and Ba, 2014).
 We cross-validate over the following set of hyper-parameters.
@@
 {100, 0, 1e-4} {100, 10, 1, 0.1} {32, 64, 128} 50
-During cross validation, we choose the model whose predictions have the best Kendall rank correlation co-efficient (Kendall, 1948) with the target.
-We do this because we do not care about matching $U_{k}$ exactly for each sub-model $k$; we rather care that our learned $\beta$ predicts the correct rankings amounts sub-models, which would denote that $\beta$ reasonably models relative module importances.
-In general, we use $\ell_{1}$-norm regularization on $\beta$ for all experiments. For the Phi-2 experiment in Section 3.2.3, we find that $\ell_{2}$-norm works better.
-C.3 Forward Pass Only / Semi-structured pruning Experiments
-Table 10 show the Bonsai hyperparameters we used for the experiments in Section 3.2.1.
+During cross validation, we choose the model whose predictions have the best Kendall rank correlation coefficient (Kendall, 1948) with the target.
+We do this because we do not care about matching $U_{k}$ exactly for each sub-model $k$; we rather care that our learned $\beta$ predicts the correct rankings amoung sub-models, which would denote that $\beta$ reasonably models relative module importances.
+In general, we use $\ell_{1}$-norm regularization on $\beta$ for all experiments. For the Phi-2 experiment in Section 4.3, we find that $\ell_{2}$-norm works better.
+A.2 Forward Pass Only / Semi-structured pruning Experiments
+Table 11 shows the Bonsai hyperparameters we used for the experiments in Section 4.1.
 Table 10: Bonsai hyper-params for forward only experiments
@@
 LR rank LoRA-$\alpha$ $\lambda$ (Distill Weight) LoRA Modules
-1e-4 128 4$\times$×rank 0.01 All Modules
+[…]
 For Wanda(Sun et al., 2023), we use the default hyper-parameters specified by the paper repo here for pruning.
@@
 For fine-tuning the Phi-2 model on Wikitext-2, we use the same hyper-parameters as Bonsai in Table 11.
-C.4 Experiments comparing to Gradient based structured pruning
+A.3 Experiments comparing to Gradient based structured pruning
 We compare to LoRA-Prune and LLM-Pruner.
@@
 Whilst we use 1 A6000 GPU (48G) for all experiments, LoRA-Prune uses A100 GPU (80G) for pruning LLaMA-1 7B.
-All Bonsai hyper-parameters are the same as Appendix C.3 except for $\mathrm{ns}_{\mathrm{sub-models}}$ which we set to $1000$.
-C.5 Phi-2 pruning experiment details
-For the experiment in Section 3.2.3, All Bonsai hyper-parameters are the same as Appendix C.3 except for the following changes:
+All Bonsai hyper-parameters are the same as Appendix A.2 except for $\mathrm{ns}_{\mathrm{sub-models}}$ which we set to $1000$.
+A.4 Phi-2 pruning experiment details
+For the experiment in Section 4.3, All Bonsai hyper-parameters are the same as Appendix A.2 except for the following changes:
   • $\mathrm{ns}_{\mathrm{sub-models}}=2000$
@@
   • Due to its relatively small size, the 1.8B pruned model can be fully fine-tuned on a single A6000 GPU over 100k sequences of length 2,048 tokens from the C4 dataset instead of using LoRA.
-Appendix D Impact of regression and perturbation ablation details
-For the experiment in Section 6, All Bonsai hyper-parameters are the same as Appendix C.3 except $p_{\mathrm{iter}}=0.1$ to speed up pruning.
+Appendix B Impact of regression and perturbation ablation details
+For the experiment in Appendix I, All Bonsai hyper-parameters are the same as Appendix A.2 except $p_{\mathrm{iter}}=0.1$ to speed up pruning.
 A simple alternative to Bonsai is to leverage the prior $\rho$, computed from the unperturbed parent model, and make pruning decisions exclusively according to this.
@@
 Note that we use a Wanda based metric to define $\rho$ for this experiment.
-Module level analogues of the unstructured pruning metrics we explore are defined in Appendix H.
+module-level analogues of the unstructured pruning metrics we explore are defined in Appendix E.
 Table 12: Experiment on linear regression to estimate module importances. Wikitext-2 Perplexity. LLaMA-2 7B pruned to 50% sparsity
-Linear Regression Relative Speepdup w/o Post-Pruning Adaptation w Post-Pruning Adaptation
+Linear Regression Relative Speedup w/o Post-Pruning Adaptation w Post-Pruning Adaptation
 No $2.06$ $146.57$ $9.68$
 Yes $1.77$ $61.63$ $9.15$
-Appendix E Varying the pruning fraction per-iteration
-For the experiment in Section 6, All Bonsai hyper-parameters are the same as Appendix C.3 except we vary $p_{\mathrm{iter}}$.
+Appendix C Varying the pruning fraction per-iteration
+For the experiment in Appendix I, All Bonsai hyper-parameters are the same as Appendix A.2 except we vary $p_{\mathrm{iter}}$.
 Table 13: Varying the fraction pruned at a time. Wikitext-2 Perplexity. LLaMA-2 7B pruned to 50% sparsity
-Prune Frac Relative Speepdup w/o Post-Pruning Adaptation w Post-Pruning Adaptation
+Prune Frac Relative Speedup w/o Post-Pruning Adaptation w Post-Pruning Adaptation
 0.05 $1.58$ $19.47$ $8.89$
@@
 0.20 $1.67$ $209.44$ $9.57$
-Appendix F Varying the number of calibration data points for pruning
-All Bonsai hyper-parameters are the same as Appendix C.3 except we vary $\mathrm{ns}_{\mathrm{data}}$ and $p_{\mathrm{iter}}=0.1$ to speed up pruning.
-Table 13: How many data-points to consider during forward passes. Wikitext-2 Perplexity. Llama-2 7B pruned to 50% sparsity
+Appendix D Varying the number of calibration data points for pruning & PPA
+For these two categories of experiments, all Bonsai hyper-parameters are the same as Appendix A.2 except we vary $\mathrm{ns}_{\mathrm{data}}$ and $p_{\mathrm{iter}}=0.1$ to speed up pruning in the former.
+Table 14: How many data points to consider during forward passes. Wikitext-2 Perplexity. Llama-2 7B pruned to 50% sparsity
 $\mathrm{ns}_{\mathrm{data}}$ w/o Adapt w Adapt
@@
 32 $61.63$ $9.15$
-Appendix G Post-pruning adaptation
-For this experiment, All Bonsai hyper-parameters are the same as Appendix C.3.
-Appendix H Impact of prior
-For this experiment, All Bonsai hyper-parameters are the same as Appendix C.3 except we vary $\rho$.
-H.1 $\rho$ is Activation Magnitude
+Appendix E Impact of prior
+For this experiment, All Bonsai hyper-parameters are the same as Appendix A.2 except we vary $\rho$.
+E.1 $\rho$ is Activation Magnitude
 MLP / Fully Connected Module: Let $d$ be the intermediate dimension of the MLP to be pruned.
@@
 $\left(\rho\in\mathbb{R}^{d}\right)\propto\hat{\mathbf{a}}=\frac{1}{B}\sum_{b}\mathrm{Mean}\bigg(\big|\mathbf{a}_{b}\big|,\mathrm{axis=}0\bigg)$ (4)
-Self-Attention Module: For any data-sample sequence $b$, the output of the self attention module before the final output projection is $\mathbf{a}\in\mathbb{R}^{B\times S\times d_{h}\times h}$ where $h$ is the number of attention heads and $d_{h}$ is the size of each head’s output.
+Self-Attention Module: For any data-sample sequence $b$, the output of the self-attention module before the final output projection is $\mathbf{a}\in\mathbb{R}^{B\times S\times d_{h}\times h}$ where $h$ is the number of attention heads and $d_{h}$ is the size of each head’s output.
 We can flatten $\mathbf{a}\in\mathbb{R}^{B\times S\times d_{h}\times h}\rightarrow\mathbb{R}^{BSd_{h}\times h}$ and then use the same formula as Equation 5 above to calculate $\rho$.
 $\left(\rho\in\mathbb{R}^{h}\right)\propto\hat{\mathbf{a}}$ (5)
-H.2 $\rho$ is Wanda (Sun et al., 2023)
+E.2 $\rho$ is Wanda (Sun et al., 2023)
 MLP / Fully Connected Module: Let $d$ be the intermediate dimension of the MLP to be pruned.
 Let $W\in\mathbb{R}^{d\times o}$ be the output projection matrix for the MLP.
-For any data-sample sequence $b$, we flatten model activation before the final output, $\mathbf{a}\in\mathbb{R}^{B\times S\times d}\rightarrow\mathbb{R}^{BS\times d}$ and then compute the following metric which is a module level analogue of Wanda:
+For any data-sample sequence $b$, we flatten model activation before the final output, $\mathbf{a}\in\mathbb{R}^{B\times S\times d}\rightarrow\mathbb{R}^{BS\times d}$ and then compute the following metric which is a module-level analogue of Wanda:
 $\begin{split}\left(\rho\in\mathbb{R}^{d}\right)\propto\hat{\mathbf{a}}&=\frac{1}{o}\sum_{o}\mathbf{a}^{o}\\ \mathbf{a}^{o}&=\bigg|W[:,o]\bigg|\odot\mathrm{RootMeanSquare}\bigg(\mathbf{a},\mathrm{axis=}0\bigg)\end{split}$ (6)
 Self-Attention Module: Let $W\in\mathbb{R}^{d\times o}$ be the output projection matrix for the self-attention module.
-For any data-sample sequence $b$, the output of the self attention module before the final output projection is $\mathbf{a}\in\mathbb{R}^{B\times S\times d_{h}\times h}$ where $h$ is the number of attention heads and $d_{h}$ is the size of each head’s output.
+For any data-sample sequence $b$, the output of the self-attention module before the final output projection is $\mathbf{a}\in\mathbb{R}^{B\times S\times d_{h}\times h}$ where $h$ is the number of attention heads and $d_{h}$ is the size of each head’s output.
 We can flatten $\mathbf{a}\in\mathbb{R}^{B\times S\times d_{h}\times h}\rightarrow\mathbb{R}^{BSd_{h}\times h}$ and then use the same formula as Equation 5 above to calculate $\rho\in\mathbb{R}^{h}$.
-H.3 $\rho$ is Fluctuation-based (FLAP) An et al. (2024)
+E.3 $\rho$ is Fluctuation-based (FLAP) (An et al., 2024)
 MLP / Fully Connected Module: Let $d$ be the intermediate dimension of the MLP to be pruned.
@@
 For any data-sample sequence $b$, we compute the sample variance of each input feature across batches and weight it with the squared norm of the corresponding column of the weight matrix:
-$\left(\rho\in\mathbb{R}^{d}\right)\propto S_{:,j}=\frac{1}{N-1}\sum_{n=1}^{N}(X_{n,j,:}-\bar{X}_{:,j,:})^{2}\cdot||W_{:,j}||_{2}^{2}$ (7)
-where $\bar{X}_{:,j,:}$ represents the average of the $j$-th channel for all samples, and $||W_{:,j}||_{2}^{2}$ denotes the squared norm of the $j$-th column of the weight matrix.
+[…]
+where $\bar{X}_{:,j,:}$ represents the average of the $j$-th channel for all samples, and $W_{:,j}||_{2}^{2}$ denotes the squared norm of the $j$-th column of the weight matrix.
 Self-Attention Module: We compute the fluctuation at the head level, weighted by the corresponding weights in the output projection matrix:
-$\left(\rho\in\mathbb{R}^{h}\right)\propto S_{:,j}=\frac{1}{N-1}\sum_{n=1}^{N}(X_{n,j,:}-\bar{X}_{:,j,:})^{2}\cdot||W_{:,j}||_{2}^{2}$ (8)
+[…]
 For our improved fluctuation metric (fluct.2.0), we track the mean and second moment separately over multiple batches, computing variance as:
@@
 This approach provides greater numerical stability when accumulating statistics across multiple forward passes.
-Appendix I How many perturbative samples are reasonable?
-For this experiment, All Bonsai hyper-parameters are the same as Appendix C.3 except $p_{\mathrm{iter}}=0.1$ to speed up pruning.
+Appendix F How many perturbative samples are reasonable?
+For this experiment, All Bonsai hyper-parameters are the same as Appendix A.2 except $p_{\mathrm{iter}}=0.1$ to speed up pruning.
 Table 15: Varying the number of sub-models generated. Wikitext-2 Perplexity. LLaMA-2 7B pruned to 50% sparsity
@@
 50 NaN $9.24$
-Using $\mathrm{ns}_{\mathrm{sub-models}}=50$ results in an model with NaN perplexity on the Wikitext validation set.
-We posit that this is because of the LLaMA models are half precision, and removing the wrong modules can result in activations going outside of the FP16 dynamic range for unique data-points.
+Using $\mathrm{ns}_{\mathrm{sub-models}}=50$ results in a model with NaN perplexity on the Wikitext validation set.
+We posit that this is because of the LLaMA models are in half precision, and removing the wrong modules can result in activations going outside of the FP16 dynamic range for unique data points.
 Note that we are able to recover good performance of the model after fine-tuning though (we do not observe NaNs with the Wikitext-2 training data).
 This indicates that Bonsai actually recovers good modules even using as few samples as 50 sub-models.
-Appendix J Mistral-7B Experiment Details
+Appendix G LlaMA-3-8B Experiments
+In our efforts toward generalization, we also validate Bonsai on LLaMA-3-8B and use the same experimental setup as our LLaMA-2-7B experiments (hyperparameters in Appendix A.2) and prune to 50% sparsity on Wikitext-2.
+Table 16: LLaMA-3-8B at 50% sparsity on Wikitext-2
+Method Wikitext-2 PPL
+LLaMA-3-8B (base) 6.14
+Bonsai (w/o PPA) 21.36
+Bonsai (w/ PPA) 10.37
+The results show comparable trends to LLaMA-2-7B in the main text: Bonsai operates within accessible memory budgets while achieving competitive accuracy after post-pruning adaptation.
+This confirms that Bonsai’s approach generalizes across LLaMA model versions.
+Appendix H Mistral-7B Experiment Details
 In addition to the primary experiments on the LLaMA and Phi-2 models, supplementary experiments were performed on the Mistral-7B Jiang et al.
@@
 We factor these differences into consideration in the implementation of Bonsai for Mistral.
-For the experiments that produced the results below, all Bonsai hyper-parameters are the same as Appendix C.3.
-Table 15 presents the test perplexity results for Mistral-7B under different pruning methods.
-Considering the fully-structured sparsity nature of Bonsai, it achieves a test perplexity of 47.5 without post-pruning adaptation, with 1.66$\times$ inference speedup.
-After performing post-pruning adaptation on our pruned Mistral-7B, perplexity dropped drastically to 10.08.
-Note that the reported results of Wanda-pruned Mistral-7B are not fine-tuned afterward; if they were, their results would be marginally better than Bonsai’s results.
-However, as shown in Table 2, latency speedup would have dropped rapidly, while Bonsai stays the same at 1.66$\times$.
+For the experiments that produced the results below, all Bonsai hyper-parameters are the same as Appendix A.2.
+[…]
 Table 17: Test perplexity of Mistral-7B model on Wikitext-2 across fully-structured Bonsai and semi-structured Wanda methods.
@@
 We notice a recurring theme: when an attention layer is significantly altered, it leads to compensation in the next layers within the sequence.
-This adaptive behavior, termed the ”Hydra effect” by (McGrath et al., 2023), implies that the layers within a language model interact in a way that changes in one layer prompt adjustments in another.
+This adaptive behavior, termed the "Hydra effect" by (McGrath et al., 2023), implies that the layers within a language model interact in a way that changes in one layer prompt adjustments in another.
 (McGrath et al., 2023) specifically mentioned that when one attention layer was removed from a language model, the model was still able to self-repair and produce similar outputs; but it did so by relying more heavily on other layers.
 Figure 5: Mistral’s pruned attention layers.
-The heavily pruned layers are usually preceded by or sandwiched between lightly-pruned layers, exhibiting the self-repairing ”Hydra effect” McGrath et al.
+The heavily pruned layers are usually preceded by or sandwiched between lightly-pruned layers, exhibiting the self-repairing "Hydra effect" McGrath et al.
 (2023).
-Should Bonsai prune iteratively?
+Appendix I Should Bonsai prune iteratively?
 Table 18 demonstrates the benefits of using Bonsai in an iterative fashion.
@@
 The performance gap between values of $p_{\mathrm{iter}}$ persists even after post-pruning adaptation, indicating that slower pruning allows for more accurate estimates of module importance.
-Table 16: Varying $p_{\mathrm{iter}}$. Wikitext-2 perplexity of LLaMA-2 7B pruned to 50% sparsity. See Appendix E for experiment details.
+Table 18: Varying $p_{\mathrm{iter}}$. Wikitext-2 perplexity of LLaMA-2 7B pruned to 50% sparsity. See Appendix C for experiment details.
 $p_{\mathrm{iter}}=0.05$ $p_{\mathrm{iter}}=0.1$ $p_{\mathrm{iter}}=0.2$